### v1

# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

Abstract.
We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are proper subgroups of the associated affine Weyl groups.
Key words and phrases:
Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group
1. Introduction
The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study of the relation of integrable systems and 2-dimensional topological field theories [8, 9]. This geometric structure and its generalizations play a central role in topological field theory, Gromov-Witten theory, singularity theory and in the theory of integrable systems (see [1, 2, 3, 4, 5, 9, 13, 15, 18, 19, 22, 24] and the references therein).
We consider in this paper the construction of a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups. These structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector field. The simplest example in this class corresponds to the affine Weyl group of type $A_{1}$, which naturally appears in the study of Hodge integrals and their remarkable connections to two important discrete integrable systems: the Volterra hierarchy and the $q$-deformed KdV hierarchy—both fundamental in soliton theory [21, 22].
Our approach parallels the construction of a Frobenius manifold structure on the orbit space of a Coxeter group or on a certain extension of an affine Weyl group [9, 12, 14, 25, 26, 28, 29]. For these cases, a flat pencil of metrics, based on which the Frobenius manifold structure is constructed, is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space. The key to these constructions lies in the linear dependence of both the contravariant components of the metric and its Levi-Civita connection on a distinguished coordinate of the orbit space—this special coordinate naturally gives rise to the flat unit vector field of the Frobenius manifold. However, for the generalized Frobenius manifolds that we are considering, the unit vector field is not flat, implying the absence of such a special coordinate. To overcome this, we proceed as follows: Given an affine Weyl group $W_{a}$ acting on an $\ell$-dimensional Euclidean space $V$, we introduce a parameter $\lambda$ into the invariant Fourier polynomials, yielding the coordinates $z^{1},\dots,z^{\ell}$ on the orbit space. The projected contravariant metric $\left(g^{\alpha\beta}(z)\right)$ then depends on $\lambda$. The key idea is to choose the $\lambda$-dependence of the invariant Fourier polynomials such that $\left(g^{\alpha\beta}(z)\right)$ varies linearly with $\lambda$.
In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds (including their monodromy groups), and provide some explicit examples. Further examples will appear in a subsequent paper [20]. The paper is organized as follows. In Sect. 2, we introduce the invariant $\lambda$-Fourier polynomial ring and study its properties. In Sect. 3, we construct generalized Frobenius manifold structures on orbit spaces of affine Weyl groups, and summarize the main results in Theorem 3.9. In Sect. 4, we analyze properties of the monodromy groups of these generalized Frobenius manifolds. In Sect. 5, we present some representative examples. In Sect. 6, we conclude the paper with remarks and further directions.
2. Invariant $\lambda$-Fourier polynomial rings and proper generators
Let us begin with several basic settings.
2.1. Affine coordinates
[…]
$\alpha_{j}^{\vee}=\frac{2\alpha_{j}}{(\alpha_{j},\alpha_{j})},\quad j=1,\dots,\ell,$
then the fundamental weights $\omega_{1},\dots,\omega_{\ell}\in V$ of $R$ are defined by the relations
$(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij},\quad i,j=1,\dots,\ell.$
We fix a weight
$\omega=\sum_{j=1}^{\ell}m_{j}\omega_{j},\quad m_{j}\in\mathbb{Z}_{\geq 0}$ (2.1)
of $R$, and introduce affine coordinates $(x^{1},\dots,x^{\ell};c)$ on $V$ by
$\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee},$ (2.2)
where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of elements of $V$:
$\mathbf{x}=c\omega+\hat{\mathbf{x}}=(x^{1}+\theta_{1}c)\alpha_{1}^{\vee}+\dots+(x^{\ell}+\theta_{\ell}c)\alpha_{\ell}^{\vee},$ (2.3)
with $\hat{\mathbf{x}}=x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$, and
$\theta_{j}=(\omega_{j},\omega),~{}j=1,\dots,\ell.$ (2.4)
Denote by $W_{a}(R)$ the affine Weyl group of $R$. It acts on $V$ by affine transformations
$\mathbf{x}\mapsto\sigma(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\sigma\in W,\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
where $W=W(R)$ is the Weyl group of $R$ generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha_{j}^{\vee},\mathbf{x})\alpha_{j},\quad\mathbf{x}\in V,\ j=1,\dots,\ell.$
The action
$\sigma\left(c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}\right)=c\omega+\tilde{x}^{1}\alpha_{1}^{\vee}+\dots+\tilde{x}^{\ell}\alpha_{\ell}^{\vee}$ (2.6)
of $\sigma\in W_{a}(R)$ on $V$ yields the change of affine coordinates
$\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c).$
2.2. The Invariant $\lambda$-Fourier polynomial ring and its generators
Let us consider the Fourier function ring
$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{m(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}\right\}.$
We introduce a gradation on $\mathscr{F}$ by the assignment
$\deg\mathrm{e}^{2\pi ix^{j}}=\theta_{j},\quad\deg\mathrm{e}^{2\pi ic}=-1,$ (2.7)
where $\theta_{j}$ is defined in (2.4). Then for any $t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}$ and $\zeta\in\mathbb{C}$ we have
$\deg\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}+\zeta)}=-t_{0}+t_{1}\theta_{1}+\dots+t_{\ell}\theta_{\ell}.$ (2.8)
In other words, for any given linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and constant $m_{0}\in\mathbb{C}^{\times}$ we have
$\deg\left(m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}\right)=\psi(\theta_{1},\dots,\theta_{\ell};-1).$ (2.9)
The action of $W_{a}(R)$ on the affine coordinates of $V$ that is defined by (2.6) leads to a right action of $W_{a}(R)$ on $\mathscr{F}$, which is given by
$\sigma:~{}\mathscr{F}\to\mathscr{F},~{}f(x^{1},\dots,x^{\ell};c)\mapsto(f\circ\sigma)(x^{1},\dots,x^{\ell};c)$ (2.10)
for any given $\sigma\in W_{a}(R)$.
Lemma 2.1.
The affine Weyl group $W_{a}(R)$ acts homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
Proof.
Let $m(x^{1},\dots,x^{\ell};c)$ be a nonzero monimial, which can be represented in the form
$m(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$
with a linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and a nonzero constant $m_{0}$. For any given translation
$\sigma:\mathbf{x}\mapsto\mathbf{x}+\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
of $W_{a}(R)$, its action on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is given by
$(x^{1},\dots,x^{\ell};c)\mapsto(x^{1}+n_{1},\dots,x^{\ell}+n_{\ell};c),$
so the associated action on $m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$ yields $m_{0}^{\prime}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$, which preserves the linear function $\psi(x^{1},\dots,x^{\ell};c)$ and only changes $m_{0}$ to another constant
$m_{0}^{\prime}=m_{0}\mathrm{e}^{2\pi i(t_{1}n_{1}+\dots+t_{\ell}n_{\ell})}.$
In particular, the action by $\sigma$ preserves the degree of $m(x^{1},\dots,x^{\ell};c)$.
On the other hand, the action of any given $\sigma\in W(R)$ on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is linear, hence
$(\sigma m)(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i(\psi\circ\sigma)(x^{1},\dots,x^{\ell};c)}.$
From the relations
$\sigma(\mathbf{0})=\mathbf{0},\quad(-1)\omega+\theta_{1}\alpha_{1}^{\vee}+\dots+\theta_{\ell}\alpha_{\ell}^{\vee}=\mathbf{0},$
it follows that $\sigma$ preserves $(\theta_{1},\dots,\theta_{\ell};-1)$. Therefore by using (2.9) we arrive at
$\deg(\sigma m)=(\psi\circ\sigma)(\theta_{1},\dots,\theta_{\ell};-1)=\psi(\theta_{1},\dots,\theta_{\ell};-1)=\deg m.$
The lemma is proved. ∎
Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. Let
$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
where $N_{j}=\#\{\sigma\in W\mid\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
Proposition 2.2.
The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are homogeneous of degree 00.
Proof.
The invariance of $Y_{j}$ under the action of $W_{a}(R)$ follows from their definition. For any $\sigma\in W_{a}(R)$, we know from Lemma 2.1 that
$\deg\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}=\deg\mathrm{e}^{2\pi i(\omega_{j},\mathbf{x})}=\deg\mathrm{e}^{2\pi i(\theta_{j}c+x^{j})}=-\theta_{j}+\theta_{j}=0,$
thus we arrive at $\deg Y_{j}=0$. The proposition is proved. ∎
Let us proceed to consider generators of the $W_{a}(R)$-invariant ring $\mathscr{F}^{W}$. Since $\mathscr{F}^{W}$ is too large to consider its generators, we are to introduce a smaller subring of $\mathscr{F}$ which is named the $\lambda$-Fourier polynomial ring.
Denote
$\kappa=\gcd\{(\omega,\alpha_{r})~{}|~{}r=1,\dots,\ell\},$ (2.12)
where $\gcd$ refers to the greatest common divisor. This makes sense since
$(\omega,\alpha_{r})=m_{r}(\omega_{r},\alpha_{r})=m_{r}\frac{(\alpha_{r},\alpha_{r})}{2},\quad r=1,\dots,\ell$
are nonnegative rational numbers. Let
$\lambda=\mathrm{e}^{-2\pi i\kappa c},$ (2.13)
then we have
$\deg\lambda=\kappa.$ (2.14)
In particular, if we take $\omega=\omega_{k}$, then $\kappa=(\omega_{k},\alpha_{k})$ and $\lambda=\mathrm{e}^{-2\pi ic(\omega_{k},\alpha_{k})}$. Define the $\lambda$-Fourier polynomial ring $\mathscr{A}$ as a subring of $\mathscr{F}$ by
$\mathscr{A}=\mathbb{C}[\lambda]\otimes\mathbb{C}\bigl{[}\mathrm{e}^{2\pi ix^{j}},\mathrm{e}^{-2\pi ix^{j}}\mid j=1,\dots,\ell\bigr{]}\subset\mathscr{F}.$
Explicitly, $\mathscr{A}$ is composed of elements of the form
$\mathrm{e}^{2\pi i\left(-a_{0}\kappa c+a_{1}x^{1}+\dots+a_{\ell}x^{\ell}\right)},\quad a_{0}\in\mathbb{Z}_{\geq 0},a_{1},\dots,a_{\ell}\in\mathbb{Z}$
and their linear combinations. We denote by $\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$ the $W_{a}(R)$-invariant $\lambda$-Fourier polynomial ring. Then we have $\lambda=\mathrm{e}^{-2\pi i\kappa c}\in\mathscr{A}^{W}$.
Let
$y^{j}(\mathbf{x})=\mathrm{e}^{-2\pi i\theta_{j}c}~{}Y_{j}(\mathbf{x}),~{}j=1,\dots,\ell.$ (2.15)
From Proposition 2.2 it follows that $y^{j}$ is homogeneous of degree $\theta_{j}$.
Lemma 2.3.
We have $y^{j}\in\mathscr{A}^{W}$, $j=1,\dots,\ell$.
Proof.
It is obvious that $y^{j}$ is $W_{a}(R)$-invariant, so we only need to check that $y^{j}\in\mathscr{A}$. It suffices to show that for any $\sigma\in W(R)$, the monomial
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}\in\mathscr{A}.$ (2.16)
Observe that
$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~{}=$ $~{}-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{*}(\omega_{j}),\mathbf{x})$
$=$ $~{}-2\pi ic(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
Since $\sigma^{*}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
$\sigma^{*}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
from which it follows that
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
where $m$ is a non-negative integer. We also have
$\left(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
where $c_{sr}$ are the $(s,r)$-entries of the Cartan matrix of the root system $R$. Therefore, there exists $a_{0}\in\mathbb{Z}_{\geq 0}$ and $a_{1},\dots a_{\ell}\in\mathbb{Z}$ such that
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}=\mathrm{e}^{2\pi i(-a_{0}\kappa c+a_{1}x^{1}+\dots+a_{\ell}x^{\ell})}\in\mathscr{A}.$
The lemma is proved. ∎
Example 2.1 ($(R,\omega)=(A_{2},\omega_{2})$).
We consider the case where the root system $R=A_{2}$ and the weight $\omega=\omega_{2}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}$
of the root system $A_{2}$ in the hyperplane $V$ of the Euclidean space $\mathbb{R}^{3}$ spanned by $\alpha_{1},\alpha_{2}$. We have
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
The affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
$\mathbf{x}=c\omega_{2}+x^{1}\alpha^{\vee}_{1}+x^{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
The Weyl group $W$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
From the definition (2.11) of $Y_{j}$, we have
$Y_{1}$ $=\mathrm{e}^{2\pi i\left(x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x^{2}-x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x^{2}+\frac{2}{3}c\right)},$
$Y_{2}$ $=\mathrm{e}^{2\pi i\left(x^{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x^{1}-x^{2}-\frac{1}{3}c\right)}.$
Here $\kappa=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
Then we have
$y^{1}$ $=e^{-\frac{2}{3}\pi ic}Y_{1}=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})}+\lambda\mathrm{e}^{-2\pi ix^{2}},$
$y^{2}$ $=e^{-\frac{4}{3}\pi ic}Y_{2}=\mathrm{e}^{2\pi ix^{2}}+\lambda\mathrm{e}^{-2\pi ix^{1}}+\lambda\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
In the next subsection we are to show that $y^{1},\dots,y^{\ell}$ and $\lambda$ generate $\mathscr{A}^{W}$.
2.3. The $\varphi_{\lambda}$-transformation and the null transformation
The parameter $\lambda$ of the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ is defined by (2.13), so it lies on the unit circle. However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$.
Definition 2.1.
For any fixed $\lambda\in\mathbb{C}$, the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ yield a transformation
$\varphi_{\lambda}\colon(x^{1},\dots,x^{\ell})\mapsto(y^{1},\dots,y^{\ell}).$
We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null transformation.
Example 2.2 ($(R,\omega)=(A_{2},\omega_{2})$).
From the Example 2.1, we know that the $\varphi_{1}$-transformation of $(y^{1},y^{2})$ is given by
$y^{1}\big{|}_{\lambda=1}$ $=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})}+\mathrm{e}^{-2\pi ix^{2}},$
$y^{2}\big{|}_{\lambda=1}$ $=\mathrm{e}^{2\pi ix_{2}}+\mathrm{e}^{-2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
The corresponding null transformation is given by
$y^{1}\big{|}_{\lambda=0}$ $=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})},$
$y^{2}\big{|}_{\lambda=0}$ $=\mathrm{e}^{2\pi ix^{2}}.$
Note that $y^{2}$ has only one term $\mathrm{e}^{2\pi ix^{2}}$, this result will be generalized in Theorem 2.7.
Proposition 2.4.
The Jacobian determinant of the null transformation is given by
$\mathbf{J}_{0}(\hat{\mathbf{x}})=e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+},\,(\beta,\omega)=0}\left(1-\mathrm{e}^{-2\pi i(\beta,\hat{\mathbf{x}})}\right),$ (2.21)
where $\Phi^{+}$ is the set of all positive roots of $R$.
Proof.
From [6] and [12] we know that the Jacobian determinant of $\varphi_{\lambda}$ can be represented in the form
$\mathbf{J}_{\lambda}(\hat{\mathbf{x}})~{}=$ $~{}\mathrm{e}^{-2\pi ic\sum_{j}(\omega_{j},\omega)+\pi i\sum_{\alpha\in\Phi^{+}}\left(c(\alpha,\omega)+(\alpha,\hat{\mathbf{x}})\right)}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right)$
$=$ $~{}e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right).$ (2.22)
Here we use the fact that [6]
$\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
Since $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is a polynomial of $\lambda$, its constant term w.r.t. $\lambda$ yields (2.21). The proposition is proved. ∎
Corollary 2.5.
The Jacobian determinant of the null transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor. In particular, $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent.
Theorem 2.6.
We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
Proof.
For any given $f\in\mathscr{A}^{W}$, we need to show that it can be represented in terms of a polynomial of $y^{1},\dots,y^{\ell}$ and $\lambda$. By using Lemma 2.1, we know that $f$ can be decomposed as the sum of homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is homogeneous of degree $n$.
From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big{|}_{\lambda=1}$,…,$y^{\ell}\big{|}_{\lambda=1}$ as follows:
$f|_{\lambda=1}=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}(y^{1}\big{|}_{\lambda=1})^{\beta_{1}}\dots(y^{\ell}\big{|}_{\lambda=1})^{\beta_{\ell}},$ (2.24)
so it follows from the homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
$f=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}\lambda^{n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}}(y^{1})^{\beta_{1}}\cdots(y^{\ell})^{\beta_{\ell}},$ (2.25)
here $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
$f=\sum_{j\in Z}\lambda^{j}P_{j}(y^{1},\dots,y^{\ell}),$ (2.26)
where $P_{j}$ are polynomials. Let $q$ be the smallest integer such that $P_{q}=0$. If $q<0$, then by viewing $f$ and $y^{1},\dots,y^{\ell}$ as functions of $x^{1},\dots,x^{\ell}$ and $\lambda$, and by comparing the $\lambda^{q}$ term in (2.26) we arrive at
$P_{q}(y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0})=0.$ (2.27)
This leads to a contradiction with the fact, as shown by Corollary 2.5, that $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent. Thus $q$ must be a non-negative integer, and the proposition is proved. ∎
Definition 2.2.
We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
2.4. Leading terms of the basic generators
From the definition (2.15) of $y^{j}$, we know that it contains a term
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},{\mathbf{x}})}=\mathrm{e}^{2\pi ix^{j}},$ (2.28)
which we call the leading term of $y^{j}$. This term also appears in $y^{j}|_{\lambda=0}$, and the following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ only contain their leading terms.
Theorem 2.7.
Let us denote
$S=\{r\in\{1,\dots,\ell\}\mid m_{r}=(\omega,\alpha_{r}^{\vee})>0\}.$ (2.29)
Then we have
$y^{j}\big{|}_{\lambda=0}=\mathrm{e}^{2\pi ix^{j}},\quad j\in S.$ (2.30)
Proof.
For any given $j\in S$, from (2.11) and (2.15) we obtain
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ (\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (2.31)
Assume that $\sigma\in W$ satisfies the condition
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0,$ (2.32)
then from (2.18) it follows that
$\sum_{r=1}^{\ell}n_{r}m_{r}(\alpha_{r},\omega_{r})=0.$
By using the fact that all terms on the left hand side of the above equality are nonnegative and $m_{j}>0$, we obtain $n_{j}=0$, and
$(\sigma^{*}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
Since $\sigma^{*}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{*}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
The theorem is proved. ∎
Corollary 2.8.
The following identity holds true:
$\sum_{r=1}^{\ell}m_{r}\log\left(y^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.35)
2.5. Proper generators
In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection some other generators of $\mathscr{A}^{W}$.
Definition 2.3.
Suppose $z^{1},\dots,z^{\ell}\in\mathscr{A}^{W}$. We say that $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators of $\mathscr{A}^{W}$ if $\deg z^{j}=\theta_{j}$ and
$z^{j}\big{|}_{\lambda=0}=y^{j}\big{|}_{\lambda=0},\quad j=1,\dots,\ell.$ (2.36)
Lemma 2.9.
Let $z^{1},\dots,z^{\ell}\in\mathscr{A}^{W}$, then these functions form a set of proper generators of $\mathscr{A}^{W}$ if and only if for any $1\leq j\leq\ell$ there exists a polynomial $s^{j}$ of $\lambda$ and $\{y^{r}\big{|}~{}1\leq r\leq\ell,~{}\theta_{r}<\theta_{j}\}$ such that either $s^{j}=0$ or $\deg s^{j}=\theta_{j}-\kappa$, and
$z^{j}=y^{j}+\lambda s^{j}.$ (2.37)
Proof.
If $z^{1},\dots,z^{\ell}$ satisfy the relation (2.37), then it is obvious to see that they form a set of proper generators of $\mathscr{A}^{W}$.
Now we suppose that $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators of $\mathscr{A}^{W}$, then for any $1\leq j\leq\ell$ we have $(z^{j}-y^{j})\in\mathscr{A}^{W}$ and
$\left(z^{j}-y^{j}\right)\big{|}_{\lambda=0}=0.$
Hence there exists $s^{j}\in\mathscr{A}^{W}$ such that $z^{j}-y^{j}=\lambda s^{j}$, and either $s^{j}=0$ or $\deg s^{j}=\theta_{j}-\kappa$. By using Theorem 2.6 and by degree counting, we know that $s^{j}$ is a polynomial of $\lambda$ and $y^{r}$’s whose degrees are less than $\theta_{j}$. The lemma is proved. ∎
Proposition 2.10.
Let $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$, then
$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell},\lambda].$ (2.38)
Proof.
It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for $j=1,\dots,\ell$. Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, and let $y^{k}$ be one of those with the lowest degree. It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$. However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}\not\in C[z^{1},\dots,z^{\ell},\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$. Then we know from Lemma 2.9 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, which leads to a contradiction. The proposition is proved. ∎
Corollary 2.11.
For any set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we have
$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}_{\lambda}(\hat{\mathbf{x}}),$ (2.39)
where $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is defined in (2.22), thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
Proof.
For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ satisfying the condition $\theta_{r}<\theta_{j}$. Then from the relation
$\frac{\partial z^{j}}{\partial x^{p}}=\frac{\partial y^{j}}{\partial x^{p}}+\lambda\sum_{\theta_{r}<\theta_{j}}\frac{\partial s^{j}}{\partial y^{r}}\frac{\partial y^{r}}{\partial x^{p}}$ (2.40)
we arrive at (2.39). The corollary is proved. ∎
For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we will also call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null transformation. Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation, i.e.,
$z^{j}\big{|}_{\lambda=0}=y^{j}\big{|}_{\lambda=0},\quad j=1,\dots,\ell.$ (2.41)
In particular, for any element $r$ of the set $S$ defined by (2.29) we have
$z^{r}\big{|}_{\lambda=0}=\mathrm{e}^{2\pi ix^{r}}.$ (2.42)
Hence from Corollary 2.35 it follows that
$\sum_{r=1}^{\ell}m_{r}\log\left(z^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.43)
In the next section, we will impose an additional condition on proper generators of $\mathscr{A}^{W}$ in order to construct generalized Frobenius manifold structures on the orbits space of the affine Weyl group $W_{a}(R)$.
3. The construction of generalized Frobenius manifold structures
In this section, we are to construct a generalized Frobenius manifold structure starting from a given irreducible reduced root system $R$ in the $l$-dimensional Euclidean space $V$, a fixed weight $\omega$ of the form (2.1), and a set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$ satisfying a certain additional condition.
3.1. The orbit space and pencil generators
By definition, $\mathbb{C}[z^{1},\cdots,z^{\ell}]$ is a subalgebra of $\mathscr{A}^{W}$. Denote the spectrum of this subalgebra by $\mathcal{M}$, then $\mathcal{M}\cong\mathbb{C}^{\ell}$. Taking $x^{1},\cdots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
$\varphi_{\lambda}\colon V\otimes\mathbb{C}\to\mathcal{M},\quad(x^{1},\dots,x^{\ell})\mapsto(z^{1},\dots,z^{\ell}),\quad\lambda\in\mathbb{C}.$ (3.1)
Here the $\varphi_{\lambda}$-transformation is defined in the last section. We call $\mathcal{M}$ the orbit space of the affine Weyl group $W_{a}(R)$ w.r.t. the proper generators $z^{1},\dots,z^{\ell}$. In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to endow $\mathcal{M}$ with a structure of generalized Frobenius manifold.
[…]
$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\ \mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
on $V\otimes\mathbb{C}$. Here by a contravariant metric we mean a symmetric non-degenerate bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$.
Proposition 3.1.
The numbers $\theta_{j}$ defined by (2.4) can be represented in the form
$\theta_{j}=a^{jr}m_{r},\quad j=1,\dots,\ell.$ (3.3)
Here and in what follows, we assume summation over repeated upper and lower indices.
Proof.
From $\omega=\sum_{s}m_{s}\omega_{s}=\sum_{p}\theta_{p}\alpha_{p}^{\vee}$ it follows that
$\sum_{p=1}^{\ell}\theta_{p}(\alpha_{p}^{\vee},\alpha_{s}^{\vee})=(\omega,\alpha_{s}^{\vee})=m_{s},$ (3.4)
thus (3.3) holds true. The proposition is proved. ∎
The $\varphi_{\lambda}$-transformation induces a collection of pushing forward $\{(\varphi_{\lambda})_{*}\}_{\lambda\in\mathbb{C}}$, which transforms $(a^{ij})$ to a contravariant metric $(\varphi_{\lambda})_{*}a$ on $\mathcal{M}$. We denote
$g_{\lambda}=\frac{1}{4\pi^{2}}(\varphi_{\lambda})_{*}a=\bigl{(}g_{\lambda}^{ij}\bigr{)},$
where
$g_{\lambda}^{ij}=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{r}}a^{rs}\frac{\partial z^{j}}{\partial x^{s}},\quad i,j=1,\dots,\ell.$ (3.5)
Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s.
Proposition 3.2.
We have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$.
Proof.
By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ is invariant w.r.t. the action of $W_{a}(R)$. Since the invariance of $g^{ij}_{\lambda}$ w.r.t. translations is obvious, we consider actions by elements of the Weyl group $W(R)$. Write $z^{j}$ as $z^{j}=z^{j}(x^{1},\dots,x^{\ell};c)$. For any given $\sigma\in W(R)$, since $z^{j}(x^{1},\dots,x^{\ell};c)=z^{j}(\sigma(x^{1},\dots,x^{\ell};c))$, we have
$\frac{\partial z^{j}}{\partial x^{r}}(x^{1},\dots,x^{\ell};c)=\frac{\partial z^{j}}{\partial x^{p}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial\tilde{x}^{p}}{\partial x^{r}},$ (3.6)
where $\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c)$. Note that $\frac{\partial\tilde{x}^{p}}{\partial x^{r}}$ are constants since $\sigma$ is a linear transformation. Therefore,
$g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{r}}(x^{1},\dots,x^{\ell};c)a^{rs}\frac{\partial z^{j}}{\partial x^{s}}(x^{1},\dots,x^{\ell};c)$
$\quad=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{p}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial z^{j}}{\partial x^{q}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}.$
Since $\sigma$ is an orthogonal transformation, we know that
$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq}.$
Hence $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$. The proposition is proved. ∎
The above proposition shows in particular that $g^{ij}_{\lambda}$ are polynomials of $\lambda$. In order to construct generalized Frobenius manifold structures on the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.
Definition 3.1.
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$. We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl{(}g^{ij}_{\lambda}\bigr{)}$ depends linearly on $\lambda$, i.e., $g^{ij}_{\lambda}$ can be represented in the form
$g^{ij}_{\lambda}=g^{ij}+\lambda\eta^{ij},$
where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the contravariant components $\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j}$ of the Levi-Civita connection of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
$\Gamma^{ij}_{g_{\lambda},k}=\Gamma^{ij}_{g,k}+\lambda\Gamma^{ij}_{\eta,k}.$ (3.7)
From the above definition we know that a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil[10] of metrics $g,\eta$. Note that the flat metric $g$ is independent of the choice of proper generators, since
$g=g_{0}=\frac{1}{4\pi^{2}}(\varphi_{0})_{*}a$ (3.8)
is induced merely by the null-transformation which does not depend on the choice of proper generators.
We will present examples of pencil generators in Sect. 5 and in a subsequent paper [20]. Unless otherwise specified, in what follows we always assume that $\{z^{1},\dots,z^{\ell}\}$ is a set of pencil generators of $\mathscr{A}^{W}$. We will end this section by proving that we can construct a generalized Frobenius manifold structure on $\mathcal{M}$ with the given flat pencil of metrics.
3.2. Homogeneous flat coordinates
We are to show in this subsection the existence of a system of homogeneous flat coordinates for the flat metric $\eta$.
Flat coordinates of $\eta$ satisfy the system of equations
$\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\quad i,j=1,\dots,\ell.$ (3.9)
Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of $\eta$’s Levi-Cività connection, they are given by
$\Gamma_{\eta,ij}^{k}=\frac{1}{2}\eta^{kr}\left(\frac{\partial\eta_{ir}}{\partial z^{j}}+\frac{\partial\eta_{jr}}{\partial z^{i}}-\frac{\partial\eta_{ij}}{\partial z^{r}}\right).$ (3.10)
From Definition 2.3 we know that
$\deg g^{ij}=\theta_{i}+\theta_{j},\quad\deg\eta^{ij}=\theta_{i}+\theta_{j}-\kappa,$ (3.11)
which lead to
$\deg g_{ij}=-\theta_{i}-\theta_{j},\quad\deg\eta_{ij}=\kappa-\theta_{i}-\theta_{j},\quad\deg\Gamma_{\eta,ij}^{k}=\theta_{k}-\theta_{i}-\theta_{j}.$ (3.12)
We are to use the above homogeneity property to prove the existence of homogeneous flat coordinates of $\eta$. To this end, let us first clarify the solution space of the system of equations (3.9) that we are considering. Let $\widehat{\mathcal{M}}$ be the universal covering of $\mathcal{M}\setminus D$ with
$D=\{z\in\mathcal{M}~{}\big{|}~{}\det\eta(z)=0\}.$ (3.13)
We consider solutions of the system of equations (3.9) on the space
$\mathscr{Q}(\widehat{\mathcal{M}})=\mathscr{O}(\widehat{\mathcal{M}})/\mathbb{C},$
where $\mathscr{O}(\widehat{\mathcal{M}})$ refers to the holomorphic function ring on $\mathcal{M}$. Denote the solution space of (3.9) by
$\mathcal{K}=\left\{t\in\mathscr{Q}(\widehat{\mathcal{M}})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
It follows from the theory of differential equations that $\mathcal{K}$ is an $\ell$-dimensional space, since a solution is uniquely determined by the values of its $\ell$ partial derivatives at an arbitrary point.
Now we introduce the Euler vector field
$E=\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial}{\partial z^{r}}.$ (3.14)
Then for any homogeneous function $f$ of degree $n$, the following identity holds true:
$\mathcal{L}_{E}f=nf,$ (3.15)
here $\mathcal{L}_{E}$ is the Lie derivative along $E$. Hence homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees. Now we consider the action of $\mathcal{L}_{E}$ on the solution space $\mathcal{K}$.
Proposition 3.3.
The space $\mathcal{K}$ is $\mathcal{L}_{E}$-invariant, i.e., $\mathcal{L}_{E}(\mathcal{K})\subseteq\mathcal{K}$.
Proof.
Denote
$\mathcal{L}_{ij}=\frac{\partial^{2}}{\partial z^{i}\partial z^{j}}-\Gamma_{\eta,ij}^{k}\frac{\partial}{\partial z^{k}}.$ (3.16)
By using (3.12) we obtain
$\mathcal{L}_{E}\Gamma_{\eta,ij}^{k}=(\theta_{k}-\theta_{i}-\theta_{j})\Gamma_{\eta,ij}^{k},$
from which it follows that
$\left[\mathcal{L}_{E},\mathcal{L}_{ij}\right]=-(\theta_{i}+\theta_{j})\mathcal{L}_{ij}.$ (3.17)
Hence if $t\in\ker\mathcal{L}_{ij}$, then $\mathcal{L}_{E}t\in\ker\mathcal{L}_{ij}$. The proposition is proved. ∎
In what follows, we always assume that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable. This implies that $\eta$ admits $\ell$ homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$. Let
$d_{\alpha}=\deg t^{\alpha},\quad\alpha=1,\dots,\ell,$ (3.18)
then we have
$E=\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial t^{\alpha}}{\partial z^{r}}\frac{\partial}{\partial t^{\alpha}}=\sum_{\alpha=1}^{\ell}d_{\alpha}t^{\alpha}\frac{\partial}{\partial t^{\alpha}}.$ (3.19)
Now let us consider the flat pencil of metrics $g$ and $\eta$ in the flat coordinates $t^{1},\dots,t^{\ell}$. We denote their components by $g^{\alpha\beta}$ and $\eta^{\alpha\beta}$ with Greek indices respectively. Recall that we use $g^{ij}$ and $\eta^{ij}$ with Latin indices to denote the components of $g$ and $\eta$ in the coordinates $z^{1},\dots,z^{\ell}$. Since $\eta^{\alpha\beta}$ are constants, from
$\deg\eta^{\alpha\beta}=d_{\alpha}+d_{\beta}-\kappa$ (3.20)
it follows that if $d_{\alpha}+d_{\beta}\not=\kappa$ then $\eta^{\alpha\beta}$ must vanish. Thus by performing, if necessary, a degree preserving linear transformation, we can choose the flat coordinates $t^{1},\dots,t^{n}$ of $\eta$ such that
$\eta^{\alpha\beta}=\delta_{\beta,\alpha^{*}}.$ (3.21)
Here * is an involution
:$*\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
which satisfies the relation $d_{\alpha}+d_{\alpha^{*}}=\kappa$. Usually, we prefer to choose the flat coordinates so that $\alpha^{*}=\ell+1-\alpha$.
We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$. Since
$\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}E=\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}\sum_{k=1}^{\ell}\theta_{k}z^{k}\frac{\partial}{\partial z^{k}}=\left(\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,ik}^{j}+\theta_{i}\delta^{j}_{i}\right)\frac{\partial}{\partial z^{j}},$ (3.22)
the matrix $A$ has elements
$A_{j}^{i}=\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,jk}^{i}+\theta_{j}\delta^{i}_{j}.$ (3.23)
It has eigenvectors $\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T}$ with eigenvalues $d_{\alpha}$, i.e.,
$A\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}=d_{\alpha}\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}.$ (3.24)
Remark 3.1.
When we choose $\omega$ to be a fundamental weight, we conjecture that all $d_{\alpha}$ are positive numbers. In what follows, we will continue our construction by assuming this fact.
3.3. The Frobenius algebra
Denote by $\Gamma_{\alpha\gamma}^{\beta}$ the Christoffel symbols of the Levi-Civita connection of $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$, and by
$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta}$ (3.25)
the contravariant components of the Levi-Civita connection of $g$. Then from (3.7) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the contravariant components of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$.
We define an algebra structure on $T^{*}\mathcal{M}$ by
$\mathrm{d}t^{\alpha}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\Gamma_{\gamma}^{\alpha\beta}\mathrm{d}t^{\gamma},\quad\alpha,\beta=1,\dots,\ell,$ (3.26)
[…]
$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}.$ (3.27)
Proposition 3.4.
The 1-form (3.27) is a left unity of the algebra structure defined by (3.26) on $T^{*}\mathcal{M}$.
Proof.
From the definition of the multiplication (3.26) it follows that
$\omega_{e}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}~{}\nabla_{g}^{\omega_{e}}\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}.$ (3.28)
Here $\nabla_{g}$ (resp. $\nabla^{g}$) is the contravariant (resp. covariant dual) Levi-Civita connection, and $g^{\sharp}$ is the canonical isomorphism
$g^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto g(\omega,\cdot).$ (3.29)
So it suffices to show that
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}=\mathrm{d}t^{\beta}.$ (3.30)
Note that the null-transformation $\varphi_{0}$ transforms the metric $a$ defined in (3.2) to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$. Since the relation between the coordinates $z^{1},\dots,z^{\ell}$ and $t^{1},\dots,t^{\ell}$ is independent of $\lambda$, in the following proof of the validity of (3.30) we can assume that $z^{i}$ and $t^{\alpha}$ are related with $x^{1},\dots,x^{\ell}$ via the null-transformation $\varphi_{0}$. From (2.43) we obtain
$\varphi_{0}^{*}\omega_{e}=-\frac{2}{\kappa}\pi i\mathrm{d}(\omega,\hat{\mathbf{x}}),$ (3.31)
it follows that
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}$ $=-2\pi i\frac{1}{d_{\beta}}\sum_{s=1}^{\ell}m_{s}\nabla^{g}_{g^{\sharp}\mathrm{d}x^{s}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\mathrm{d}x^{p}\right)$
$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}(\mathrm{d}x^{r},\mathrm{d}x^{s})\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
$=\frac{1}{d_{\beta}}\left(\sum_{r=1}^{\ell}\frac{1}{2\pi i}\theta_{r}\frac{\partial}{\partial x^{r}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\right)\mathrm{d}x^{p}.$
Here we used the relation (3.3) for the third equality. Note that $t^{\beta}\bigr{|}_{\lambda=0}$ is a Fourier polynomial of degree $d_{\beta}$ which does not depend on $\lambda$, so $\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}$ and
$\sum_{r=1}^{\ell}\frac{1}{2\pi i}~{}\theta_{r}\frac{\partial}{\partial x^{r}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)=\left.d_{\beta}\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}.$ (3.32)
Thus we arrive at
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}=\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\mathrm{d}x^{p}=\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}.$ (3.33)
The proposition is proved. ∎
Corollary 3.5.
Denote $\omega_{e}=e_{\alpha}\mathrm{d}t^{\alpha}$, then the following relations hold true:
$e_{\alpha}g^{\alpha\beta}=\frac{1}{\kappa}d_{\beta}t^{\beta},\quad e_{\alpha}\Gamma_{\gamma}^{\alpha\beta}=\frac{1}{\kappa}d_{\beta}\delta_{\gamma}^{\beta}.$ (3.34)
Proof.
The second relation of (3.34) follows immediately from Proposition 3.4. To prove the first relation of (3.34) we can assume, as in the proof of Proposition 3.4, that the coordinates $t^{1},\dots,t^{\ell}$ and $x^{1},\dots,x^{\ell}$ are related via the null-transformation $\varphi_{0}$. Thus we have
$e_{\alpha}g^{\alpha\beta}$ $=g(\omega_{e},\mathrm{d}t^{\beta})=-\frac{2}{\kappa}\pi ig(\mathrm{d}(\omega,\hat{\mathbf{x}}),\mathrm{d}t^{\beta})$
$=\frac{\kappa}{2\pi i}\sum_{s=1}^{\ell}m_{s}(\mathrm{d}x^{r},\mathrm{d}x^{s})\frac{\partial t^{\beta}}{\partial x^{r}}=\frac{1}{\kappa}\sum_{s=1}^{\ell}\frac{\theta_{s}}{2\pi i}\frac{\partial t^{\beta}}{\partial x^{s}}=d_{\beta}t^{\beta}.$
The corollary is proved. ∎
Proposition 3.6 ([9], Appendix D).
The Christoffel symbols $\Gamma_{\gamma}^{\alpha\beta}$ are uniquely determined by the metric $g$ via the relations
$\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}=\Gamma_{\gamma}^{\alpha\beta}+\Gamma_{\gamma}^{\beta\alpha},\quad g^{\alpha\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=g^{\beta\epsilon}\Gamma_{\epsilon}^{\alpha\gamma}.$ (3.35)
Moreover, we have the following properties for the flat pencil of metrics:
$\eta^{\alpha\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=\eta^{\beta\epsilon}\Gamma_{\epsilon}^{\alpha\gamma},\quad\frac{\partial\Gamma_{\gamma}^{\alpha\beta}}{\partial t^{\delta}}=\frac{\partial\Gamma_{\delta}^{\alpha\beta}}{\partial t^{\gamma}},\quad\Gamma_{\epsilon}^{\alpha\beta}\Gamma^{\epsilon\gamma}_{\delta}=\Gamma_{\epsilon}^{\alpha\gamma}\Gamma^{\epsilon\beta}_{\delta}.$ (3.36)
It follows from the first two relations of (3.36) that there exist functions $f^{\beta}$ such that
$\eta^{\alpha\epsilon}\frac{\partial^{2}f^{\beta}}{\partial t^{\epsilon}\partial t^{\gamma}}=\Gamma^{\alpha\beta}_{\gamma},$ (3.37)
where $\deg f^{\beta}=\kappa+d_{\beta}$
Lemma 3.7.
The Christoffel symbols satisfy the following relations:
$\Gamma_{\alpha}^{\beta\gamma}=\frac{d_{\gamma}}{d_{\beta}+d_{\gamma}}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}}.$ (3.38)
Proof.
By multiplying $e_{\alpha}$ on both sides of the second relation of (3.35) and by using (3.34) we obtain
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=d_{\gamma}g^{\beta\gamma},$ (3.39)
so it follows from (3.37) that
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}\right)=(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}=d_{\gamma}g^{\beta\gamma}.$ (3.40)
Thus we have
$(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial^{2}f^{\gamma}}{\partial t^{\nu}\partial t^{\alpha}}=d_{\gamma}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}},$ (3.41)
which leads to the validity of (3.38). The lemma is proved. ∎
Theorem 3.8.
[…]
Proof.
We need to show that the algebra is commutative and associative, and the operation of multiplication is compatible with the bilinear form. From (3.38) it follows that
$\frac{1}{d_{\gamma}}\Gamma^{\beta\gamma}_{\alpha}=\frac{1}{d_{\beta}}\Gamma^{\gamma\beta}_{\alpha},$ (3.42)
which leads to the commutativity of the algebra. Now by using the third relation of (3.36) we have
$\left(\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta}\right)\cdot\operatorname{d\!}t^{\gamma}=\frac{\kappa}{d_{\beta}}\Gamma^{\alpha\beta}_{\xi}\operatorname{d\!}t^{\xi}\cdot\operatorname{d\!}t^{\gamma}=\frac{\kappa}{d_{\beta}}\frac{\kappa}{d_{\gamma}}\Gamma^{\alpha\beta}_{\xi}\Gamma^{\xi\gamma}_{\zeta}\operatorname{d\!}t^{\zeta}$
$\quad=\frac{\kappa}{d_{\alpha}}\frac{\kappa}{d_{\gamma}}\Gamma^{\beta\alpha}_{\xi}\Gamma^{\xi\gamma}_{\zeta}\operatorname{d\!}t^{\zeta}=\frac{\kappa}{d_{\alpha}}\frac{\kappa}{d_{\gamma}}\Gamma^{\beta\gamma}_{\xi}\Gamma^{\xi\alpha}_{\zeta}\operatorname{d\!}t^{\zeta}$
$\quad=\left(\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\right)\cdot\operatorname{d\!}t^{\alpha}=\operatorname{d\!}t^{\alpha}\cdot\left(\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\right),$
so the associativity of the algebra holds true. Finally, from the first relation of (3.36) we obtain
$\langle\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta},\operatorname{d\!}t^{\gamma}\rangle=\frac{\kappa}{d_{\beta}}\eta^{\gamma\xi}\Gamma^{\alpha\beta}_{\xi}=\frac{\kappa}{d_{\beta}}\eta^{\alpha\xi}\Gamma^{\gamma\beta}_{\xi}=\frac{\kappa}{d_{\gamma}}\eta^{\alpha\xi}\Gamma^{\beta\gamma}_{\xi}=\langle\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\rangle,$
which verifies the compatibility of the multiplication with the bilinear form. The proposition is proved. ∎
The Frobenius algebra on $T^{*}\mathcal{M}$ induces a Frobenius algebra structure on $T(M\setminus D)$ via the isomorphism
$\eta^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.43)
where $D$ is define by (3.13). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T(M\setminus D)$ is defined by
$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}),$ (3.44)
with the bilinear form
$\langle X,Y\rangle=\eta^{-1}(X,Y)=\eta(\eta^{\flat}(X),\eta^{\flat}(Y))$ (3.45)
and the unity
$e=\eta^{\sharp}(\omega_{e}),$ (3.46)
here $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.14), (3.20) and (3.27) it follows that
$\mathcal{L}_{E}e=-\kappa e.$
3.4. The generalized Frobenius manifold structure on $\mathcal{M}\setminus D$
The notion of a generalized Frobenius manifold is a certain generalization of Dubrovin’s definition of Frobenius manifolds [9], it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
Definition 3.2 ([9, 22]).
[…]
  1. FM1.
[…]
  2. FM2.
    Define a $3$-tensor $c$ by $c(u,v,w)=\langle u\cdot v,w\rangle$, then the $4$-tensor
    $\nabla c:(W,X,Y,Z)\to(\nabla_{W}c)(X,Y,Z)$ (3.47)
    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$.
  3. FM3.
    There exist an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y.$ (3.48)
    $\mathcal{L}_{E}\langle X,Y\rangle$ $=\langle\mathcal{L}_{E}X,Y\rangle+\langle X,\mathcal{L}_{E}Y\rangle+(2-d)\langle X,Y\rangle.$ (3.49)
    Here $\nabla$ is the Levi-Civita connection of the flat metric, and $d$ is a constant whic is called the charge of $M$.
Let us show that the Frobenius algebra on $T(\mathcal{M}\setminus D)$ that is defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}\setminus D$ with the Euler vector field
$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.50)
and with charge $d=1$. The validity of the axiom FM1 follows from the flatness of the metric $\eta$, and that of the axiom FM2 follows from the relation (3.42) and the second relation of (3.36). To check the quasi-homogeneity of the Frobenius manifold structure required by the axiom FM3, we need to verify equations (3.48) and (3.49) which can be represented in the flat coordinates $t^{1},\dots,t^{\ell}$ as follows:
$\partial_{\tilde{E}}c_{\alpha\beta}^{\gamma}$ $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}+\frac{d_{\gamma}}{\kappa}\right)c_{\alpha\beta}^{\gamma},$ (3.51)
00 0 $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.52)
Here
$c_{\alpha\beta}^{\gamma}=\frac{\kappa}{d_{\rho}}\eta_{\alpha\nu}\eta_{\beta\rho}\eta^{\gamma\zeta}\Gamma^{\nu\rho}_{\zeta}$ (3.53)
are the structure constants of the Frobenius algebra
$\frac{\partial}{\partial t^{\alpha}}\cdot\frac{\partial}{\partial t^{\beta}}=c_{\alpha\beta}^{\gamma}\frac{\partial}{\partial t^{\gamma}},$ (3.54)
and $(\eta_{\alpha\beta})=(\eta^{\alpha\beta})^{-1}$ with $\eta^{\alpha\beta}=\eta(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})$. The validity of the identities follows from (3.20) and the fact that $\deg\Gamma^{\alpha\beta}_{\gamma}=d_{\alpha}+d_{\beta}-d_{\gamma}$. Thus we have the following theorem.
Theorem 3.9 (Main).
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}\setminus D$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.44). Moreover, the unit vector field $e$ and the Euler vector field $\tilde{E}$ of this generalized Frobenius are given by (3.46) and (3.50) respectively, and the intersection form coincides with $g$.
Proof.
We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}\setminus D$ coincides with $g$, i.e.,
$g(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})=i_{\tilde{E}}(\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta}).$ (3.55)
Indeed, by using Lemma 3.7 we obtain
$i_{\tilde{E}}(\mathrm{d}t^{\alpha}\cdot\mathrm{d}t^{\beta})=\frac{\kappa\tilde{E}^{\gamma}}{d_{\beta}}\Gamma^{\alpha\beta}_{\gamma}=\sum_{\gamma}\frac{d_{\gamma}t^{\gamma}}{d_{\beta}}\Gamma^{\alpha\beta}_{\gamma}=\sum_{\gamma}\frac{d_{\gamma}t^{\gamma}}{d_{\alpha}+d_{\beta}}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}=g^{\alpha\beta}.$ (3.56)
Here we used the fact that $\deg g^{\alpha\beta}=d_{\alpha}+d_{\beta}$. The theorem is proved. ∎
Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widehat{\mathcal{M}\setminus D}$ of $\mathcal{M}\setminus D$, hence $\pi_{1}(\mathcal{M}\setminus D)$ acts on those functions. Denote the stable subgroup of this action by
$H=\{\gamma\in\pi_{1}(\mathcal{M}\setminus D)\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),~{}\forall p\in\widehat{\mathcal{M}\setminus D}\},$ (3.57)
then $H$ corresponds to a covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$. Concretely, we may define an equivalence relation $\sim$ on $\widehat{\mathcal{M}\setminus D}$ by
$p\sim p^{\prime}\iff\pi(p)=\pi(p^{\prime}),\quad t^{\alpha}(p)=t^{\alpha}(p^{\prime}),\quad\alpha=1,\dots,\ell.$ (3.58)
Here $\pi\colon\widehat{\mathcal{M}\setminus D}\to\mathcal{M}\setminus D$ is the canonical projection. Then we have
¯$\overline{\mathcal{M}}=\widehat{\mathcal{M}\setminus D}/\sim.$ (3.59)
We are to show in Corollary 4.3 that $t^{1},\dots,t^{\ell}$ are coordinates of $\overline{\mathcal{M}}$. Therefore, the generalized Frobenius manifold structure that we constructed above is defined on $\overline{\mathcal{M}}$ instead of on $\mathcal{M}\setminus D$. If there is no ambiguity, we usually identify $\overline{\mathcal{M}}$ with $\mathcal{M}\setminus D$ since they are locally isomorphic. However, when we discuss the monodromy group of the generalized Frobenius manifold structure in the next section, it is necessary to view this generalized Frobenius manifold as $\overline{\mathcal{M}}$. We will prove in the next section that $z^{1},\dots,z^{\ell}$ are polynomials of $t^{1},\dots,t^{\ell}$, so we have the canonical projection
$\overline{\pi}:~{}\overline{\mathcal{M}}\to\mathcal{M}\setminus D.$
Since $\overline{\mathcal{M}}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we will also denote this generalized Frobenius manifold by
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$ (3.60)
If there is no confusion on the selection of the pencil generators, we simply write it as
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega).$ (3.61)
4. Polynomial Property and the Monodromy Groups
In this section, we consider the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Note that the monodromy groups of the Frobenius manifold structures that are constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups coincide with the underlined groups. In the present case the monodromy group $\mathrm{Mono}(\overline{\mathcal{M}})$ can be embedded into the affine Weyl group $W_{a}(R)$. More precisely, we will show that the image of this embedding satisfies
$\mathrm{Mono}(\overline{\mathcal{M}})~{}\leq~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell},$ (4.1)
where
$\mathrm{Stab}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}.$ (4.2)
To this end, we first need to prove Theorem 4.2 which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
4.1. Periods of the generalized Frobenius manifold
Let us consider flat coordinates of the flat metric $g+\lambda\eta$ of $\overline{\mathcal{M}}$. They satisfy the equations
$\nabla_{g+\lambda\eta}\mathrm{d}u(\lambda)=0$ (4.3)
and are called periods of the generalized Frobenius manifold [9, 22]. Denote
$M_{\lambda}=(\overline{\mathcal{M}}\times\mathbb{C}^{\times})\setminus{\Sigma_{\lambda}},$ (4.4)
where
${\Sigma}_{\lambda}=\{(z,\lambda)\in M\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
Then solutions of the equation (4.3) live on
$\mathscr{Q}(\widehat{M_{\lambda}})=\mathscr{O}(\widehat{M_{\lambda}})/\mathbb{C}.$
In the coordinates $(z^{1},\dots,z^{\ell})$, the equation (4.3) reads
$\frac{\partial^{2}u}{\partial z^{i}\partial z^{j}}-\Gamma_{\lambda,ij}^{k}\frac{\partial u}{\partial z^{k}}=0,$ (4.5)
Here $\Gamma_{\lambda,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$. Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$. By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ such that
$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}\right).$ (4.6)
Therefore, the refined flat coordinates
$\check{x}^{r}(z,\lambda):=x^{r}+\theta_{r}c=x^{r}+\frac{\theta_{r}}{\kappa}\log\lambda,\quad r=1,\dots,\ell$ (4.7)
depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are homogeneous of degree zero, i.e., they satisfy the equation
$\kappa\lambda\frac{\partial u}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial u}{\partial z^{j}}=0.$ (4.8)
We denote by $\mathcal{S}_{\lambda}$ the space of periods of $\overline{\mathcal{M}}$ which are homogeneous of degree zero, then $\dim\mathcal{S}_{\lambda}=\ell$, and
$\mathcal{S}_{\lambda}=\mathrm{span}_{\mathbb{C}}\left\{\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)\right\}.$ (4.9)
Now we proceed to find another basis of $\mathcal{S}_{\lambda}$ that is associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
Proposition 4.1.
For each homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a solution
$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ (4.10)
of the equations (4.5) and (4.8), where $h^{\alpha}(z,\rho)$ can be analytically continued to $\rho=0$ such that $h^{\alpha}(z,0)=t^{\alpha}(z)$.
Proof.
Recall that we made the assumption in Remark 3.1 that $d_{\alpha}>0$, so from (3.20) we have $0<d_{\alpha}<\kappa$. Let us looking for a solution of (4.5) and (4.8) of the form $u(z,\rho)=\rho^{d_{\alpha}/\kappa}h(z,\rho)$ satisfying the initial condition $h(z,0)=t^{\alpha}(z)$. Denote $\xi_{j}=\frac{\partial h}{\partial z^{j}}$, then these equations take the form
$\frac{\partial\xi_{j}}{\partial z^{i}}-\Gamma_{ij}^{k}(\rho)\xi_{k}$ $=0,$ (4.11)
$-\kappa\rho\frac{\partial\xi_{j}}{\partial\rho}+\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial\xi_{j}}{\partial z^{r}}$ $=(d_{\alpha}-\theta_{j})\xi_{j},$ (4.12)
where $\Gamma_{ij}^{k}(\rho)=\Gamma_{1/\rho,ij}^{k}$, which are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$. By using the equations (4.11) we can rewrite the equations (4.12) as
$\kappa\rho\frac{\partial\vec{\xi}}{\partial\rho}=\left(A(\rho)-d_{\alpha}\right)\vec{\xi},$ (4.13)
where
$\vec{\xi}=\vec{\xi}(\rho)=(\xi_{1},\cdots,\xi_{\ell})^{T},\quad A_{j}^{k}(\rho)=\sum_{r=1}^{\ell}\theta_{r}z^{r}\Gamma_{rj}^{k}(\rho)+\theta_{j}\delta_{j}^{k}.$
Expand $A(\rho)$ and $\zeta{\xi}(\rho)$ into power series of $\rho$ as follows:
$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{0}+\rho\vec{\xi}_{1}+\rho^{2}\vec{\xi}_{2}+\dots,$
here $A=A(0)=(A^{k}_{j}(0))$ coincides with the matrix $A$ defined in (3.23), and
$\vec{\xi}_{0}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T},$ (4.14)
then the equation (4.13) is equivalent to
$(A-d_{\alpha})\vec{\xi}_{0}=\mathbf{0},\quad(A-d_{\alpha}-\kappa)\vec{\xi}_{1}=-A_{1}\vec{\xi}_{0},$ (4.15)
$(A-d_{\alpha}-2\kappa)\vec{\xi}_{2}=-A_{2}\vec{\xi}_{0}-A_{1}\vec{\xi}_{1},\dots$ (4.16)
From (3.24) we know that the first equation of (4.15) is satisfied by $\vec{\xi}_{0}$ that is defined by (4.14). Since the eigenvalues of $A$ are positive and less than $\kappa$, we can solve the other linear equations of (4.15) and (4.16) to obtain a unique solution $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ recursively. From the fact that $\vec{\xi}_{0}$ gives a solution of the equations (4.11), and from the compatibility of of the equations of (4.11) and (4.12) we know that $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition. ∎
4.2. The polynomiality property
The periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\overline{\mathcal{M}}$ given by Proposition 4.1 are linearly independent, so they yield another basis of $\mathcal{S}_{\lambda}$. Thus there exist constants $c_{\alpha}^{\beta}$ and $s^{\beta}$ such that
$\check{x}^{\beta}(z,\rho)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
Since the functions $Z_{r}=Z_{r}(\mathrm{e}^{2\pi i\check{x}^{1}},\dots,\mathrm{e}^{2\pi i\check{x}^{\ell}})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$. Hence we have
$z^{r}=\rho^{-\theta_{r}/\kappa}Z_{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}(v^{1}(z,\rho))^{n_{1}}\cdots(v^{\ell}(z,\rho))^{n_{\ell}},$ (4.18)
where $a_{n_{1},\dots,n_{\ell}}$ are some constants.
Theorem 4.2.
For each $r=1,\cdots,\ell$, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$.
Proof.
It follows from Proposition 4.1 that
$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\rho\to 0.$
Thus we have
$z^{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}\left(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa})\right).$ (4.19)
Let us denote
$\theta=\inf\{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}~{}\big{|}~{}a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.20)
Suppose $\theta>\theta_{r}$, then it follows from (4.19) that
$z^{r}=\rho^{-\theta_{r}/\kappa}o(\rho^{\theta_{r}/\kappa})=o(1),\quad\rho\to 0,$ (4.21)
this is not true due to our definition of proper generators. On the other hand, if $\theta<\theta_{r}$, then from (4.19) we know that
$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\rho\to 0.$ (4.22)
Since $(\theta-\theta_{r})/\kappa<0$, we must have
$\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}=0.$
However, since $t^{1},\cdots,t^{\ell}$ are functionally independent, all those coefficients $a_{n_{1},\cdots,n_{\ell}}$ must vanish. This contradicts with the definition (4.20) of $\theta$. Thus we have $\theta=\theta_{r}$, and it follows that
$z^{r}=\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}.$ (4.23)
In particular, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$. The theorem is proved. ∎
Remark 4.1.
This theorem shows that the proper generators are polynomials of the flat coordinates of $\eta$. It deserves to note that this property also holds true for the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and on the extended affine Weyl groups [12, 14]. For the Frobenius manifold structures constructed in [9] and [12], the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s, while for the ones constructed in [14], $t^{\alpha}$ may involve radical expressions of $z^{j}$’s. For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell})$ for $\ell\geq 5$.
From Theorem 4.2, one can verify that each $g^{\alpha\beta}$ is a rational function of $t^{1},\cdots,t^{\ell}$. Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\overline{\mathcal{M}}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
$\eta^{\gamma\nu}\frac{\partial^{3}F}{\partial t^{\alpha}\partial t^{\beta}\partial t^{\nu}}=c^{\gamma}_{\alpha\beta},\quad\mathcal{L}_{\tilde{E}}F=2F+\textrm{quadratic functions of}\ t^{1},\dots,t^{\ell},$
up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$, here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.53). From the relations (3.38) and (3.53) it follows that we can choose the potential $F$ such that
$\eta^{\alpha\nu}\eta^{\beta\xi}\frac{\partial^{2}F}{\partial t^{\nu}\partial t^{\xi}}=\frac{\kappa}{d_{\alpha}+d_{\beta}}g^{\alpha\beta}.$ (4.24)
Therefore, all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
Corollary 4.3.
Define the covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$ as (3.59). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\overline{\mathcal{M}}$.
Proof.
To prove that $(t^{1},\cdots,t^{\ell})$ are global coordinates, it suffices to show that
$(t^{1},\cdots,t^{\ell})\colon\overline{\mathcal{M}}\to\mathbb{C}^{\ell}$ (4.25)
is injective. Suppose $p,p^{\prime}\in\overline{\mathcal{M}}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~{}\alpha=1,\cdots,\ell$, then by using Theorem 4.2 we know that
$z^{r}(t^{1}(p),\cdots,t^{\ell}(p))=z^{r}(t^{1}(p^{\prime}),\cdots,t^{\ell}(p^{\prime})).$
This implies that $\overline{\pi}(p)=\overline{\pi}(p^{\prime})$, where $\overline{\pi}:\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is the canonical projection. Therefore $p\sim p^{\prime}$, i.e., they are identified on $\overline{\mathcal{M}}$. Hence (4.25) is injective, and we proved the corollary. ∎
In conclusion, we have the following commutative diagram:
[…]
In other words, upon taking coordinates $(t^{1},\cdots,t^{\ell})$ on $\overline{\mathcal{M}}$ and $(z^{1},\cdots,z^{\ell})$ on $\mathcal{M}\setminus D$, the projection $\overline{\pi}$ coincides with the natural map $\mathrm{Spec}(\mathbb{C}[t^{1},\cdots,t^{\ell}])\to\mathrm{Spec}(\mathbb{C}[z^{1},\cdots,z^{\ell}])$.
4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and it can be extended directly to generalized Frobenius manifolds. Let us first recall this defintion.
Let $M$ be an $\ell$-dimensional generalized Frobenius manifold, of which the flat metric is $\eta$ and the intersection form is $g$, and let $t^{1},\cdots,t^{\ell}$ be flat coordinates of $\eta$. Consider the system of equations
$\frac{\partial^{2}u}{\partial t^{\alpha}\partial t^{\beta}}=\Gamma_{\alpha\beta}^{\epsilon}\frac{\partial u}{\partial t^{\epsilon}},\quad\alpha,\beta=1,\dots,\ell$ (4.26)
for the flat coordinates of $g$. Here $\Gamma_{\alpha\beta}^{\epsilon}$ are the Christoffel symbols of the Levi-Civita connection of $g$. Denote $\Sigma=\{t\in M~{}|~{}\det g(t)=0\}$, then the solutions $u$ live on the universal covering $\widehat{M\setminus\Sigma}$ of $M\setminus\Sigma$. Define the solution space
[…]
then we know that $\dim\mathcal{S}=n$, and there exist flat coordinates $u^{1},\cdots,u^{\ell}$ of $g$ such that
$\mathcal{S}=\mathrm{span}\{u^{1},\cdots,u^{\ell}\}.$ (4.28)
For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, we have a transformation
$\gamma^{*}u^{\alpha}=A_{\beta}^{\alpha}(\gamma)u^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.29)
on $\mathcal{S}$ by analytic continuation. Denote by $E^{\ell}$ the dual space of $\mathcal{S}$, then it has coordinates $u^{1},\cdots,u^{\ell}$. Hence the transformation (4.29) yields a transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$. Since $E^{\ell}$ is an $\ell$-dimensional complexified Euclidean space $\mathbb{R}^{\ell}\otimes\mathbb{C}$ equipped with the metric $g$, and the action by $\pi_{1}(M\setminus\Sigma)$ preserves the metric $g$, this transformation on $E^{\ell}$ is an isometry. Hence we have obtained a representation
$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}.$ (4.30)
Definition 4.1 ([9], Appendix G.).
The image of $r$ given in (4.30) is called the monodromy group of $M$, which is denoted by
$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.31)
Now we proceed to compute the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Let
$\overline{\Sigma}=\{t\in\overline{\mathcal{M}}\mid\det\left(g^{\alpha\beta}(t)\right)=0\},$
and $\mathcal{U}(\mathcal{M})$ be the universal covering space of ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$. The flat coordiantes of $g$ are $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$, and they are functions on $\mathcal{U}(\mathcal{M})$. Moreover, $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so we can view the complexified Euclidean space $E^{\ell}$ as $V\otimes\mathbb{C}$. Hence we obtain a representation
$r\colon\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.32)
where
$\gamma^{*}x^{i}\bigr{|}_{\lambda=0}=A_{j}^{i}(\gamma)x^{j}\bigr{|}_{\lambda=0}+B^{i}(\gamma).$ (4.33)
Here the image $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ is the monodromy group of $\overline{\mathcal{M}}$. However, this group is difficult to compute, and we will instead consider a group bigger than it.
Recall that the covering map
$\overline{\pi}\colon\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ (4.34)
is nondegerated, so from
$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)$ (4.35)
we know that $\left(g^{\alpha\beta}(t)\right)$ degenerates if and only if $\left(g^{ij}(z(t)\right)$ degenerates. In other words, denote
$\Sigma^{\prime}=\{z\in\mathcal{M}\setminus D\mid\det\left(g^{ij}(z)\right)=0\},$
then we have $(\overline{\pi})^{-1}(\Sigma^{\prime})=\overline{\Sigma}$, and the map
$\overline{\pi}:\overline{\mathcal{M}}\setminus\overline{\Sigma}\to\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ (4.36)
is also a covering map, hence
$\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\leq\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})).$ (4.37)
Since ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$ is a covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, they share the same universal covering space $\mathcal{U}(\mathcal{M})$, and both $\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})$ and $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ act on $\mathcal{U}(\mathcal{M})$. Note that $g^{ij}(z)$ are polynomials of $z^{1},\cdots,z^{\ell}$, so in particular they are single valued functions on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. This implies that the action of $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ on $\mathcal{U}(\mathcal{M})$ also leaves the metric $g$ invariant, so the representation (4.32) can be extended to a representation
$r\colon\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.38)
Thus it follows from (4.37) that
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))~{}\leq~{}r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))),$ (4.39)
i.e., the monodromy group of $\overline{\mathcal{M}}$ is a subgroup of $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$. We are to show in what follows that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
Recall that
$\varphi_{0}\colon V\otimes\mathbb{C}\to\mathcal{M}$ (4.40)
is a quotient map. By eliminating a divisor, we obtain a covering map
$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})\to\mathcal{M}\setminus(D\cup\Sigma^{\prime}).$ (4.41)
Since $\mathcal{U}(\mathcal{M})$ is the universal covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, it is also a covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$, and the covering map is given by
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\colon\mathcal{U}(\mathcal{M})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime}),\quad p\mapsto(x^{1}\bigr{|}_{\lambda=0}(p),\cdots,x^{\ell}\bigr{|}_{\lambda=0}(p)).$ (4.42)
In conclusion, we have the following commutative diagram of covering spaces:
$\mathcal{U}(\mathcal{M})$¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathrm{pr}_{1}$$\overline{\pi}$$\mathrm{pr}_{2}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\varphi_{0}$
All the maps in the diagram are covering maps. From the property of universal covering spaces, we have
$\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))=\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\right).$ (4.43)
Here $\mathrm{Deck}(X/Y)$ is the deck transformation group on $X$ over $Y$, where $X$ is a covering space over $Y$. Relation (4.43) implies that, every element $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ induces a deck transformation on $\mathcal{U}(\mathcal{M})$. We have the following lemma for $r(\gamma)$.
Lemma 4.4.
Let $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, and let us view $\gamma$ as a deck transformation on $\mathcal{U}(\mathcal{M})$. Then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ such that
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\circ\gamma=p\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.44)
In other words, we have the following commutative diagram.
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$\exists!~{}p=r(\gamma)$$\varphi_{0}$$\varphi_{0}$$\gamma$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$
Proof.
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, the map $p$ satisfying (4.44) is uniquely determined. Moreover, since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M})$, $p=r(\gamma)$ satisfies (4.44). Thus in order to prove the lemma, it suffices to verify that $r(\gamma)$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$. As $\gamma$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, we have
$\varphi_{0}\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}=\varphi_{0}\circ\hat{\mathbf{x}}\Big{|}_{\lambda=0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.45)
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, it follows that
$\varphi_{0}=\varphi_{0}\circ r(\gamma).$
Therefore, $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. The lemma is proved. ∎
Proposition 4.5.
The group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ can be represented as
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Iso}(V\otimes\mathbb{C})\cap\mathrm{Deck}\bigg{(}\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\bigg{)}.$ (4.46)
The right hand side consists of all the isometries on $V\otimes\mathbb{C}$ whose restrictions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ are deck transformations over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
Proof.
For any $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, $r(\gamma)$ is of course in $\mathrm{Iso}(V\otimes\mathbb{C})$. Moreover, by using Lemma 4.4, we know that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
On the other hand, suppose $p$ be an isometry of $V\otimes\mathbb{C}$ and also be a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. By the general lifting lemma of covering spaces, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M})\to\mathcal{U}(\mathcal{M})$, since
$p_{*}\left(\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\right)_{*}\pi_{1}(\mathcal{U}(\mathcal{M}))=\{e\}.$ (4.47)
Therefore, we have the following commutative diagram:
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$p$$\varphi_{0}$$\varphi_{0}$$\sigma$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$
Thus by using Lemma 4.4, we have $p=r(\sigma)$. The proposition is proved. ∎
Remark 4.2.
Note that $z^{1},\cdots,z^{\ell}$ are coordinates on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, so intuitively speaking, $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(z^{1},\cdots,z^{\ell})$. Similarly, $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(t^{1},\cdots,t^{\ell})$. From Theorem 4.2 we know that, if $t^{1},\cdots,t^{\ell}$ are preserved under some transformations, so are $z^{1},\cdots,z^{\ell}$. This fact implies the relation (4.39).
We are ready to present the following theorem.
Theorem 4.6.
For the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$, the group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\rtimes\mathbb{Z}^{\ell}$. More precisely,
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))\cong\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.48)
Here $\mathrm{Stab}(\omega)$ is the stabilizer of $\omega$ in $W(R)$.
Proof.
Due to Proposition 4.5, we only need to determine all the isometric actions on $x^{i}\bigr{|}_{\lambda=0}$ that preserve $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., preserve $z^{1},\cdots,z^{\ell}$. From (2.31) and (2.41) it follows that
$z^{j}\bigr{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (4.49)
Here
$H_{j}=\{\sigma\in W(R)\mid(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\}$
is a subgroup of $W(R)$. For any given element $\tilde{\phi}\in\mathrm{Iso}(E^{\ell})$, we decompose it into linear and the affine parts as follows:
$\tilde{\phi}(\mathbf{t})=\phi(\mathbf{t})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.50)
It acts on the the Fourier polynomials $z^{j}$ by
$\tilde{\phi}^{*}z^{j}\big{|}_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.51)
$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.52)
so it preserves $z^{j}\bigr{|}_{\lambda=0}$, $j=1,\cdots,\ell$ if and only if $\phi$ induces a permutation on $H_{j}$ by right action $\sigma\to\sigma\phi$ and
$(\sigma^{*}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
is always an integer for each $\sigma\in H_{j}$, $j=1,\cdots,\ell$. These conditions are equivalent to
$\phi\in\bigcap_{j=1}^{\ell}H_{j},\quad v^{1},\cdots,v^{\ell}\in\mathbb{Z}.$ (4.53)
This implies that
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\bigcap_{j=1}^{\ell}H_{j}\rtimes\mathbb{Z}^{\ell}.$ (4.54)
Now let us compute $\bigcap_{j=1}^{\ell}H_{j}$. Note that $\sigma\in H_{j}$, $j=1,\cdots,\ell$ if and only if
$(\omega_{j},\omega)=(\omega_{j},\sigma(\omega)),\quad j=1,\cdots,\ell.$ (4.55)
Since $\omega_{j},j=1,\cdots,\ell$ span $V\otimes\mathbb{C}$, (4.55) is equivalent to $\sigma(\omega)=\omega$. Therefore,
$\bigcap_{j=1}^{\ell}H_{j}=\mathrm{Stab}(\omega).$ (4.56)
It follows from (4.54) and (4.56) that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$. The theorem is proved. ∎
By using (4.39), we have the following corollary.
Corollary 4.7.
The monodromy group of the generalized Frobenius manifold $\mathcal{M}(R,\omega)$ is isomorphic to a subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
It comes out a natural question to ask: is $\mathrm{Mono}(\mathcal{M}(R,\omega))$ exactly isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$? Unfortunately, the answer to this question is negative, and we will see counterexamples in Sec. 5. However, in some special cases, the isomorphism holds true. One simple case is when ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. In this case we have ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}\cong(\mathcal{M}\setminus D)\setminus\Sigma^{\prime}$, hence
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))=r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.57)
We will also see examples which belong to this case in Sec. 5.
5. Examples
[…]
We start with the simplest example.
Example 5.1 ($(R,\omega)=(A_{1},\omega_{1})$).
Let $R$ be the root system of type $A_{1}$, $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
$\alpha_{1}=e_{1}-e_{2}.$
We take $\omega$ to be the fundamental weight $\omega_{1}=\frac{1}{2}\alpha^{\vee}_{1}=\frac{1}{2}\alpha_{1}$. Let us introduce the coordinate $x_{1}$ on $V$ by
$\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1},\quad\mathbf{x}\in V.$
In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{1}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{1},\mathbf{x})\alpha_{1},\quad\mathbf{x}\in V.$
Then we have
$\sigma_{1}(\mathbf{x})=c\omega_{1}+(-c-x_{1})\alpha^{\vee}_{1},\quad x_{1}\mapsto-c-x_{1}.$
The invariant Fourier polynomial has the form
$Y_{1}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
and we take
$y_{1}=e^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
We have
$(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2},$
and
$g^{11}_{\lambda}=\frac{1}{4\pi^{2}}(\operatorname{d\!}y_{1},\operatorname{d\!}y_{1})=\frac{1}{4\pi^{2}}\frac{\partial y_{1}}{\partial x_{1}}(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})\frac{\partial y_{1}}{\partial x_{1}}=-\frac{1}{2}y_{1}^{2}+2\lambda.$
Therefore, here $\{y_{1}\}$ is a pencil generator, and we obtain a flat pencil
$\eta^{11}=2,\quad g^{11}=-\frac{1}{2}y_{1}^{2}.$
Now we consider the flat coordinate $t_{1}=y_{1}/\sqrt{2}$ of $\eta$, then in the new coordinate the flat pencil has the form
$\eta^{11}=1,\quad g^{11}=-\frac{1}{2}t_{1}^{2}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{1},\omega_{1})$ with potential
$F=-\frac{1}{24}t_{1}^{4}.$
The Euler vector field and the unit vector field are given by
$E=\frac{1}{2}t_{1}\frac{\partial}{\partial t_{1}},\quad e=-\frac{1}{t_{1}}\frac{\partial}{\partial t_{1}}.$
Moreover, the flat coordinate of $g$ is given by
$x^{1}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}(\log(\sqrt{2}t_{1})).$
It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $t_{1}$. Hence the monodromy group of the generalized Frobenius manifold is given by
$\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathbb{Z}.$
We considered the construction of generalized Frobenius manifolds for different types of affine Weyl groups with appropriately chosen weights. In what follows we will give some representative examples, which are chosen based on the following points:
  1. 1.
    In Example 5.2, the basic generators $y^{1},\cdots,y^{\ell}$ are pencil generators. The covering space ¯$\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is trivial, thus
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$
  2. 2.
    In Example 5.3, the basic generators $y^{1},\cdots,y^{\ell}$ are also pencil generators. However, unlike the previous example, the monodromy group
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{\ell}$
    is a proper subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
  3. 3.
    In Example 5.4, the basic generators $y^{1},y^{2},y^{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y^{1},y^{2}$ and $y^{3}$, i.e.,
    $z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=y_{3}.$
    Here $a_{1},a_{2}$ are some constants, and we find three possible values for them.
  4. 4.
    In Example 5.5, $y^{1},y^{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
    $z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
  5. 5.
    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight. This complicates the construction of pencil generators and the generalized Frobenius manifold structure.
These examples involve most issues that we may come across in the computations. Something interesting issues will also happen when we consider other types of affine Weyl groups. More details will be presented in our subsequent publications.
Example 5.2 ($(R,\omega)=(A_{2},\omega_{2})$).
We take the root system $R$ of type $A_{2}$. Let $V$ be the hyperplane of $\mathbb{R}^{3}$ spanned by
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}.$
We have the coroots and fundamental weights
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
We choose $\omega=\omega_{2}$. Let us introduce the coordinates $x_{1},x_{2}$ of $V$ by
$\mathbf{x}=c\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
The invariant $\lambda$-Fourier polynomials are given by
$Y_{1}$ $=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{2}-x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{2}+\frac{2}{3}c\right)},$
$Y_{2}$ $=\mathrm{e}^{2\pi i\left(x_{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{1}-x_{2}-\frac{1}{3}c\right)}.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$ (5.1)
We take
$y_{1}$ $=\lambda^{\frac{1}{3}}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\lambda\mathrm{e}^{-2\pi ix_{2}},$
$y_{2}$ $=\lambda^{\frac{2}{3}}Y_{2}=\mathrm{e}^{2\pi ix_{2}}+\lambda\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}.$
The metric on $V$ that is induced from $\mathbb{R}^{3}$ is given by
$\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix},\quad(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})={\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}}^{-1}=\frac{1}{3}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}.$
In the coordinates $y_{1},y_{2}$ we have
$\left(g_{\lambda}^{ij}\right)=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{2}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&3\lambda-\frac{1}{3}y_{1}y_{2}\\[3.0pt] 3\lambda-\frac{1}{3}y_{1}y_{2}&2\lambda y_{1}-\frac{2}{3}y_{2}^{2}\end{pmatrix}.$
Then $y^{1},y^{2}$ are pencil generators, and the corresponding flat pencil is given by
$\left(\eta^{ij}\right)=\begin{pmatrix}0&3\\ 3&2y_{1}\end{pmatrix},\quad\left(g^{ij}\right)=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&-\frac{1}{3}y_{1}y_{2}\\[3.0pt] -\frac{1}{3}y_{1}y_{2}&-\frac{2}{3}y_{2}^{2}\end{pmatrix}.$
The metric $\eta$ has the flat coordinates
$t_{1}=\frac{1}{3}y_{1},\quad t_{2}=y_{2}-\frac{1}{6}y_{1}^{2},\quad$
in which the metrics $\eta$ and $g$ have the form
$\left(\eta^{ij}\right)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\left(g^{ij}\right)=\begin{pmatrix}\frac{1}{9}\left(2t_{2}-3t_{1}^{2}\right)&-t_{1}t_{2}+\frac{1}{2}t_{1}^{3}\\[4.0pt] -t_{1}t_{2}+\frac{1}{2}t_{1}^{3}&-\frac{1}{6}\left(2t_{2}-3t_{1}^{2}\right)^{2}\end{pmatrix}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{2},\omega_{2})$ with potential
$F=\frac{1}{18}t_{2}^{3}-\frac{1}{4}t_{1}^{2}t_{2}^{2}+\frac{1}{8}t_{1}^{4}t_{2}-\frac{3}{80}t_{1}^{6},$ (5.2)
the Euler vector field
$E=\frac{1}{3}t_{1}\frac{\partial}{\partial t_{1}}+\frac{2}{3}t_{2}\frac{\partial}{\partial t_{2}},$
and the unit vector field
$e=-\frac{2}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{1}}-\frac{6t_{1}}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{2}}.$
We also have $e^{\flat}=-\mathrm{d}\log y_{2}$. Note that in this example, not only $y_{1},y_{2}$ are polynomials of $t^{1}$ and $t^{2}$, but also $t^{1},t^{2}$ are polynomials of $y_{1}$ and $y_{2}$. This implies that the monodromy between $t^{1}$, $t^{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. Therefore,
$\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))=\mathrm{Stab}(\omega_{2})\rtimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\rtimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
More explicitly, $\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))$ is generated by
$\sigma_{1}\colon$ $x_{1}\to x_{2}-x_{1},\quad x_{2}\to x_{2},$
$T_{1}\colon$ $x_{1}\to x_{1}+1,\quad x_{2}\to x_{2},$
$T_{2}\colon$ $x_{1}\to x_{1},\quad x_{2}\to x_{2}+1.$
Example 5.3 ($(R,\omega)=(C_{3},\omega_{3})$).
Let $R$ be the root system of type $C_{3}$, and $V=\mathbb{R}^{3}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=2e_{3}.$
Then we have the coroots
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3}.$
The fundamental weights are given by
$\omega_{1}=\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\alpha^{\vee}_{3},\quad\omega_{2}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+2\alpha^{\vee}_{3},\quad\omega_{3}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+3\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{3}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c\omega_{3}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
Since $\kappa=2$, we have
$\lambda=\mathrm{e}^{-4\pi ic},$
then the $\lambda$-Fourier invariant polynomials are given by
$y_{1}=\lambda^{\frac{1}{2}}(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{3}{2}}\xi_{1}\xi_{2}\xi_{3},$
where
$\xi_{1}$ $=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi ix_{1}}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi ix_{1}},\quad\xi_{2}=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi i(x_{2}-x_{1})},$
$\xi_{3}$ $=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi i(x_{3}-x_{2})}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi i(x_{3}-x_{2})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&1\end{pmatrix}^{-1}=\begin{pmatrix}1&1&1\\ 1&2&2\\ 1&2&3\end{pmatrix}.$
In the coordinates $y_{1},y_{2},y_{3}$ we have
$\left(g_{\lambda}^{ij}\right)$ $=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=\begin{pmatrix}-y_{1}^{2}+2y_{2}+12\lambda&-y_{1}y_{2}+3y_{3}+8\lambda y_{1}&-y_{1}y_{3}+4\lambda y_{2}\\ -y_{1}y_{2}+3y_{3}+8\lambda y_{1}&-2y_{2}^{2}+2y_{1}y_{3}-8\lambda y_{2}+8\lambda y_{1}^{2}&-2y_{2}y_{3}+4\lambda y_{1}y_{2}-12\lambda y_{3}\\ -y_{1}y_{3}+4\lambda y_{2}&-2y_{2}y_{3}+4\lambda y_{1}y_{2}-12\lambda y_{3}&-3y_{3}^{2}-8\lambda y_{1}y_{3}+4\lambda y_{2}^{2}\end{pmatrix}.$
It follows that $y_{1},y_{2},y_{3}$ are pencil generators, and the corresponding flat pencil is given by
$\left(\eta^{ij}\right)$ $=\begin{pmatrix}12&8y_{1}&4y_{2}\\ 8y_{1}&-8y_{2}+8y_{1}^{2}&4y_{1}y_{2}-12y_{3}\\ 4y_{2}&4y_{1}y_{2}-12y_{3}&-8y_{1}y_{3}+4y_{2}^{2}\end{pmatrix},$
$\left(g^{ij}\right)$ $=\begin{pmatrix}-y_{1}^{2}+2y_{2}&-y_{1}y_{2}+3y_{3}&-y_{1}y_{3}\\ -y_{1}y_{2}+3y_{3}&-2y_{2}^{2}+2y_{1}y_{3}&-2y_{2}y_{3}\\ -y_{1}y_{3}&-2y_{2}y_{3}&-3y_{3}^{2}\end{pmatrix}.$
The flat coordinates $t_{1},t_{2},t_{3}$ of $\eta$ are given by the polynomials
$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{2}^{3}+t_{1}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
In these flat coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
Thus we obtain the a generalized Frobenius manifold $\mathcal{M}(C_{3},\omega_{3})$ with potential
$F=-\frac{1}{9}t_{1}^{3}t_{2}-\frac{1}{36}t_{2}^{4}-\frac{1}{3}t_{1}t_{2}^{2}t_{3}-\frac{1}{6}t_{1}^{2}t_{3}^{2}-\frac{1}{9}t_{2}t_{3}^{3}.$
The Euler vector field and the unit vector field are given by
$E=$ $\,\frac{1}{2}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{2}t_{3}\frac{\partial}{\partial t_{3}},$
$e=$ $\,\frac{3}{2\left(t_{1}^{3}+t_{2}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right)}$
×$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
We also have $e^{\flat}=-\frac{1}{2}\mathrm{d}\log y_{3}$.
Note that, if we take the flat coordinates $v_{1},v_{2},v_{3}$ of $\eta$ via a linear transformation
$t_{1}$ $=\frac{1}{\sqrt{3}}(v_{1}+\zeta v_{2}+\zeta^{2}v_{3}),$
$t_{2}$ $=\frac{1}{\sqrt{3}}(v_{1}+v_{2}+v_{3}),$
$t_{3}$ $=\frac{1}{\sqrt{3}}(v_{1}+\zeta^{2}v_{2}+\zeta v_{3}),$
where $\zeta=\frac{-1+\sqrt{3}i}{2}$, then in these coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},$
and the metric $g$ has the form
$\left(g^{\alpha\beta}\right)=\begin{pmatrix}-v_{1}^{2}&0&0\\ 0&-v_{2}^{2}&0\\ 0&0&-v_{3}^{2}\end{pmatrix}.$
The potential of the generalized Frobenius manifold is given by
$F=-\frac{1}{12}v_{1}^{4}-\frac{1}{12}v_{2}^{4}-\frac{1}{12}v_{3}^{4},$
the Euler vector field and the unit vector field have the form
$E=\frac{1}{2}v^{1}\frac{\partial}{\partial v^{1}}+\frac{1}{2}v^{2}\frac{\partial}{\partial v^{2}}+\frac{1}{2}v^{3}\frac{\partial}{\partial v^{3}},\quad e=-\frac{1}{2v^{1}}\frac{\partial}{\partial v^{1}}-\frac{1}{2v^{2}}\frac{\partial}{\partial v^{2}}-\frac{1}{2v^{3}}\frac{\partial}{\partial v^{3}}.$
It is clear that, up to a factor, we have the isomorphism
$\mathcal{M}(C_{3},\omega_{3})~{}\cong~{}\mathcal{M}(A_{1},\omega_{1})^{\oplus 3}.$
Moreover, since the flat coordinates of $g$ has the form
$x_{1}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\log(2v_{1}),\quad x_{2}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\left(\log(2v_{2})+\log(2v_{1})\right),$
$x_{3}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\left(\frac{1}{2}\log(2v_{3})+\frac{1}{2}\log(2v_{2})+\frac{1}{2}\log(2v_{1})\right),$
we know that the monodromy group is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
$\mathrm{Mono}(\mathcal{M}(C_{3},\omega_{3}))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{3}~{}<~{}\mathrm{Stab}(\omega_{3})\rtimes\mathbb{Z}^{3}.$ (5.3)
Here $\mathrm{Stab}(\omega_{3})$ is nontrivial since $\sigma_{1},\sigma_{2}\in\mathrm{Stab}(\omega_{3})$. This example shows that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ is not always isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
Example 5.4 ($(R,\omega)=(B_{3},\omega_{1})$).
Let $R$ be the root system of type $B_{3}$, and $V=\mathbb{R}^{3}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}.$
Then we have the coroots
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=2e_{3}.$
The fundamental weights are given by
$\omega_{1}=\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{1}{2}\alpha^{\vee}_{3},\quad\omega_{2}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+\alpha^{\vee}_{3},\quad\omega_{3}=\frac{1}{2}\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{3}{4}\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{1}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
$y_{1}=\lambda(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{1}{2}}\tilde{\xi}_{1}\tilde{\xi}_{2}\tilde{\xi}_{3}.$
Here
$\xi_{1}$ $=\lambda^{-1}\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}},\quad\xi_{2}=\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\mathrm{e}^{-2\pi i(x_{2}-x_{1})},$
$\xi_{3}$ $=\mathrm{e}^{2\pi i(2x_{3}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{3}-x_{2})},$
and
$\tilde{\xi}_{1}$ $=\lambda^{-\frac{1}{2}}e^{\pi ix_{1}}+\lambda^{\frac{1}{2}}\mathrm{e}^{-\pi ix_{1}},\quad\tilde{\xi}_{2}=e^{\pi i(x_{2}-x_{1})}+e^{-\pi i(x_{2}-x_{1})},$
$\tilde{\xi}_{3}$ $=e^{\pi i(2x_{3}-x_{2})}+e^{-\pi i(2x_{3}-x_{2})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((dx_{i},dx_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-2\\ 0&-2&4\end{pmatrix}^{-1}=\begin{pmatrix}1&1&\frac{1}{2}\\ 1&2&1\\ \frac{1}{2}&1&\frac{3}{4}\end{pmatrix}.$
In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. Let us consider the proper generators with the form
$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=z_{3},$
here $a_{1},a_{2}$ are two constants. Then there are three possible choices of $a_{1}$ and $a_{2}$ such that $z_{1},z_{2}$ and $z_{3}$ form a set of pencil generators, they are given by
$a_{1}=6,\quad a_{2}=-12.$ (5.4)
$a_{1}=2,\quad a_{2}=4,$ (5.5)
$a_{1}=-2,\quad a_{2}=4.$ (5.6)
We first consider the $a_{1}=6,a_{2}=-12$ case. In this case we have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}12z_{1}+2z_{2}&3z_{3}^{2}-16z_{1}&6z_{3}\\ 3z_{3}^{2}-16z_{1}&-32z_{2}-12z_{3}^{2}-64z_{1}&-24z_{3}\\ 6z_{3}&-24z_{3}&0\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}$ $=-\frac{1}{1920z_{3}^{7/3}}\left(1280z_{1}^{2}+640z_{1}z_{2}+80z_{2}^{2}-480z_{1}z_{3}^{2}+120z_{2}z_{3}^{2}-27z_{3}^{4}\right),$
$t_{2}$ $=\frac{1}{8z_{3}}(16z_{1}+4z_{2}-z_{3}^{2}),$
$t_{3}$ $=z_{3}^{1/3}.$
In these coordinates we have
$(\eta^{ij})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},1)$ with potential
$F=$ $\,\frac{t_{1}^{2}t_{2}}{3t_{3}}-\frac{t_{1}^{2}\ t_{3}^{2}}{12}-\frac{2t_{1}t_{2}^{3}}{27t_{3}^{2}}-\frac{1}{36}\ t_{1}t_{2}^{2}t_{3}+\frac{1}{360}t_{1}t_{2}\ t_{3}^{4}-\frac{t_{1}t_{3}^{7}}{30240}$
$\,+\frac{t_{2}^{5}}{135\ t_{3}^{3}}-\frac{t_{2}^{4}}{432}+\frac{t_{2}^{3}\ t_{3}^{3}}{2160}-\frac{t_{2}^{2}t_{3}^{6}}{8640}+\frac{t_{2}\ t_{3}^{9}}{172800}-\frac{t_{3}^{12}}{7603200}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{5}{6}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{6}t_{3}\frac{\partial}{\partial t_{3}},$
$e$ $=-\frac{6}{D}\left(\left(40t_{1}+20t_{2}t_{3}^{2}+t_{3}^{5}\right)\frac{\partial}{\partial t_{1}}+\left(40t_{2}+20t_{3}^{3}\right)\frac{\partial}{\partial t_{2}}+40t_{3}\frac{\partial}{\partial t_{3}}\right).$
Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
Now we consider the $a_{1}=2,a_{2}=4$ case. We have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}4z_{1}+2z_{2}&3z_{3}^{2}-4z_{2}&4z_{3}\\ 3z_{3}^{2}-4z_{2}&16z_{2}-4z_{3}^{2}&0\\ 4z_{3}&0&0\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}=\frac{48z_{1}+12z_{2}-5z_{3}^{2}}{48\sqrt{z_{3}}},\quad t_{2}=\frac{1}{4}\sqrt{4z_{2}-z_{3}^{2}},\quad t_{3}=\frac{1}{2}\sqrt{z_{3}}.$
In these coordinates we have
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},2)$ with potential
$F=$ $\,-\frac{t_{2}^{4}}{48}+\frac{t_{1}^{3}}{24t_{3}}-\frac{1}{4}t_{1}t_{2}^{2}t_{3}-\frac{t_{1}^{2}t_{3}^{2}}{12}+\frac{t_{2}^{2}t_{3}^{4}}{24}+\frac{t_{1}t_{3}^{5}}{180}-\frac{t_{3}^{8}}{2268}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{3}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{4}t_{3}\frac{\partial}{\partial t_{3}},$
$e$ $=\frac{1}{3t_{2}^{2}-6t_{1}t_{3}-2t_{3}^{4}}\left(\left(6t_{1}+8t_{3}^{3}\right)\frac{\partial}{\partial t_{1}}-6t_{2}\frac{\partial}{\partial t_{2}}+6t_{3}\frac{\partial}{\partial t_{3}}\right).$
We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
Finally, let us consider the $a_{1}=-2,a_{2}=4$ case. We have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}-4z_{1}+2z_{2}&3z_{3}^{2}-8z_{2}&2z_{3}\\ 3z_{3}^{2}-8z_{2}&4z_{3}^{2}&8z_{3}\\ 2z_{3}&8z_{3}&16\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}=-\frac{48z_{1}+8z_{2}-5z_{3}^{2}}{48\left(4z_{2}-z_{3}^{2}\right)^{1/4}},\quad t_{2}=\frac{1}{4}z_{3},\quad t_{3}=\frac{1}{2}\left(4z_{2}-z_{3}^{2}\right)^{1/4}.$
In these coordinates we have
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix},$
and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$. In all cases, the monodromy groups are given by
$\mathrm{Mono}(\mathcal{M}(B_{3},\omega_{1},k))=\mathrm{Stab}(\omega_{1})\rtimes\mathbb{Z}^{3}=\langle\sigma_{2},\sigma_{3}\rangle\rtimes\mathbb{Z}^{3}\cong D_{4}\rtimes\mathbb{Z}^{3},\quad k=1,2,3.$
Example 5.5 ($(R,\omega)=(G_{2},\omega_{2})$).
Let $R$ be the root system of type $G_{2}$, and $V$ be the subspace of $\mathbb{R}^{3}$ spanned by the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=-2e_{1}+e_{2}+e_{3}.$
We have the coroots
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\frac{1}{3}\alpha_{2}.$
The fundamental weights are given by
$\omega_{1}=2\alpha^{\vee}_{1}+3\alpha^{\vee}_{2},\quad\omega_{2}=3\alpha^{\vee}_{1}+6\alpha^{\vee}_{2}.$
Let us take $\omega=\omega_{2}$ and introduce coordinates $x_{1},x_{2}$ of $V$ by
$\mathbf{x}=c\,\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}=(x_{1}+3c)\alpha^{\vee}_{1}+(x_{2}+6c)\alpha^{\vee}_{2},$
Since $\kappa=3$, we have
$\lambda=\mathrm{e}^{-6\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
$y_{1}=$ $\,\mathrm{e}^{-6\pi ic}\left(\mathrm{e}^{2\pi i(x_{1}+3c)}+\mathrm{e}^{-2\pi i(x_{1}+3c)}+\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{1}-x_{2}-3c)}+\mathrm{e}^{-2\pi i(x_{1}-x_{2}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}$
$+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
$y_{2}=$ $\,e^{-12\pi ic}\left(\mathrm{e}^{2\pi i(x_{2}+6c)}+\mathrm{e}^{-2\pi i(x_{2}+6c)}+\mathrm{e}^{2\pi i(3x_{1}-2x_{2}-3c)}+\mathrm{e}^{-2\pi i(3x_{1}-2x_{2}-3c)}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{2}-3x_{1}-3c)}+\mathrm{e}^{-2\pi i(x_{2}-3x_{1}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}$
$\,+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1\\ -1&\frac{2}{3}\end{pmatrix}^{-1}=\begin{pmatrix}2&3\\[3.0pt] 3&6\end{pmatrix}.$
In the $y_{1},y_{2}$ coordinates the metric does not depend at most linearly on $\lambda$, hence $y_{1}$ and $y_{2}$ are not pencil generators. Proper generators $z_{1}$ and $z_{2}$ may be introduced by
$z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
Then in the $z_{1},z_{2}$ coordinates we have
$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$
The metric $\eta$ has flat coordinates
$t_{1}=\frac{z_{1}}{2\left(4z_{2}-z_{1}^{2}\right)^{\frac{1}{6}}},\quad t_{2}=-\frac{1}{3}\left(4z_{2}-z_{1}^{2}\right)^{\frac{1}{6}},$
in these coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$
We thus obtain a generalized Frobenius manifold with potential
$F=\frac{t_{1}^{4}}{2916t_{2}^{2}}-\frac{1}{8}t_{1}^{2}t_{2}^{2}+\frac{81t_{2}^{6}}{320}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{2}{3}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{3}t_{2}\frac{\partial}{\partial t_{1}},$
$e$ $=-\frac{2}{3t_{2}\left(4t_{1}^{2}+81t_{2}^{4}\right)}\left((4t_{1}^{2}+243t_{2}^{4})\frac{\partial}{\partial t_{1}}+4t_{1}t_{2}\frac{\partial}{\partial t_{2}}\right).$
In this case, $\mathrm{Stab}(\omega_{1})=\langle\sigma_{1}\rangle$, where
$\sigma_{1}(x_{1},x_{2})=(-x_{1}+x_{2},x_{2}).$
However, one has
$t_{1}=\frac{z_{1}}{2\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}},\quad t_{2}=-\frac{1}{3}\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}.$
Hence the actions of $\sigma_{1}$ on $t_{1}$ and $t_{2}$ is not trivial. In fact,
[…]
Here the value of $3]{-1}$ depends on the choice of the monodromy region of $f(z)=z^{1/3}$. Therefore we have
$\mathrm{Mono}(\mathcal{M}(G_{2},\omega_{2}))=\{e\}\rtimes\mathbb{Z}^{2}.$
Example 5.6 ($(R,\omega)=(A_{3},\omega_{1}+\omega_{3})$).
Let $R$ be the root system of type $A_{3}$, $V$ be the subspace of $\mathbb{R}^{4}$ spanned by the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}-e_{4}.$
We have $\alpha^{\vee}_{i}=\alpha_{i},i=1,2,3$, and the fundamental weights are given by
$\omega_{1}=\frac{3}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{1}{4}\alpha^{\vee}_{3},\quad\omega_{2}=\frac{1}{2}\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{1}{2}\alpha^{\vee}_{3},\quad\omega_{3}=\frac{1}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{3}{4}\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{1}+\omega_{3}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c(\omega_{1}+\omega_{3})+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3},\quad\mathbf{x}\in V.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
We have the $\lambda$-invariant Fourier polynomials
$y_{1}=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{2})}+\lambda^{2}\mathrm{e}^{-2\pi ix_{3}}+\lambda\mathrm{e}^{-2\pi i(x_{2}-x_{3})},$
$y_{2}=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{2}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{3})}+\lambda^{2}\mathrm{e}^{-2\pi i(x_{1}-x_{2}+x_{3})}$
$\,+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{3})}+\mathrm{e}^{2\pi i(x_{1}-x_{2}+x_{3})},$
$y_{3}=$ $\,\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{2\pi ix_{3}}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
The metric on $V$ that is induced from $\mathbb{R}^{4}$ is given by
$\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix},\quad\left(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j}\right)={\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)}^{-1}=\frac{1}{4}\begin{pmatrix}3&2&1\\ 2&4&2\\ 1&2&3\end{pmatrix}.$
In the coordinates $y_{1},y_{2},y_{3}$ we have
$\left(g_{\lambda}^{ij}\right)$ $=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=\begin{pmatrix}-\frac{3}{4}y_{1}^{2}+2\lambda y_{2}&-\frac{1}{2}y_{1}y_{2}+3\lambda y_{3}&-\frac{1}{4}y_{1}y_{3}+4\lambda^{2}\\ -\frac{1}{2}y_{1}y_{2}+3\lambda y_{3}&-y_{2}^{2}+2y_{1}y_{3}+4\lambda^{2}&-\frac{1}{2}y_{2}y_{3}+3\lambda y_{1}\\ -\frac{1}{4}y_{1}y_{3}+4\lambda^{2}&-\frac{1}{2}y_{2}y_{3}+3\lambda y_{1}&-\frac{3}{4}y_{3}^{2}+2\lambda y_{2}\end{pmatrix}.$
We introduce the pencil generators
$z_{1}=y_{1}+4\lambda,\quad z_{2}=y_{2}-6\lambda,\quad z_{3}=y_{3}+4\lambda.$
In the new coordinates, we obtain a flat pencil of metrics $\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right)$ with
$\left(\eta^{ij}\right)=\begin{pmatrix}6z_{1}+2z_{2}&-3z_{1}+2z_{2}+3z_{3}&z_{1}+z_{3}\\ -3z_{1}+2z_{2}+3z_{3}&-8z_{1}-12z_{2}-8z_{3}&3z_{1}+2z_{2}-3z_{3}\\ z_{1}+z_{3}&3z_{1}+2z_{2}-3z_{3}&2z_{2}+6z_{3}\end{pmatrix}.$
The metric $\eta$ has the flat coordinates $t_{1},t_{2},t_{3}$ which are given by
$z_{1}$ $=\frac{1}{6}t_{1}^{4}+t_{1}^{2}t_{2}+\frac{1}{2}t_{2}^{2}+t_{1}t_{3},$
$z_{2}$ $=\frac{2}{3}t_{1}^{4}-t_{2}^{2}-2t_{1}t_{3},$
$z_{3}$ $=\frac{1}{6}t_{1}^{4}-t_{1}^{2}t_{2}+\frac{1}{2}t_{2}^{2}+t_{1}t_{3};$
in these coordinates the metrics $\eta$ has the form
$\left(\eta^{ij}\right)$ $=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{3},\omega_{1}+\omega_{3})$ with potential
$F=$ $-\frac{t_{1}^{8}}{18144}+\frac{t_{1}^{5}\ t_{3}}{720}-\frac{t_{2}^{6}}{96t_{1}^{4}}-\frac{t_{1}^{4}\ t_{2}^{2}}{288}+\frac{t_{2}^{4}t_{3}}{16\ t_{1}^{3}}-\frac{t_{2}^{2}t_{3}^{2}}{8t_{1}^{2}}$
$-\frac{t_{1}^{2}\ t_{3}^{2}}{24}-\frac{1}{24}t_{1}t_{2}^{2}\ t_{3}+\frac{t_{3}^{3}}{24t_{1}}-\frac{t_{2}^{4}}{96},$
the Euler vector field
$E=\frac{1}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{3}{4}t_{3}\frac{\partial}{\partial t_{3}},$
and the unit vector field
$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{3}t_{1}^{2}-36t_{2}^{2}t_{1}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{2}t_{3}t_{1}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right.$
$\,\left.+\left(-8t_{1}^{7}-60t_{3}t_{1}^{4}+120t_{2}^{2}t_{1}^{3}-72t_{3}^{2}t_{1}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$
where
$D=\left(t_{1}^{4}-6t_{2}t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{2}\ t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right).$
We also have $e^{\flat}=-(\mathrm{d}\log z_{1}+\mathrm{d}\log z_{3})$. The monodromy group is given by
$\mathrm{Mono}(\mathcal{M}(A_{3},\omega_{1}+\omega_{3}))=\mathrm{Stab}(\omega_{1}+\omega_{3})=\langle\sigma_{2}\rangle\rtimes\mathbb{Z}^{3}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
6. Conclusions
Given a root system $R$, a weight $\omega$ and a set of pencil generators $z^{1},\cdots,z^{\ell}$, we constructed a generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$. We proved that $z^{1},\cdots,z^{\ell}$ depend polynomially on the flat coordinates $t^{1},\cdots,t^{\ell}$ of the generalized Frobenius manifold, we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold can be embedded into the subgroup $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with an appropriately chosen weight $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail. For the case $(A_{\ell},\omega_{\ell})$, the generalized Frobenius manifold structure that we will present in [20] coincides with the one constructed in [7], where this generalized Frobenius manifold structure is constructed in terms of a super-potential given by a Fourier polynomial. It is shown in [7] that the bihamiltonian structure that is given by the associated flat pencil of metrics coincides with the dispersionless limit of the bihamiltonian structure of the q-deformed Gelfand-Dickey hierarchy [16, 17].
Finally, let us outline some potential directions for further research.
  1. i)
    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrary weight $\omega$?
  2. ii)
    Precise description of monodromy groups: While we have shown that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ embeds into $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$, can this subgroup be characterized more explicitly?
  3. iii)
    Relation to other Frobenius manifold constructions: How do these generalized Frobenius manifolds relate to earlier constructions on the orbit spaces of the finite and the extended affine Weyl groups?
  4. iv)
    Almost duality and geometric interpretations: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
  5. v)
    Legendre transformations: What are the properties of Legendre transformations [23, 27] for these generalized Frobenius manifolds?
  6. vi)
    Associated integrable systems: Which integrable hierarchies already known in the theory of solitons are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?

### v3

@@
 Abstract.
-We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are proper subgroups of the associated affine Weyl groups.
+We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are parabolic subgroups of the associated affine Weyl groups.
 Key words and phrases:
-Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group
+Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group.
 1. Introduction
-The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study of the relation of integrable systems and 2-dimensional topological field theories [8, 9].
+The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study on the relation of integrable systems and 2-dimensional topological field theories [8, 9].
 This geometric structure and its generalizations play a central role in topological field theory, Gromov-Witten theory, singularity theory and in the theory of integrable systems (see [1, 2, 3, 4, 5, 9, 13, 15, 18, 19, 22, 24] and the references therein).
-We consider in this paper the construction of a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups.
-These structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector field.
+In this paper, we construct on the orbit spaces of affine Weyl groups a class of generalized Frobenius manifold structures.
+These generalized structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector fields.
 The simplest example in this class corresponds to the affine Weyl group of type $A_{1}$, which naturally appears in the study of Hodge integrals and their remarkable connections to two important discrete integrable systems: the Volterra hierarchy and the $q$-deformed KdV hierarchy — both fundamental in soliton theory[21, 22].
 Our approach parallels the construction of a Frobenius manifold structure on the orbit space of a Coxeter group or on a certain extension of an affine Weyl group [9, 12, 14, 25, 26, 28, 29].
-For these cases, a flat pencil of metrics, based on which the Frobenius manifold structure is constructed, is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space.
+For these cases, a flat pencil of metrics is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space, based on which the Frobenius manifold structure is constructed.
 The key to these constructions lies in the linear dependence of both the contravariant components of the metric and its Levi-Civita connection on a distinguished coordinate of the orbit space — this special coordinate naturally gives rise to the flat unit vector field of the Frobenius manifold.
@@
 To overcome this, we proceed as follows: Given an affine Weyl group $W_{a}$ acting on an $\ell$-dimensional Euclidean space $V$, we introduce a parameter $\lambda$ into the invariant Fourier polynomials, yielding the coordinates $z^{1},\dots,z^{\ell}$ on the orbit space.
-The projected contravariant metric $\left(g^{\alpha\beta}(z)\right)$ then depends on $\lambda$.
-The key idea is to choose the $\lambda$-dependence of the invariant Fourier polynomials such that $\left(g^{\alpha\beta}(z)\right)$ varies linearly with $\lambda$.
-In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds (including their monodromy groups), and provide some explicit examples.
-Further examples will appear in a subsequent paper [20].
+The projected contravariant metric $\left(g_{\lambda}^{\alpha\beta}(z)\right)$ then depends on $\lambda$.
+Our approach is to choose the invariant $\lambda$-Fourier polynomials such that $\left(g_{\lambda}^{\alpha\beta}(z)\right)$ and the Christoffel symbols of their Levi-Civita connections depend linearly on $\lambda$.
+We prove in this paper that as long as this linear dependence holds, there exists a corresponding generalized Frobenius manifold structure on a dense open subset of the orbit space.
+In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds including their monodromy groups, and provide some explicit examples.
+Further examples will appear in our subsequent paper [20].
 The paper is organized as follows.
@@
 2.1. Affine coordinates
-[…]
+Let $R$ be an irreducible reduced root system in an $\ell$-dimensional Euclidean space $V$ with inner product (,)($\cdot\,$$\cdot$), and $\alpha_{1},\dots,\alpha_{\ell}$ be a fixed basis of simple roots.
+Denote by $\alpha_{1}^{\vee},\dots,\alpha_{\ell}^{\vee}$ the corresponding coroots given by
 $\alpha_{j}^{\vee}=\frac{2\alpha_{j}}{(\alpha_{j},\alpha_{j})},\quad j=1,\dots,\ell,$
@@
 $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee},$ (2.2)
-where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of elements of $V$:
+where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of vectors in $V$:
 $\mathbf{x}=c\omega+\hat{\mathbf{x}}=(x^{1}+\theta_{1}c)\alpha_{1}^{\vee}+\dots+(x^{\ell}+\theta_{\ell}c)\alpha_{\ell}^{\vee},$ (2.3)
@@
 Denote by $W_{a}(R)$ the affine Weyl group of $R$. It acts on $V$ by affine transformations
-$\mathbf{x}\mapsto\sigma(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\sigma\in W,\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
+$\phi\colon\mathbf{x}\mapsto\phi_{0}(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\phi_{0}\in W(R),\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
 where $W=W(R)$ is the Weyl group of $R$ generated by the reflections
 $\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha_{j}^{\vee},\mathbf{x})\alpha_{j},\quad\mathbf{x}\in V,\ j=1,\dots,\ell.$
-The action
-$\sigma\left(c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}\right)=c\omega+\tilde{x}^{1}\alpha_{1}^{\vee}+\dots+\tilde{x}^{\ell}\alpha_{\ell}^{\vee}$ (2.6)
-of $\sigma\in W_{a}(R)$ on $V$ yields the change of affine coordinates
-$\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c).$
-2.2. The Invariant $\lambda$-Fourier polynomial ring and its generators
-Let us consider the Fourier function ring
-$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{m(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}\right\}.$
-We introduce a gradation on $\mathscr{F}$ by the assignment
+2.2. The invariant $\lambda$-Fourier polynomial ring and basic generators
+In this subsection, we are to introduce the $\lambda$-Fourier polynomial ring $\mathscr{A}$ and its invariant subring $\mathscr{A}^{W}$ w.r.t.
+the action of the affine Weyl group $W_{a}(R)$, and define the set of basic generators $\{y^{1},\dots,y^{\ell}\}$.
+To this end, we first introduce a bigger ring $\mathscr{F}$, called the Fourier function ring, as follows:
+$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{f(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{C}\right\}.$
+For any $f\in\mathscr{F}$, we can uniquely write it in the form $f(\mathbf{x};c)$. Then we have a $W_{a}(R)$-action on $\mathscr{F}$ given by
+$(\phi^{*}f)(\mathbf{x};c)=f(\phi(\mathbf{x});c),\quad\phi\in W_{a}(R).$ (2.6)
+Note that if we write $f$ as a $f(\hat{\mathbf{x}};c)$, then the action is given by
+$(\phi^{*}f)(\hat{\mathbf{x}};c)=(\phi^{*}f)(\mathbf{x}-c\omega;c)=f(\phi(\mathbf{x})-c\omega;c),\quad\phi\in W_{a}(R),$ (2.7)
+which in general does not coincide with $f(\phi(\hat{\mathbf{x}});c)$. We introduce a $\mathbb{C}$-gradation on $\mathscr{F}$ by the assignment
 $\deg\mathrm{e}^{2\pi ix^{j}}=\theta_{j},\quad\deg\mathrm{e}^{2\pi ic}=-1,$ (2.8)
-where $\theta_{j}$ is defined in (2.4). Then for any $t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}$ and $\zeta\in\mathbb{C}$ we have
-$\deg\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}+\zeta)}=-t_{0}+t_{1}\theta_{1}+\dots+t_{\ell}\theta_{\ell}.$ (2.8)
-In other words, for any given linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and constant $m_{0}\in\mathbb{C}^{\times}$ we have
-$\deg\left(m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}\right)=\psi(\theta_{1},\dots,\theta_{\ell};-1).$ (2.9)
-The action of $W_{a}(R)$ on the affine coordinates of $V$ that is defined by (2.6) leads to a right action of $W_{a}(R)$ on $\mathscr{F}$, which is given by
-$\sigma:~{}\mathscr{F}\to\mathscr{F},~{}f(x^{1},\dots,x^{\ell};c)\mapsto(f\circ\sigma)(x^{1},\dots,x^{\ell};c)$ (2.10)
-for any given $\sigma\in W_{a}(R)$.
+where $\theta_{j}$ is defined in (2.4). Then for any given linear function
+$\psi(\mathbf{x};c)=t_{1}(x^{1}+\theta_{1}c)+\dots+t_{\ell}(x^{\ell}+\theta_{\ell}c)+t_{0}c,\quad t_{1},\dots,t_{\ell}\in\mathbb{C}$ (2.9)
+and constant $A_{0}\in\mathbb{C}^{\times}$, we have
+$\deg\left(A_{0}~\mathrm{e}^{2\pi i\psi(\mathbf{x};c)}\right)=\psi(\mathbf{0};-1).$ (2.10)
+Here $\mathbf{0}$ is the zero vector in $V$.
 Lemma 2.1.
-The affine Weyl group $W_{a}(R)$ acts homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
+The affine Weyl group $W_{a}(R)$ acts quasi-homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
 Proof.
-Let $m(x^{1},\dots,x^{\ell};c)$ be a nonzero monimial, which can be represented in the form
-$m(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$
-with a linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and a nonzero constant $m_{0}$. For any given translation
-$\sigma:\mathbf{x}\mapsto\mathbf{x}+\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
-of $W_{a}(R)$, its action on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is given by
-$(x^{1},\dots,x^{\ell};c)\mapsto(x^{1}+n_{1},\dots,x^{\ell}+n_{\ell};c),$
-so the associated action on $m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$ yields $m_{0}^{\prime}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$, which preserves the linear function $\psi(x^{1},\dots,x^{\ell};c)$ and only changes $m_{0}$ to another constant
-$m_{0}^{\prime}=m_{0}\mathrm{e}^{2\pi i(t_{1}n_{1}+\dots+t_{\ell}n_{\ell})}.$
-In particular, the action by $\sigma$ preserves the degree of $m(x^{1},\dots,x^{\ell};c)$.
-On the other hand, the action of any given $\sigma\in W(R)$ on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is linear, hence
-$(\sigma m)(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i(\psi\circ\sigma)(x^{1},\dots,x^{\ell};c)}.$
-From the relations
-$\sigma(\mathbf{0})=\mathbf{0},\quad(-1)\omega+\theta_{1}\alpha_{1}^{\vee}+\dots+\theta_{\ell}\alpha_{\ell}^{\vee}=\mathbf{0},$
-it follows that $\sigma$ preserves $(\theta_{1},\dots,\theta_{\ell};-1)$. Therefore by using (2.9) we arrive at
-$\deg(\sigma m)=(\psi\circ\sigma)(\theta_{1},\dots,\theta_{\ell};-1)=\psi(\theta_{1},\dots,\theta_{\ell};-1)=\deg m.$
-The lemma is proved. ∎
-Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. Let
-$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
-where $N_{j}=\#\{\sigma\in W\mid\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
+Let $f(\mathbf{x};c)\in\mathscr{F}$ be a nonzero monomial, which can be represented in the form
+$f(\mathbf{x};c)=A_{0}\mathrm{e}^{2\pi i\psi(\mathbf{x};c)}$
+with a linear function $\psi$ of the form (2.9) and a nonzero constant $A_{0}$. On one hand, for any given translation
+$\phi\colon\mathbf{x}\mapsto\mathbf{x}+\alpha,\quad\alpha=\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
+of $W_{a}(R)$, its action on $f$ is given by
+$(\phi^{*}f)(\mathbf{x};c)=A_{0}\mathrm{e}^{2\pi i\psi(\mathbf{x}+\alpha;c)}=\mathrm{e}^{2\pi i\psi(\alpha;0)}f(\mathbf{x};c).$
+Therefore, $\phi^{*}f$ and $f$ merely differ by a nonzero factor, hence the action by $\phi$ preserves the degree.
+On the other hand, for any $\phi\in W(R)$ is given by
+$\deg(\phi^{*}f)=\deg\left(A_{0}~\mathrm{e}^{2\pi i\psi(\phi(\mathbf{x});c)}\right)=\psi(\phi(\mathbf{0});c).$
+Since $\phi$ is linear, we have $\phi(\mathbf{0})=\mathbf{0}$, hence $\deg(\phi^{*}f)=\deg f$. The lemma is proved. ∎
+Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. We introduce the invariant functions $Y_{1},\dots,Y_{\ell}\in\mathscr{F}^{W}$ by
+$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
+where $N_{j}=\#\{\sigma\in W(R)\colon\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
 Proposition 2.2.
-The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are homogeneous of degree 00.
+The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are quasi-homogeneous of degree 0.
 Proof.
@@
 thus we arrive at $\deg Y_{j}=0$. The proposition is proved. ∎
-Let us proceed to consider generators of the $W_{a}(R)$-invariant ring $\mathscr{F}^{W}$.
-Since $\mathscr{F}^{W}$ is too large to consider its generators, we are to introduce a smaller subring of $\mathscr{F}$ which is named the $\lambda$-Fourier polynomial ring.
-Denote
+Let us proceed to introduce the $\lambda$-Fourier polynomial ring $\mathscr{A}$ and its invariant subring $\mathscr{A}^{W}$ w.r.t. the action of $W_{a}(R)$. Denote
 $\kappa=\gcd\{(\omega,\alpha_{r})~|~r=1,\dots,\ell\},$ (2.12)
@@
 $\deg\lambda=\kappa.$ (2.14)
-In particular, if we take $\omega=\omega_{k}$, then $\kappa=(\omega_{k},\alpha_{k})$ and $\lambda=\mathrm{e}^{-2\pi ic(\omega_{k},\alpha_{k})}$.
 Define the $\lambda$-Fourier polynomial ring $\mathscr{A}$ as a subring of $\mathscr{F}$ by
@@
 $\mathrm{e}^{2\pi i\left(-a_{0}\kappa c+a_{1}x^{1}+\dots+a_{\ell}x^{\ell}\right)},\quad a_{0}\in\mathbb{Z}_{\geq 0},a_{1},\dots,a_{\ell}\in\mathbb{Z}$
-and their linear combinations.
-We denote by $\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$ the $W_{a}(R)$-invariant $\lambda$-Fourier polynomial ring.
+and their linear combinations. We call the subring of $\mathscr{A}$ defined by
+$\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$
+the invariant $\lambda$-Fourier polynomial ring.
 Then we have $\lambda=\mathrm{e}^{-2\pi i\kappa c}\in\mathscr{A}^{W}$.
+Note that $W_{a}(R)$ does not act on $\mathscr{A}$, since results of the action may involve negative power terms of $\lambda$, which are not in $\mathscr{A}$.
 Let
-$y^{j}(\mathbf{x})=\mathrm{e}^{-2\pi i\theta_{j}c}~{}Y_{j}(\mathbf{x}),~{}j=1,\dots,\ell.$ (2.15)
-From Proposition 2.2 it follows that $y^{j}$ is homogeneous of degree $\theta_{j}$.
+$y^{j}(\mathbf{x};c)=\mathrm{e}^{-2\pi i\theta_{j}c}~Y_{j}(\mathbf{x}),~j=1,\dots,\ell.$ (2.15)
+It follows from Proposition 2.2 that $y^{j}$ is quasi-homogeneous of degree $\theta_{j}$.
 Lemma 2.3.
@@
 Observe that
-$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~{}=$ $~{}-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{*}(\omega_{j}),\mathbf{x})$
-$=$ $~{}-2\pi ic(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
-Since $\sigma^{*}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
-$\sigma^{*}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
+$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~=$ $~-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{-1}(\omega_{j}),\mathbf{x})$
+$=$ $~-2\pi ic(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{-1}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
+Since $\sigma^{-1}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
+$\sigma^{-1}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
 from which it follows that
-$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
+$(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
 where $m$ is a non-negative integer. We also have
-$\left(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
+$\left(\sigma^{-1}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
 where $c_{sr}$ are the $(s,r)$-entries of the Cartan matrix of the root system $R$. Therefore, there exists $a_{0}\in\mathbb{Z}_{\geq 0}$ and $a_{1},\dots a_{\ell}\in\mathbb{Z}$ such that
@@
 The lemma is proved. ∎
+Theorem 2.4.
+We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
+A proof of this theorem will be given in the next subsection. We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
 Example 2.1 ($(R,\omega)=(A_{2},\omega_{2})$).
-We consider the case where the root system $R=A_{2}$ and the weight $\omega=\omega_{2}$. We take the simple roots
-$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}$
-of the root system $A_{2}$ in the hyperplane $V$ of the Euclidean space $\mathbb{R}^{3}$ spanned by $\alpha_{1},\alpha_{2}$. We have
+Let $e_{1},e_{2},e_{3}$ be an the orthonormal basis of the Euclidean space $\mathbb{R}^{3}$, and $V$ be the hyperplane spanned by
+$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}.$
+Then we can take $\alpha_{1},\alpha_{2}$ as a basis of simple roots for the root system $R$ of type $A_{2}$ in $V$. We have
 $\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
-The affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
+Let us take $\omega=\omega_{2}$, then the affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
 $\mathbf{x}=c\omega_{2}+x^{1}\alpha^{\vee}_{1}+x^{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
-The Weyl group $W$ acting on $V$ is generated by the reflections
-$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
 From the definition (2.11) of $Y_{j}$, we have
@@
 $Y_{2}$ $=\mathrm{e}^{2\pi i\left(x^{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x^{1}-x^{2}-\frac{1}{3}c\right)}.$
-Here $\kappa=1$, hence
-$\lambda=\mathrm{e}^{-2\pi ic}.$
-Then we have
+Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$, and the basic generators are given by
 $y^{1}$ $=e^{-\frac{2}{3}\pi ic}Y_{1}=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})}+\lambda\mathrm{e}^{-2\pi ix^{2}},$
 $y^{2}$ $=e^{-\frac{4}{3}\pi ic}Y_{2}=\mathrm{e}^{2\pi ix^{2}}+\lambda\mathrm{e}^{-2\pi ix^{1}}+\lambda\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
-In the next subsection we are to show that $y^{1},\dots,y^{\ell}$ and $\lambda$ generate $\mathscr{A}^{W}$.
-2.3. The $\varphi_{\lambda}$-transformation and the null transformation
+2.3. The $\varphi_{\lambda}$-transformation and the null-transformation
 The parameter $\lambda$ of the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ is defined by (2.13), so it lies on the unit circle.
-However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$.
+However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$ in particular.
 Definition 2.5.
-For any fixed $\lambda\in\mathbb{C}$, the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ yield a transformation
+For any fixed $\lambda\in\mathbb{C}$, the basic generators $y^{1},\dots,y^{\ell}$ yield a transformation
 $\varphi_{\lambda}\colon(x^{1},\dots,x^{\ell})\mapsto(y^{1},\dots,y^{\ell}).$
-We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null transformation.
+We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null-transformation.
 Example 2.2 ($(R,\omega)=(A_{2},\omega_{2})$).
@@
 $y^{2}\big|_{\lambda=1}$ $=\mathrm{e}^{2\pi ix_{2}}+\mathrm{e}^{-2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
-The corresponding null transformation is given by
+The corresponding null-transformation is given by
 $y^{1}\big|_{\lambda=0}$ $=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})},$
@@
 Proposition 2.6.
-The Jacobian determinant of the null transformation is given by
-$\mathbf{J}_{0}(\hat{\mathbf{x}})=e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+},\,(\beta,\omega)=0}\left(1-\mathrm{e}^{-2\pi i(\beta,\hat{\mathbf{x}})}\right),$ (2.21)
-where $\Phi^{+}$ is the set of all positive roots of $R$.
+The Jacobian determinant of the null-transformation is given by
+[…]
+where $R^{+}$ is the set of all positive roots of $R$ and $C$ is a nonzero constant.
 Proof.
 From [6] and [12] we know that the Jacobian determinant of $\varphi_{\lambda}$ can be represented in the form
-$\mathbf{J}_{\lambda}(\hat{\mathbf{x}})~{}=$ $~{}\mathrm{e}^{-2\pi ic\sum_{j}(\omega_{j},\omega)+\pi i\sum_{\alpha\in\Phi^{+}}\left(c(\alpha,\omega)+(\alpha,\hat{\mathbf{x}})\right)}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right)$
-$=$ $~{}e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right).$ (2.22)
+[…]
+[…]
 Here we use the fact that [6]
-$\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
-Since $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is a polynomial of $\lambda$, its constant term w.r.t. $\lambda$ yields (2.21). The proposition is proved. ∎
+$\frac{1}{2}\sum_{\alpha\in R^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
+Since $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is a polynomial of $\lambda$, we have $\mathbf{J}_{0}(\hat{\mathbf{x}})=\mathbf{J}(\hat{\mathbf{x}};0)$, so the constant term of $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ w.r.t.
+$\lambda$ yields (2.21).
+The proposition is proved.
+∎
 Corollary 2.7.
-The Jacobian determinant of the null transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor.
+The Jacobian determinant of the null-transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor.
 In particular, $y^{1}\big|_{\lambda=0},\dots,y^{\ell}\big|_{\lambda=0}$ are functionally independent.
-Theorem 2.6.
-We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
-Proof.
+We now turn to the proof of Theorem 2.4.
+Proof of Theorem 2.4.
 For any given $f\in\mathscr{A}^{W}$, we need to show that it can be represented in terms of a polynomial of $y^{1},\dots,y^{\ell}$ and $\lambda$.
-By using Lemma 2.1, we know that $f$ can be decomposed as the sum of homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is homogeneous of degree $n$.
-From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big{|}_{\lambda=1}$,…,$y^{\ell}\big{|}_{\lambda=1}$ as follows:
+By using Lemma 2.1, we know that $f$ can be decomposed as the sum of quasi-homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is quasi-homogeneous of degree $n$.
+From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big|_{\lambda=1}$$y^{\ell}\big|_{\lambda=1}$ as follows:
 $f|_{\lambda=1}=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}(y^{1}\big|_{\lambda=1})^{\beta_{1}}\dots(y^{\ell}\big|_{\lambda=1})^{\beta_{\ell}},$ (2.24)
-so it follows from the homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
+so it follows from the quasi-homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
 $f=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}\lambda^{n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}}(y^{1})^{\beta_{1}}\cdots(y^{\ell})^{\beta_{\ell}},$ (2.25)
-here $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
+where each $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
 $f=\sum_{j\in Z}\lambda^{j}P_{j}(y^{1},\dots,y^{\ell}),$ (2.26)
@@
 This leads to a contradiction with the fact, as shown by Corollary 2.7, that $y^{1}\big|_{\lambda=0},\dots,y^{\ell}\big|_{\lambda=0}$ are functionally independent.
-Thus $q$ must be a non-negative integer, and the proposition is proved.
+Thus $q$ must be a non-negative integer, and the theorem is proved.
 ∎
-Definition 2.2.
-We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
 2.4. Leading terms of the basic generators
@@
 which we call the leading term of $y^{j}$.
-This term also appears in $y^{j}|_{\lambda=0}$, and the following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ only contain their leading terms.
+Since not involving $\lambda$, this term also appears in $y^{j}|_{\lambda=0}$.
+The following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ consist merely of their leading terms.
 Theorem 2.8.
-Let us denote
-$S=\{r\in\{1,\dots,\ell\}\mid m_{r}=(\omega,\alpha_{r}^{\vee})>0\}.$ (2.29)
+Denote
+$S=\{r\in\{1,\dots,\ell\}\colon m_{r}=(\omega,\alpha_{r}^{\vee})>0\},$ (2.29)
 then we have
@@
 For any given $j\in S$, from (2.11) and (2.15) we obtain
-$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ (\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (2.31)
+$y^{j}\big|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ (\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$.
+(2.31)
 Assume that $\sigma\in W$ satisfies the condition
-$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0,$ (2.32)
+$(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=0,$ (2.32)
 then from (2.18) it follows that
@@
 By using the fact that all terms on the left hand side of the above equality are nonnegative and $m_{j}>0$, we obtain $n_{j}=0$, and
-$(\sigma^{*}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
-Since $\sigma^{*}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{*}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
-$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
+$(\sigma^{-1}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
+Since $\sigma^{-1}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{-1}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
+$y^{j}\big|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ \omega_{j}=\sigma^{-1}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W(R)\\ \omega_{j}=\sigma^{-1}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
 The theorem is proved. ∎
@@
 The following identity holds true:
-$\sum_{r=1}^{\ell}m_{r}\log\left(y^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.35)
-2.5. Proper generators
-In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection some other generators of $\mathscr{A}^{W}$.
+[…]
+2.5. Proper factors and proper generators
+In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection the notion proper generators and proper factors of $\mathscr{A}^{W}$.
+Consider the subalgebra $\mathscr{R}=\mathbb{C}[y^{1},\dots,y^{\ell}]$ of $\mathscr{A}^{W}$. It has the following two properties:
+  • $\mathscr{R}$ is a graded subring of $\mathscr{A}$.
+This means that for any $f\in\mathscr{R}$, if we decompose it into quasi-homogeneous components as $f_{1}+\cdots+f_{n}$, then each $f_{k}\in\mathscr{R}$.
+This property follows from the quasi-homogeneity of $y^{1},\dots,y^{\ell}$.
+  • The tensor product decomposition $\mathscr{A}^{W}=\mathscr{R}\otimes\mathbb{C}[\lambda]$ holds. This means that the natural homomorphism
+    $h\colon\mathscr{R}\otimes\mathbb{C}[\lambda]\to\mathscr{A}^{W},\quad f\otimes g\mapsto fg,\quad f\in\mathscr{R},\,g\in\mathbb{C}[\lambda]$
+    is an isomorphism.
+These properties give rise to the following definition.
+Definition 2.10.
+A subalgebra $\mathscr{P}$ of $\mathscr{A}^{W}$ is called a proper factor, if it is graded and $\mathscr{A}^{W}=\mathscr{P}\otimes\mathbb{C}[\lambda]$.
+In what follows, we are to find a set of generators for each proper factor $\mathscr{P}$, called proper generators.
 Definition 2.11.
@@
 Let $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$, then
-$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell},\lambda].$ (2.38)
+$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda].$ (2.38)
 Proof.
-It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for $j=1,\dots,\ell$.
-Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, and let $y^{k}$ be one of those with the lowest degree.
-It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$.
-However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}\not\in C[z^{1},\dots,z^{\ell},\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$.
-Then we know from Lemma 2.9 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, which leads to a contradiction.
+It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$ for $j=1,\dots,\ell$.
+Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, and let $y^{k}$ be one of those with the lowest degree.
+It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$.
+However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}$ that do not belong to $C[z^{1},\dots,z^{\ell};\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$.
+Then we know from Lemma 2.12 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, which leads to a contradiction.
 The proposition is proved.
@@
 For any set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we have
-$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}_{\lambda}(\hat{\mathbf{x}}),$ (2.39)
-where $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is defined in (2.22), thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
+$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}(\hat{\mathbf{x}};\lambda),$ (2.39)
+where $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is defined by (2.22). Thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
 Proof.
-For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ satisfying the condition $\theta_{r}<\theta_{j}$. Then from the relation
+For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ that satisfy the condition $\theta_{r}<\theta_{j}$.
+Thus from the relations
 $\frac{\partial z^{j}}{\partial x^{p}}=\frac{\partial y^{j}}{\partial x^{p}}+\lambda\sum_{\theta_{r}<\theta_{j}}\frac{\partial s^{j}}{\partial y^{r}}\frac{\partial y^{r}}{\partial x^{p}}$ (2.40)
 we arrive at (2.39). The corollary is proved. ∎
-For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we will also call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null transformation.
-Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation, i.e.,
-$z^{j}\big{|}_{\lambda=0}=y^{j}\big{|}_{\lambda=0},\quad j=1,\dots,\ell.$ (2.41)
+Suppose $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators, then $\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$ is a proper factor.
+In fact, it follows from the corollary above that $z^{1},\dots,z^{\ell}$ and $\lambda$ are algebraically independent (as functions of $x^{1},\dots,x^{\ell}$ and $\lambda$), hence $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell}]\otimes\mathbb{C}[\lambda]$.
+Since $z^{1},\dots,z^{\ell}$ are quasi-homogeneous, the subalgebra $\mathbb{C}[z^{1},\dots,z^{\ell}]$ is graded.
+The following theorem shows that all proper factors have such a form.
+Furthermore, each proper factor is generated by a unique set of proper generators.
+Theorem 2.15.
+Suppose $\mathscr{P}$ is a proper factor of $\mathscr{A}^{W}$, then there exists a unique set of proper generator $\{z^{1},\dots,z^{\ell}\}$ such that $\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$.
+Proof.
+By definition, we have $\mathscr{A}^{W}=\mathscr{P}\otimes\mathbb{C}[\lambda]$, hence each $y^{i}\in\mathscr{A}^{W}$ can be represented in the form
+$y^{i}=\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(\lambda),\quad p^{i}_{\alpha}\in\mathscr{P},\ T^{i}_{\alpha}(\lambda)\in\mathbb{C}[\lambda].$ (2.41)
+In particular, we have
+$y^{i}\big|_{\lambda=0}=\Bigl(\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)\Bigr)\Bigl|_{\lambda=0},$
+where $\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)\in\mathscr{P}$.
+So we define $z^{i}$ to be the $\theta_{i}$-degree component of $\sum_{\alpha}p^{i}_{\alpha}T^{i}_{\alpha}(0)$, then $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators.
+Since $\mathscr{P}$ is graded, we have $z^{i}\in\mathscr{P}$, thus $\mathbb{C}[z^{1},\dots,z^{\ell}]\subset\mathscr{P}$.
+However, it follows from the properties of proper generators that $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell}]\otimes\mathbb{C}[\lambda]$.
+Since $\mathbb{C}[\lambda]$ is faithfully flat, we have
+$\mathbb{C}[z^{1},\dots,z^{\ell}]=\mathscr{P}.$
+Now, let us prove the uniqueness of the set of proper generators.
+Suppose there is another set of proper generators $\{\tilde{z}^{1},\dots,\tilde{z}^{\ell}\}$ such that $\mathscr{P}=\mathbb{C}[\tilde{z}^{1},\dots,\tilde{z}^{\ell}$], then for each $i=1,\dots,\ell$ we have
+$z^{i}-\tilde{z}^{i}=\lambda(s^{i}-\tilde{s}^{i}).$ (2.42)
+Here we write $\tilde{z}^{i}=y^{i}+\lambda\tilde{s}^{i}$.
+Since $z^{i}-\tilde{z}^{i}\in\mathscr{P}=\mathbb{C}[z^{1},\dots,z^{\ell}]$ and $(s^{i}-\tilde{s}^{i})\in\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, if $s^{i}-\tilde{s}^{i}$ do not vanish, then (2.42) gives a nontrivial algebraic relation of $z^{1},\dots,z^{\ell}$ and $\lambda$, which contradicts with their algebraic independence.
+Therefore, $z^{i}=\tilde{z}^{i}$ for each $i=1,\dots,\ell$.
+∎
+The above theorem shows that the map
+$\{\text{proper~generators}\}\to\{\text{proper~factors}\},\quad\{z^{1},\dots,z^{\ell}\}\mapsto\mathbb{C}[z^{1},\dots,z^{\ell}].$ (2.43)
+is a bijection. As a result, we can characterize a proper factor by its proper generators.
+For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we still call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null-transformation.
+Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation.
 In particular, for any element $r$ of the set $S$ defined by (2.29) we have
@@
 Hence from Corollary 2.35 it follows that
-$\sum_{r=1}^{\ell}m_{r}\log\left(z^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.43)
+[…]
 In the next section, we will impose an additional condition on proper generators of $\mathscr{A}^{W}$ in order to construct generalized Frobenius manifold structures on the orbits space of the affine Weyl group $W_{a}(R)$.
 3. The construction of generalized Frobenius manifold structures
-In this section, we are to construct a generalized Frobenius manifold structure starting from a given irreducible reduced root system $R$ in the $l$-dimensional Euclidean space $V$, a fixed weight $\omega$ of the form (2.1), and a set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$ satisfying a certain additional condition.
+For a given irreducible reduced root system $R$ in an $\ell$-dimensional Euclidean space $V$, let us fix a weight $\omega$ of the form (2.1), a proper factor $\mathscr{P}$ of $\mathscr{A}^{W}$ with the corresponding set $\{z^{1},\dots,z^{\ell}\}$ of proper generators.
+We are to impose a certain additional condition on these proper generators, so that we construct a generalized Frobenius manifold structure on a dense open subset of the spectrum of $\mathscr{P}$.
 3.1. The orbit space and pencil generators
-By definition, $\mathbb{C}[z^{1},\cdots,z^{\ell}]$ is a subalgebra of $\mathscr{A}^{W}$.
-Denote the spectrum of this subalgebra by $\mathcal{M}$, then $\mathcal{M}\cong\mathbb{C}^{\ell}$.
-Taking $x^{1},\cdots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
+Consider the manifold $\mathcal{M}=\mathbb{C}^{\ell}$ on which $z^{1},\dots,z^{\ell}$ are global coordinates.
+Taking $x^{1},\dots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
 $\varphi_{\lambda}\colon V\otimes\mathbb{C}\to\mathcal{M},\quad(x^{1},\dots,x^{\ell})\mapsto(z^{1},\dots,z^{\ell}),\quad\lambda\in\mathbb{C}.$ (3.1)
-Here the $\varphi_{\lambda}$-transformation is defined in the last section.
+Here the $\varphi_{\lambda}$-transformation is defined in the above section.
 We call $\mathcal{M}$ the orbit space of the affine Weyl group $W_{a}(R)$ w.r.t.
 the proper generators $z^{1},\dots,z^{\ell}$.
-In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to endow $\mathcal{M}$ with a structure of generalized Frobenius manifold.
-[…]
-$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\ \mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
-on $V\otimes\mathbb{C}$. Here by a contravariant metric we mean a symmetric non-degenerate bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$.
+In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to introduce a structure of generalized Frobenius manifold on a dense open subset of $\mathcal{M}$.
+The inner product (,)($\cdot\,$$\cdot$) of the Euclidean space $V$ yields a contravariant metric
+$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\quad\mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
+on $V\otimes\mathbb{C}$.
+By a contravariant metric, we mean a non-degenerate, symmetric bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$.
+We will also call $a$ a metric on $V\otimes\mathbb{C}$.
 Proposition 3.1.
@@
 $g_{\lambda}^{ij}=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{r}}a^{rs}\frac{\partial z^{j}}{\partial x^{s}},\quad i,j=1,\dots,\ell.$ (3.5)
-Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s.
+Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s. We denote
+$\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j},\quad i,j,k=1,\dots,\ell,$ (3.6)
+and call them the Christoffel symbols of the Levi-Civita connection of the contravariant metric $g_{\lambda}$, or simply the Christoffel symbols of $g_{\lambda}$.
 Proposition 3.2.
-We have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$.
+For all indices $i,j,k$, we have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$ and $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}^{W}$.
 Proof.
-By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ is invariant w.r.t.
+We first consider the entries of the contravariant metric $g_{\lambda}$.
+By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ are invariant w.r.t.
 the action of $W_{a}(R)$.
@@
 Since $\sigma$ is an orthogonal transformation, we know that
-$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq}.$
-Hence $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$. The proposition is proved. ∎
-The above proposition shows in particular that $g^{ij}_{\lambda}$ are polynomials of $\lambda$.
-In order to construct generalized Frobenius manifold structures on the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.
+$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq},$
+hence we arrive at $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$.
+Now we consider the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ of the contravariant metric $g_{\lambda}$.
+Since $g_{\lambda}$ is invariant w.r.t.
+the action of $W_{a}(R)$, it suffices to show that $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}$.
+From Corollary 2.14 and Proposition 2.6, it follows that
+$\det\left(\frac{\partial z^{i}}{\partial x^{j}}\right)=\mathbf{J}(\hat{\mathbf{x}};\lambda)=C\cdot\mathrm{e}^{\pi i\sum_{\alpha\in R^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in R^{+}}\left(1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}\right),$
+where $\mathbf{x}=c\omega+\hat{\mathbf{x}}$. For the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$, the formula
+$\Gamma^{ij}_{g_{\lambda},k}=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{p}}\frac{\partial^{2}z^{j}}{\partial x^{q}\partial x^{r}}a^{pq}\frac{\partial x^{r}}{\partial z^{k}}$ (3.8)
+holds, hence we have
+$\Gamma^{ij}_{g_{\lambda},k}=\frac{P^{ij}_{k}(\hat{\mathbf{x}};\lambda)}{\mathbf{J}(\hat{\mathbf{x}};\lambda)},$
+where $P_{k}^{ij}\in\mathscr{A}$.
+Note that $\mathbf{J}(\hat{\mathbf{x}};\lambda)$ is anti-invariant and $\Gamma_{\lambda,k}^{ij}$ are invariant w.r.t.
+the action of the Weyl group $W(R)$, so the Fourier polynomials $P_{k}^{ij}(\hat{\mathbf{x}};\lambda)$ must be anti-invariant w.r.t.
+the action of $W(R)$.
+Note that for each $\beta\in R^{+}$, if $(\beta,\mathbf{x})\in\mathbb{Z}$,
+$\sigma_{\beta}(\mathbf{x})=\mathbf{x}-(\beta,\mathbf{x})\beta^{\vee}=\mathbf{x}+n_{1}\alpha_{1}^{\vee}+\cdots+n_{\ell}\alpha_{\ell}^{\vee},\quad n_{1},\dots,n_{\ell}\in\mathbb{Z}.$ (3.9)
+Here $\beta^{\vee}\in R^{\vee}$ is a $\mathbb{Z}$-combination of $\alpha_{1}^{\vee},\dots,\alpha_{\ell}^{\vee}$.
+It follows from the anti-invariance of $P_{k}^{ij}$ and (3.9) that, as long as $(\beta,\mathbf{x})\in\mathbb{Z}$,
+$P_{k}^{ij}(\mathbf{x})=-P_{k}^{ij}(\sigma_{\beta}(\mathbf{x}))=-P_{k}^{ij}(\mathbf{x}),$
+hence $P_{k}^{ij}$ vanishes along $(\beta,\mathbf{x})\in\mathbb{Z}$, all zeros of $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$.
+Since $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$ has no square factor in $\mathscr{A}$, it follows from Nullstellensatz’s Theorem that $P_{k}^{ij}$ is divisible by $1-\mathrm{e}^{-2\pi i\left(\beta,\mathbf{x}\right)}$.
+Therefore, it is also divisible by $\mathbf{J}(\hat{\mathbf{x}};\lambda)$, and we have $\Gamma^{ij}_{g_{\lambda},k}\in\mathscr{A}$.
+The proposition is proved.
+∎
+Remark 3.1.
+This proposition implies that the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ can be defined globally on $\mathcal{M}$, although $g_{\lambda}$ may degenerate at some points.
+The above proposition shows, in particular, that $g^{ij}_{\lambda}$ and $\Gamma_{g_{\lambda},k}^{ij}$ are polynomials of $\lambda$.
+In order to construct generalized Frobenius manifold structures on a dense open subset of the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.
 Definition 3.3.
 Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$.
-We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl{(}g^{ij}_{\lambda}\bigr{)}$ depends linearly on $\lambda$, i.e., $g^{ij}_{\lambda}$ can be represented in the form
+We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl(g^{ij}_{\lambda}\bigr)$ and its Christoffel symbols $\Gamma_{g_{\lambda},k}^{ij}$ depend at most linearly on $\lambda$, and the metric $\partial_{\lambda}g_{\lambda}^{ij}$ is non-degenerate at generic points.
+From the above definition, for a set of proper generators $\{z^{1},\dots,z^{\ell}\}$, $g^{ij}_{\lambda}$ can be represented in the form
 $g^{ij}_{\lambda}=g^{ij}+\lambda\eta^{ij},$
-where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the contravariant components $\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j}$ of the Levi-Civita connection of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
+where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
 $\Gamma^{ij}_{g_{\lambda},k}=\Gamma^{ij}_{g,k}+\lambda\Gamma^{ij}_{\eta,k}.$ (3.10)
-From the above definition we know that a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil[10] of metrics $g,\eta$.
-Note that the flat metric $g$ is independent of the choice of proper generators, since
+Therefore, a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil [10] of metrics $g,\eta$. Note that the flat metric $g$ is independent of the choice of proper generators, since
 $g=g_{0}=\frac{1}{4\pi^{2}}(\varphi_{0})_{*}a$ (3.11)
 is induced merely by the null-transformation which does not depend on the choice of proper generators.
-We will present examples of pencil generators in Sect.
-5 and in a subsequent paper [20].
-Unless otherwise specified, in what follows we always assume that $\{z^{1},\dots,z^{\ell}\}$ is a set of pencil generators of $\mathscr{A}^{W}$.
-We will end this section by proving that we can construct a generalized Frobenius manifold structure on $\mathcal{M}$ with the given flat pencil of metrics.
-3.2. Homogeneous flat coordinates
-We are to show in this subsection the existence of a system of homogeneous flat coordinates for the flat metric $\eta$.
+We will present examples of pencil generators in Sect. 5 and in a subsequent paper [20]. Hereafter, $z^{1},\dots,z^{\ell}$} denotes a set of pencil generators for $\mathscr{A}^{W}$.
+3.2. Quasi-homogeneous flat coordinates
+We are to show in this subsection the existence of a system of quasi-homogeneous flat coordinates for the flat metric $\eta$.
 Flat coordinates of $\eta$ satisfy the system of equations
 $\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\quad i,j=1,\dots,\ell.$ (3.12)
-Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of $\eta$’s Levi-Cività connection, they are given by
+Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $\eta$, they are given by
 $\Gamma_{\eta,ij}^{k}=\frac{1}{2}\eta^{kr}\left(\frac{\partial\eta_{ir}}{\partial z^{j}}+\frac{\partial\eta_{jr}}{\partial z^{i}}-\frac{\partial\eta_{ij}}{\partial z^{r}}\right).$ (3.13)
-From Definition 2.3 we know that
+From Definition 2.11 it follows that
 $\deg g^{ij}=\theta_{i}+\theta_{j},\quad\deg\eta^{ij}=\theta_{i}+\theta_{j}-\kappa,$ (3.14)
@@
 $\deg g_{ij}=-\theta_{i}-\theta_{j},\quad\deg\eta_{ij}=\kappa-\theta_{i}-\theta_{j},\quad\deg\Gamma_{\eta,ij}^{k}=\theta_{k}-\theta_{i}-\theta_{j}.$ (3.15)
-We are to use the above homogeneity property to prove the existence of homogeneous flat coordinates of $\eta$.
-To this end, let us first clarify the solution space of the system of equations (3.9) that we are considering.
-Let $\widehat{\mathcal{M}}$ be the universal covering of $\mathcal{M}\setminus D$ with
-$D=\{z\in\mathcal{M}~{}\big{|}~{}\det\eta(z)=0\}.$ (3.13)
+We are to use the above quasi-homogeneity property to prove the existence of quasi-homogeneous flat coordinates of $\eta$.
+To this end, let us first clarify the solution space of the system of equations (3.12).
+Let $\mathcal{W}$ be the universal covering space of $\mathcal{M}\setminus D_{\eta}$ with
+$D_{\eta}=\{z\in\mathcal{M}\mid\det\eta(z)=0\}.$ (3.16)
 We consider solutions of the system of equations (3.12) on the space
-$\mathscr{Q}(\widehat{\mathcal{M}})=\mathscr{O}(\widehat{\mathcal{M}})/\mathbb{C},$
-where $\mathscr{O}(\widehat{\mathcal{M}})$ refers to the holomorphic function ring on $\mathcal{M}$. Denote the solution space of (3.9) by
-$\mathcal{K}=\left\{t\in\mathscr{Q}(\widehat{\mathcal{M}})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
+$\mathscr{Q}(\mathcal{W})=\mathscr{O}(\mathcal{W})/\mathbb{C},$
+where $\mathscr{O}(\mathcal{W})$ refers to the holomorphic function ring on $\mathcal{W}$. Denote the solution space of (3.12) by
+$\mathcal{K}=\left\{t\in\mathscr{Q}(\mathcal{W})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
 It follows from the theory of differential equations that $\mathcal{K}$ is an $\ell$-dimensional space, since a solution is uniquely determined by the values of its $\ell$ partial derivatives at an arbitrary point.
@@
 $E=\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial}{\partial z^{r}}.$ (3.17)
-Then for any homogeneous function $f$ of degree $n$, the following identity holds true:
+Then for any quasi-homogeneous function $f$ of degree $n$, the following identity holds true:
 $\mathcal{L}_{E}f=nf,$ (3.18)
 here $\mathcal{L}_{E}$ is the Lie derivative along $E$.
-Hence homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees.
+Hence quasi-homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees.
 Now we consider the action of $\mathcal{L}_{E}$ on the solution space $\mathcal{K}$.
@@
 Hence if $t\in\ker\mathcal{L}_{ij}$, then $\mathcal{L}_{E}t\in\ker\mathcal{L}_{ij}$. The proposition is proved. ∎
-In what follows, we always assume that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable.
-This implies that $\eta$ admits $\ell$ homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$.
+In what follows, we assume for conciseness that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable.
+This assumption of diagonalization of the vector field $E$ implies that $\eta$ admits $\ell$ quasi-homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$.
 Let
@@
 $\deg\eta^{\alpha\beta}=d_{\alpha}+d_{\beta}-\kappa$ (3.23)
-it follows that if $d_{\alpha}+d_{\beta}\not=\kappa$ then $\eta^{\alpha\beta}$ must vanish.
+it follows that $\eta^{\alpha\beta}$ vanishes if $d_{\alpha}+d_{\beta}\not=\kappa$.
 Thus by performing, if necessary, a degree preserving linear transformation, we can choose the flat coordinates $t^{1},\dots,t^{n}$ of $\eta$ such that
@@
 Here * is an involution
-:$*\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
+$\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
 which satisfies the relation $d_{\alpha}+d_{\alpha^{*}}=\kappa$. Usually, we prefer to choose the flat coordinates so that $\alpha^{*}=\ell+1-\alpha$.
-We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$.
+We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ on each point of $\mathcal{M}\setminus D_{\eta}$, corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$.
 Since
 $\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}E=\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}\sum_{k=1}^{\ell}\theta_{k}z^{k}\frac{\partial}{\partial z^{k}}=\left(\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,ik}^{j}+\theta_{i}\delta^{j}_{i}\right)\frac{\partial}{\partial z^{j}},$ (3.25)
-the matrix $A$ has elements
+the matrix $A$ has entries
 $A_{j}^{i}=\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,jk}^{i}+\theta_{j}\delta^{i}_{j}.$ (3.26)
 It has eigenvectors $\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T}$ with eigenvalues $d_{\alpha}$, i.e.,
-$A\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}=d_{\alpha}\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}.$ (3.24)
-Remark 3.1.
-When we choose $\omega$ to be a fundamental weight, we conjecture that all $d_{\alpha}$ are positive numbers. In what follows, we will continue our construction by assuming this fact.
+$A\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}=d_{\alpha}\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}$.
+(3.27)
+Lemma 3.5.
+All $d_{\alpha}$ have positive real parts.
+The proof of this lemma will be given in Appendix A. In fact, we will show in Corollary 4.4 that all $d_{\alpha}$ are positive rational numbers.
 3.3. The Frobenius algebra
-Denote by $\Gamma_{\alpha\gamma}^{\beta}$ the Christoffel symbols of the Levi-Civita connection of $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$, and by
-$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta}$ (3.25)
-the contravariant components of the Levi-Civita connection of $g$.
-Then from (3.7) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the contravariant components of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$.
-We define an algebra structure on $T^{*}\mathcal{M}$ by
+Denote by
+$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta},\quad\alpha,\beta,\gamma=1,\dots,\ell$ (3.28)
+the Christoffel symbols of the Levi-Civita connection of the contravariant $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$.
+Then from (3.10) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$.
+Moreover, we know from Remark 3.1 that $\Gamma^{\alpha\beta}_{\gamma}$ are defined globally on $\mathcal{M}\setminus D_{\eta}$.
+We define an algebra structure on $T^{*}(\mathcal{M}\setminus D_{\eta})$ by
 $\mathrm{d}t^{\alpha}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\Gamma_{\gamma}^{\alpha\beta}\mathrm{d}t^{\gamma},\quad\alpha,\beta=1,\dots,\ell,$ (3.29)
-[…]
-$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}.$ (3.27)
+where $d_{\beta}$ is the degree of $t^{\beta}$.
+From Lemma 3.5, we know in particular that $d_{\beta}\not=0$.
+This algebra is called the dual Frobenius algebra on $T^{*}(\mathcal{M}\setminus D_{\eta})$.
+Denote
+$D_{0}=\{z\in\mathcal{M}\,|\,\exists\,i\in S,\,z^{i}=0\},\quad D=D_{\eta}\cup D_{0},$
+and $\mathcal{M}_{D}=\mathcal{M}\setminus D$, then $\mathcal{M}_{D}$ is a dense open subset of $\mathcal{M}$. Let us define a one-form
+$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}$ (3.30)
+on $\mathcal{M}_{D}$.
+We are to show in this subsection that the multiplication (3.29) endows $T^{*}\mathcal{M}_{D}$ a Frobenius algebra structure with the unity $\omega_{e}$ and the bilinear form $\langle\cdot\,$$\cdot\rangle$ given by the flat metric $\eta$.
 Proposition 3.6.
-The 1-form (3.27) is a left unity of the algebra structure defined by (3.26) on $T^{*}\mathcal{M}$.
+The one-form (3.30) is a left unity of the algebra structure defined by (3.29) on $T^{*}\mathcal{M}_{D}$.
 Proof.
-From the definition of the multiplication (3.26) it follows that
+As a closed condition, it suffices to prove this proposition on generic points of $\{\det g\not=0\}$ on $\mathcal{M}_{D}$. From the definition of the multiplication (3.29) it follows that
 $\omega_{e}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}~\nabla_{g}^{\omega_{e}}\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}.$ (3.31)
 Here $\nabla_{g}$ (resp. $\nabla^{g}$) is the contravariant (resp. covariant dual) Levi-Civita connection, and $g^{\sharp}$ is the canonical isomorphism
-$g^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto g(\omega,\cdot).$ (3.29)
+$g^{\sharp}\colon T^{*}\mathcal{M}_{D}\to T^{**}\mathcal{M}_{D}=T\mathcal{M}_{D},\quad\omega\mapsto g(\omega,\cdot).$ (3.32)
 So it suffices to show that
 $\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}=\mathrm{d}t^{\beta}.$ (3.33)
-Note that the null-transformation $\varphi_{0}$ transforms the metric $a$ defined in (3.2) to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$.
 Since the relation between the coordinates $z^{1},\dots,z^{\ell}$ and $t^{1},\dots,t^{\ell}$ is independent of $\lambda$, in the following proof of the validity of (3.33) we can assume that $z^{i}$ and $t^{\alpha}$ are related with $x^{1},\dots,x^{\ell}$ via the null-transformation $\varphi_{0}$.
+Note that the null-transformation $\varphi_{0}$ transforms the constant metric $a$ to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$.
 From (2.45) we obtain
-$\varphi_{0}^{*}\omega_{e}=-\frac{2}{\kappa}\pi i\mathrm{d}(\omega,\hat{\mathbf{x}}),$ (3.31)
-it follows that
+[…]
+so it follows that
 $\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}\bigr|_{\lambda=0}$ $=-2\pi i\frac{1}{d_{\beta}}\sum_{s=1}^{\ell}m_{s}\nabla^{g}_{g^{\sharp}\mathrm{d}x^{s}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\mathrm{d}x^{p}\right)$
-$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}(\mathrm{d}x^{r},\mathrm{d}x^{s})\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
+$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}a^{rs}\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
 $=\frac{1}{d_{\beta}}\left(\sum_{r=1}^{\ell}\frac{1}{2\pi i}\theta_{r}\frac{\partial}{\partial x^{r}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\right)\mathrm{d}x^{p}.$
@@
 $\eta^{\alpha\epsilon}\frac{\partial^{2}f^{\beta}}{\partial t^{\epsilon}\partial t^{\gamma}}=\Gamma^{\alpha\beta}_{\gamma},$ (3.40)
-where $\deg f^{\beta}=\kappa+d_{\beta}$
+where $\deg f^{\beta}=\kappa+d_{\beta}$ .
 Lemma 3.9.
 The Christoffel symbols satisfy the following relations:
-$\Gamma_{\alpha}^{\beta\gamma}=\frac{d_{\gamma}}{d_{\beta}+d_{\gamma}}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}}.$ (3.38)
+$\Gamma_{\gamma}^{\alpha\beta}=\frac{d_{\beta}}{d_{\alpha}+d_{\beta}}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}.$ (3.41)
 Proof.
-By multiplying $e_{\alpha}$ on both sides of the second relation of (3.35) and by using (3.34) we obtain
-$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=d_{\gamma}g^{\beta\gamma},$ (3.39)
+By multiplying $e_{\gamma}$ on both sides of the second relation of (3.38) and by using (3.37) we obtain
+$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\alpha\beta}=d_{\beta}g^{\alpha\beta},$ (3.42)
 so it follows from (3.40) that
-$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}\right)=(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}=d_{\gamma}g^{\beta\gamma}.$ (3.40)
+$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\alpha\nu}\frac{\partial f^{\beta}}{\partial t^{\nu}}\right)=(d_{\beta}+d_{\alpha})\eta^{\alpha\nu}\frac{\partial f^{\beta}}{\partial t^{\nu}}=d_{\beta}g^{\alpha\beta}.$ (3.43)
 Thus we have
-$(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial^{2}f^{\gamma}}{\partial t^{\nu}\partial t^{\alpha}}=d_{\gamma}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}},$ (3.41)
+$(d_{\alpha}+d_{\beta})\eta^{\alpha\nu}\frac{\partial^{2}f^{\beta}}{\partial t^{\nu}\partial t^{\gamma}}=d_{\beta}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}},$ (3.44)
 which leads to the validity of (3.41). The lemma is proved. ∎
 Theorem 3.10.
-[…]
+The algebra defined by the multiplication (3.29) on $T^{*}\mathcal{M}_{D}$ is a Frobenius algebra with the bilinear form $\langle\cdot\,$$\cdot\rangle$ given by the flat metric $\eta$ and with the unity (3.30).
 Proof.
@@
 which verifies the compatibility of the multiplication with the bilinear form. The proposition is proved. ∎
-The Frobenius algebra on $T^{*}\mathcal{M}$ induces a Frobenius algebra structure on $T(M\setminus D)$ via the isomorphism
-$\eta^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.43)
-where $D$ is define by (3.13). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T(M\setminus D)$ is defined by
-$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}),$ (3.44)
+The Frobenius algebra on $T^{*}\mathcal{M}_{D}$ induces a Frobenius algebra structure on $T\mathcal{M}_{D}$ via the isomorphism
+$\eta^{\sharp}\colon T^{*}\mathcal{M}_{D}\to T\mathcal{M}_{D},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.46)
+where $D$ is define by (3.16). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T\mathcal{M}_{D}$ is defined by
+$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}_{D}),$ (3.47)
 with the bilinear form
@@
 and the unity
-$e=\eta^{\sharp}(\omega_{e}),$ (3.46)
-here $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.14), (3.20) and (3.27) it follows that
+$e=\eta^{\sharp}(\omega_{e})=-\frac{1}{\kappa}\mathrm{grad}_{\eta}\sum_{r=1}^{\ell}m_{r}\log z^{r},$ (3.49)
+where $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.17), (3.23) and (3.30) it follows that
 $\mathcal{L}_{E}e=-\kappa e.$
-3.4. The generalized Frobenius manifold structure on $\mathcal{M}\setminus D$
-The notion of a generalized Frobenius manifold is a certain generalization of Dubrovin’s definition of Frobenius manifolds [9], it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
+3.4. The generalized Frobenius manifold structure on $\mathcal{M}_{D}$
+A generalized Frobenius manifold extends Dubrovin’s original definition of a Frobenius manifold [9]; it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
 Definition 3.11 ([9, 22]).
-[…]
+A (complex) generalized Frobenius manifold (GFM) is a complex manifold $M$ endowed on each of its tangent spaces $T_{p}M$ a Frobenius algebra $A_{p}=(\cdot,\langle\cdot\,,\cdot\rangle,e)$ which depends holomorphically on $p$, here $\cdot$ is the operation of multiplication of the algebra, $\langle\cdot\,$$\cdot\rangle$ is a symmetric non-degenerate bilinear form which is compatible with the operation of multiplication, and $e$ is the unity.
+It is required to satisfy the following axioms:
   1. FM1.
-[…]
+    The bilinear form $\langle\cdot\,$$\cdot\rangle$ yields a flat metric on $M$.
   2. FM2.
@@
     $\nabla c:(W,X,Y,Z)\to(\nabla_{W}c)(X,Y,Z)$ (3.50)
-    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$.
+    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$, where $\nabla$ is the Levi-Civita connection of the flat metric.
   3. FM3.
-    There exist an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
-    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y.$ (3.48)
+    There exists an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
+    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y,$ (3.51)
     $\mathcal{L}_{E}\langle X,Y\rangle$ $=\langle\mathcal{L}_{E}X,Y\rangle+\langle X,\mathcal{L}_{E}Y\rangle+(2-d)\langle X,Y\rangle.$ (3.52)
-    Here $\nabla$ is the Levi-Civita connection of the flat metric, and $d$ is a constant whic is called the charge of $M$.
-Let us show that the Frobenius algebra on $T(\mathcal{M}\setminus D)$ that is defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}\setminus D$ with the Euler vector field
-$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.50)
+    here $d$ is a certain constant called the charge of $M$.
+Lemma 3.12.
+The Frobenius algebra on $T\mathcal{M}_{D}$ defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}_{D}$ with the Euler vector field
+$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}t^{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.53)
 and with charge $d=1$.
+Proof.
 The validity of the axiom FM1 follows from the flatness of the metric $\eta$, and that of the axiom FM2 follows from the relation (3.45) and the second relation of (3.39).
@@
 $\partial_{\tilde{E}}c_{\alpha\beta}^{\gamma}$ $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}+\frac{d_{\gamma}}{\kappa}\right)c_{\alpha\beta}^{\gamma},$ (3.54)
-00 0 $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.52)
+$0$ $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.55)
 Here
@@
 The validity of the identities follows from (3.23) and the fact that $\deg\Gamma^{\alpha\beta}_{\gamma}=d_{\alpha}+d_{\beta}-d_{\gamma}$.
-Thus we have the following theorem.
+∎
+Now we give the main theorem of this paper.
 Theorem 3.13 (Main).
-Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}\setminus D$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.44).
+Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, and the vector field $E$ given by (3.17) is diagonalizable, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}_{D}$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.47).
 Moreover, the unit vector field $e$ and the Euler vector field $\tilde{E}$ of this generalized Frobenius are given by (3.49) and (3.53) respectively, and the intersection form coincides with $g$.
 Proof.
-We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}\setminus D$ coincides with $g$, i.e.,
+We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}_{D}$ coincides with $g$, i.e.,
 $g(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})=i_{\tilde{E}}(\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta}).$ (3.58)
@@
 Here we used the fact that $\deg g^{\alpha\beta}=d_{\alpha}+d_{\beta}$. The theorem is proved. ∎
-Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widehat{\mathcal{M}\setminus D}$ of $\mathcal{M}\setminus D$, hence $\pi_{1}(\mathcal{M}\setminus D)$ acts on those functions.
-Denote the stable subgroup of this action by
-$H=\{\gamma\in\pi_{1}(\mathcal{M}\setminus D)\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),~{}\forall p\in\widehat{\mathcal{M}\setminus D}\},$ (3.57)
-then $H$ corresponds to a covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$. Concretely, we may define an equivalence relation $\sim$ on $\widehat{\mathcal{M}\setminus D}$ by
-$p\sim p^{\prime}\iff\pi(p)=\pi(p^{\prime}),\quad t^{\alpha}(p)=t^{\alpha}(p^{\prime}),\quad\alpha=1,\dots,\ell.$ (3.58)
-Here $\pi\colon\widehat{\mathcal{M}\setminus D}\to\mathcal{M}\setminus D$ is the canonical projection. Then we have
-¯$\overline{\mathcal{M}}=\widehat{\mathcal{M}\setminus D}/\sim.$ (3.59)
-We are to show in Corollary 4.3 that $t^{1},\dots,t^{\ell}$ are coordinates of $\overline{\mathcal{M}}$.
-Therefore, the generalized Frobenius manifold structure that we constructed above is defined on $\overline{\mathcal{M}}$ instead of on $\mathcal{M}\setminus D$.
-If there is no ambiguity, we usually identify $\overline{\mathcal{M}}$ with $\mathcal{M}\setminus D$ since they are locally isomorphic.
-However, when we discuss the monodromy group of the generalized Frobenius manifold structure in the next section, it is necessary to view this generalized Frobenius manifold as $\overline{\mathcal{M}}$.
-We will prove in the next section that $z^{1},\dots,z^{\ell}$ are polynomials of $t^{1},\dots,t^{\ell}$, so we have the canonical projection
-$\overline{\pi}:~{}\overline{\mathcal{M}}\to\mathcal{M}\setminus D.$
-Since $\overline{\mathcal{M}}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we will also denote this generalized Frobenius manifold by
-¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$ (3.60)
-If there is no confusion on the selection of the pencil generators, we simply write it as
-¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega).$ (3.61)
-4. Polynomial Property and the Monodromy Groups
-In this section, we consider the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$.
+Since the generalized Frobenius manifold structure of $\mathcal{M}_{D}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we denote this generalized Frobenius manifold by
+$\mathcal{M}_{D}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$
+If there is no ambiguity in the selection of the pencil generators, we simply write it as
+$\mathcal{M}_{D}=\mathcal{M}(R,\omega).$
+4. Polynomiality Property and the Monodromy Groups
+In this section, we consider the monodromy group of the generalized Frobenius manifold $\mathcal{M}_{D}=\mathcal{M}(R,\omega)$.
 Note that the monodromy groups of the Frobenius manifold structures that are constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups coincide with the underlined groups.
-In the present case the monodromy group $\mathrm{Mono}(\overline{\mathcal{M}})$ can be embedded into the affine Weyl group $W_{a}(R)$.
-More precisely, we will show that the image of this embedding satisfies
-$\mathrm{Mono}(\overline{\mathcal{M}})~{}\leq~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell},$ (4.1)
+In the present case the monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$ can be embedded into the affine Weyl group $W_{a}(R)$.
+More precisely, we will show that the image of this embedding is
+$\mathrm{Mono}(\mathcal{M}_{D})~\cong~\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell},$ (4.1)
 where
-$\mathrm{Stab}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}.$ (4.2)
-To this end, we first need to prove Theorem 4.2 which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
-4.1. Periods of the generalized Frobenius manifold
-Let us consider flat coordinates of the flat metric $g+\lambda\eta$ of $\overline{\mathcal{M}}$. They satisfy the equations
-$\nabla_{g+\lambda\eta}\mathrm{d}u(\lambda)=0$ (4.3)
+$\mathrm{Stab}_{W}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}$ (4.2)
+is a parabolic subgroup of $W(R)$.
+Moreover, we will construct another generalized Frobenius manifold $\widehat{\mathcal{M}}_{D}$ as a covering space of $\mathcal{M}_{D}$, on which the flat coordinates $t^{1},\dots,t^{\ell}$ are globally defined, and show that its monodromy group is a subgroup of $\mathrm{Mono}(\mathcal{M}_{D})$.
+To this end, we first need to prove Theorem 4.3 by using properties of the periods of $\mathcal{M}_{D}$, which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
+4.1. Periods of the generalized Frobenius manifolds
+Let us consider flat coordinates $\tilde{v}(\lambda)$ of the metric $g+\lambda\eta$ of $\mathcal{M}_{D}$. They satisfy the equations
+$\nabla_{g+\lambda\eta}\mathrm{d}\tilde{v}(\lambda)=0,$ (4.3)
 and are called periods of the generalized Frobenius manifold [9, 22]. Denote
-$M_{\lambda}=(\overline{\mathcal{M}}\times\mathbb{C}^{\times})\setminus{\Sigma_{\lambda}},$ (4.4)
+$\mathcal{M}_{\Lambda}=(\mathcal{M}_{D}\times\mathbb{C}^{\times})\setminus{\Sigma_{\Lambda}},$ (4.4)
 where
-${\Sigma}_{\lambda}=\{(z,\lambda)\in M\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
+$\Sigma_{\Lambda}=\{(z,\lambda)\in\mathcal{M}_{D}\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
 Then solutions of the equation (4.3) live on
-$\mathscr{Q}(\widehat{M_{\lambda}})=\mathscr{O}(\widehat{M_{\lambda}})/\mathbb{C}.$
-In the coordinates $(z^{1},\dots,z^{\ell})$, the equation (4.3) reads
-$\frac{\partial^{2}u}{\partial z^{i}\partial z^{j}}-\Gamma_{\lambda,ij}^{k}\frac{\partial u}{\partial z^{k}}=0,$ (4.5)
-Here $\Gamma_{\lambda,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$.
-Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$.
-By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ such that
-$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}\right).$ (4.6)
+$\mathscr{Q}(\widetilde{\mathcal{M}}_{\Lambda})=\mathscr{O}(\widetilde{\mathcal{M}}_{\Lambda})/\mathbb{C},$
+where $\widetilde{\mathcal{M}}_{\Lambda}$ is the universal covering space of $\mathcal{M}_{\Lambda}$.
+In the coordinates $z^{1},\dots,z^{\ell}$, the equation (4.3) reads
+$\frac{\partial^{2}\tilde{v}}{\partial z^{i}\partial z^{j}}-\Gamma_{g_{\lambda},ij}^{k}\frac{\partial\tilde{v}}{\partial z^{k}}=0,$ (4.5)
+here $\Gamma_{g_{\lambda},ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$.
+Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a/4\pi^{2}$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$ via the $\varphi_{\lambda}$-transformation.
+By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ of $\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}$ such that
+$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(x^{1}+\theta_{1}c,\dots,x^{\ell}+\theta_{\ell}c\right).$ (4.6)
 Therefore, the refined flat coordinates
 $\check{x}^{r}(z,\lambda):=x^{r}+\theta_{r}c=x^{r}+\frac{\theta_{r}}{\kappa}\log\lambda,\quad r=1,\dots,\ell$ (4.7)
-depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are homogeneous of degree zero, i.e., they satisfy the equation
-$\kappa\lambda\frac{\partial u}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial u}{\partial z^{j}}=0.$ (4.8)
-We denote by $\mathcal{S}_{\lambda}$ the space of periods of $\overline{\mathcal{M}}$ which are homogeneous of degree zero, then $\dim\mathcal{S}_{\lambda}=\ell$, and
-$\mathcal{S}_{\lambda}=\mathrm{span}_{\mathbb{C}}\left\{\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)\right\}.$ (4.9)
-Now we proceed to find another basis of $\mathcal{S}_{\lambda}$ that is associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
-Proposition 4.1.
-For each homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a solution
-$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ (4.10)
-of the equations (4.5) and (4.8), where $h^{\alpha}(z,\rho)$ can be analytically continued to $\rho=0$ such that $h^{\alpha}(z,0)=t^{\alpha}(z)$.
+depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are quasi-homogeneous of degree zero, i.e., they satisfy the equation
+$\kappa\lambda\frac{\partial\tilde{v}}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial\tilde{v}}{\partial z^{j}}=0.$ (4.8)
+We denote by $\mathcal{S}_{\Lambda}$ the space of periods of $\mathcal{M}_{D}$ that are quasi-homogeneous of degree zero modulo constant functions, then $\mathcal{S}_{\Lambda}$ consists of the solutions to equations (4.5) and (4.8) up to additive constants, hence $\dim_{\mathbb{C}}\mathcal{S}_{\Lambda}=\ell$, and $\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)$ form a basis of $\mathcal{S}_{\Lambda}$.
+Now we proceed to find another basis of $\mathcal{S}_{\Lambda}$ associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
+Theorem 4.1.
+For each quasi-homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, $\alpha=1,\dots,\ell$, there exist a unique quasi-homogeneous period $\tilde{v}\in\mathcal{S}_{\Lambda}$ such that
+$\tilde{v}=\tilde{v}^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho),$ (4.9)
+where is holomorphic at $\rho=0$ with $h^{\alpha}(z,0)=t^{\alpha}(z)$, or equivalently, $h^{\alpha}(z,\rho)$ has the form
+$h^{\alpha}(z,\rho)=t^{\alpha}(z)+\sum_{j=1}^{\infty}h^{\alpha}_{j}(z)\rho^{j}.$ (4.10)
 Proof.
-Recall that we made the assumption in Remark 3.1 that $d_{\alpha}>0$, so from (3.20) we have $0<d_{\alpha}<\kappa$.
-Let us looking for a solution of (4.5) and (4.8) of the form $u(z,\rho)=\rho^{d_{\alpha}/\kappa}h(z,\rho)$ satisfying the initial condition $h(z,0)=t^{\alpha}(z)$.
+Let us look for a solution of (4.5) and (4.8) of the form $\tilde{v}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ satisfying the initial condition $h^{\alpha}(z,0)=t^{\alpha}(z)$.
 Denote $\xi_{j}=\frac{\partial h}{\partial z^{j}}$, then these equations take the form
@@
 $-\kappa\rho\frac{\partial\xi_{j}}{\partial\rho}+\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial\xi_{j}}{\partial z^{r}}$ $=(d_{\alpha}-\theta_{j})\xi_{j},$ (4.12)
-where $\Gamma_{ij}^{k}(\rho)=\Gamma_{1/\rho,ij}^{k}$, which are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$.
-By using the equations (4.11) we can rewrite the equations (4.12) as
+where $\Gamma_{ij}^{k}(\rho)$ are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$. By using the equations (4.11) we can rewrite the equations (4.12) as
 $\kappa\rho\frac{\partial\vec{\xi}}{\partial\rho}=\left(A(\rho)-d_{\alpha}\right)\vec{\xi},$ (4.13)
@@
 $\vec{\xi}=\vec{\xi}(\rho)=(\xi_{1},\cdots,\xi_{\ell})^{T},\quad A_{j}^{k}(\rho)=\sum_{r=1}^{\ell}\theta_{r}z^{r}\Gamma_{rj}^{k}(\rho)+\theta_{j}\delta_{j}^{k}.$
-Expand $A(\rho)$ and $\zeta{\xi}(\rho)$ into power series of $\rho$ as follows:
-$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{0}+\rho\vec{\xi}_{1}+\rho^{2}\vec{\xi}_{2}+\dots,$
+Note that $\rho=0$ is a Fuchsian singularity of equation (4.13), so we can solve it by expanding $A(\rho)$ and $\vec{\xi}(\rho)$ into power series of $\rho$ as follows:
+$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{(0)}+\rho\vec{\xi}_{(1)}+\rho^{2}\vec{\xi}_{(2)}+\dots,$
 here $A=A(0)=(A^{k}_{j}(0))$ coincides with the matrix $A$ defined in (3.26), and
-$\vec{\xi}_{0}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T},$ (4.14)
-then the equation (4.13) is equivalent to
-$(A-d_{\alpha})\vec{\xi}_{0}=\mathbf{0},\quad(A-d_{\alpha}-\kappa)\vec{\xi}_{1}=-A_{1}\vec{\xi}_{0},$ (4.15)
-$(A-d_{\alpha}-2\kappa)\vec{\xi}_{2}=-A_{2}\vec{\xi}_{0}-A_{1}\vec{\xi}_{1},\dots$ (4.16)
-From (3.24) we know that the first equation of (4.15) is satisfied by $\vec{\xi}_{0}$ that is defined by (4.14).
-Since the eigenvalues of $A$ are positive and less than $\kappa$, we can solve the other linear equations of (4.15) and (4.16) to obtain a unique solution $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ recursively.
-From the fact that $\vec{\xi}_{0}$ gives a solution of the equations (4.11), and from the compatibility of of the equations of (4.11) and (4.12) we know that $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition.
+$\vec{\xi}_{(0)}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T}.$ (4.14)
+Note that $\rho=0$ is a Fuchsian singularity of equation (4.13), its solutions of formal power series always admit a positive radius of convergence.
+As a result, by expanding equation (4.13) into power series of $\rho$, we only need to solve the linear equations
+$(A-d_{\alpha})\vec{\xi}_{(0)}=$ $~\mathbf{0},$ (4.15)
+$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j)}=$ $-\sum_{r=1}^{j}A_{r}\vec{\xi}_{(j-r)},\quad j=1,2,\dots.$ (4.16)
+From (3.27) we know that (4.15) automatically holds.
+Recall from Lemma 3.5 that for each $\beta=1,\dots,\ell$, $\mathrm{Re}(d_{\beta})>0$, so from (3.23) we have $0<\mathrm{Re}(d_{\beta})<\kappa$.
+This indicates that all the eigenvalues of $A$ have real parts between 0 and $\kappa$, hence we can solve (4.16) to obtain a unique solution $\vec{\xi}_{(1)},\vec{\xi}_{(2)},\dots$ recursively.
+From the fact that $\vec{\xi}_{(0)}$ gives a solution of the equations (4.11) and the compatibility of the equations of (4.11) and (4.12), we know that $\vec{\xi}_{(1)},\vec{\xi}_{(2)},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition.
 ∎
 4.2. The polynomiality property
-The periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\overline{\mathcal{M}}$ given by Proposition 4.1 are linearly independent, so they yield another basis of $\mathcal{S}_{\lambda}$.
+Since periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\mathcal{M}_{D}$ given by Theorem 4.1 are linearly independent (indeed, they are linearly independent near $\rho=0$), they form another basis of $\mathcal{S}_{\Lambda}$.
 Thus there exist constants $c_{\alpha}^{\beta}$ and $s^{\beta}$ such that
-$\check{x}^{\beta}(z,\rho)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
-Since the functions $Z_{r}=Z_{r}(\mathrm{e}^{2\pi i\check{x}^{1}},\dots,\mathrm{e}^{2\pi i\check{x}^{\ell}})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$.
-Hence we have
+$\check{x}^{\beta}\left(z,1/\rho\right)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
+Since the functions $Z_{r}=Z_{r}(\check{x}^{1},\dots,\check{x}^{\ell})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$.
+Hence we can expand them to get
 $z^{r}=\rho^{-\theta_{r}/\kappa}Z_{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}(v^{1}(z,\rho))^{n_{1}}\cdots(v^{\ell}(z,\rho))^{n_{\ell}},$ (4.18)
 where $a_{n_{1},\dots,n_{\ell}}$ are some constants.
+In order to prove the polynomial dependence of $z^{1},\dots,z^{\ell}$ on $t^{1},\dots,t^{\ell}$, we need the following lemma.
+Lemma 4.2.
+Suppose $N$ be a positive integer, $b_{1},\dots,b_{N}$ be distinct real numbers, and $p_{1},\dots,p_{N}$ be complex numbers. If
+$\lim_{\rho\to 0}\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}=0,$ (4.19)
+then $p_{1}=\cdots=p_{N}=0$.
+Proof.
+Let us define the function $h\colon\mathbb{R}\to\mathbb{C}$ as follows:
+$h(s)=\sum_{j=1}^{N}p_{j}e^{-\sqrt{-1}\,b_{j}s}.$
+Then from the assumption of the lemma we get
+$\lim_{s\to+\infty}h(s)=0,$
+from which it follows that
+$\lim_{T\to+\infty}\frac{1}{T}\int_{0}^{T}h(s)e^{\sqrt{-1}\,\alpha s}\operatorname{d\!}s=0,\quad\forall\alpha\in\mathbb{R}.$
+By taking $\alpha=b_{j}$ for $j=1,\dots,N$, we arrive at $p_{j}=0$. The lemma is proved. ∎
 Theorem 4.3.
@@
 Proof.
-It follows from Proposition 4.1 that
-$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\rho\to 0.$
-Thus we have
+It follows from Theorem 4.1 that
+$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\textrm{when}\ \rho\to 0.$
+Here and in what follows, the limit $\rho\to 0$ is taken pointwise for each $z\in\mathcal{M}_{D}$. Thus we have
 $z^{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}\left(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa})\right).$ (4.20)
 Let us denote
-$\theta=\inf\{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}~{}\big{|}~{}a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.20)
-Suppose $\theta>\theta_{r}$, then it follows from (4.19) that
+$\theta=\inf\{\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})~\big|~a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.21)
+We now show that $\theta=\theta_{r}$. Suppose $\theta>\theta_{r}$, then it follows from (4.20) that
 $z^{r}=\rho^{-\theta_{r}/\kappa}o(\rho^{\theta_{r}/\kappa})=o(1),\quad\rho\to 0,$ (4.22)
-this is not true due to our definition of proper generators. On the other hand, if $\theta<\theta_{r}$, then from (4.19) we know that
-$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\rho\to 0.$ (4.22)
-Since $(\theta-\theta_{r})/\kappa<0$, we must have
-$\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}=0.$
-However, since $t^{1},\cdots,t^{\ell}$ are functionally independent, all those coefficients $a_{n_{1},\cdots,n_{\ell}}$ must vanish.
-This contradicts with the definition (4.20) of $\theta$.
-Thus we have $\theta=\theta_{r}$, and it follows that
-$z^{r}=\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}.$ (4.23)
-In particular, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$. The theorem is proved. ∎
+However, this is not true if $z^{r}\not=0$. On the other hand, if $\theta<\theta_{r}$, then from (4.20) we know that
+$z^{r}=~\rho^{(\theta-\theta_{r})/\kappa}\sum_{\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})=\theta}\rho^{\sqrt{-1}\,\mathrm{Im}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),$
+when $\rho\to 0$.
+Since each $d_{\alpha}$ admits positive real part, we know that only finitely many vectors $(n_{1},\dots,n_{\ell})\in\mathbb{Z}_{\geq 0}^{\ell}$ satisfy the condition $\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})=\theta$.
+Suppose $b_{1},\dots,b_{N}$ be all possible distinct values of $\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})$ for those $(n_{1},\dots,n_{\ell})$, then we have
+$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}(t^{1},\dots,t^{\ell})+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\textrm{when}\ \rho\to 0,$
+where $p_{1},\dots,p_{N}$ are some polynomials of $t^{1},\dots,t^{\ell}$. Since $(\theta-\theta_{r})/\kappa<0$, we must have
+$\sum_{j=1}^{N}\rho^{\sqrt{-1}\,b_{j}}p_{j}(t^{1},\dots,t^{\ell})\to 0,\quad\textrm{when}\ \rho\to 0.$ (4.23)
+Thus, by using Lemma 4.2 we arrive at $p_{j}=0$ for $j=1,\dots,N$.
+Since $t^{1},\dots,t^{\ell}$ are functionally independent, all the coefficients $a_{n_{1},\dots,n_{\ell}}$ of the polynomials $p_{j}$ must vanish.
+This contradicts with the definition (4.21) of $\theta$, so we have $\theta=\theta_{r}$.
+Now, we can represent $z^{r}$ in the form
+$z^{r}$ $=\sum_{\mathrm{Re}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})=\theta_{r}}\rho^{\sqrt{-1}\,\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(1),$ (4.24)
+$=p_{0}(t^{1},\dots,t^{\ell})+\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}(t^{1},\dots,t^{\ell})+o(1),\quad\textrm{when}\ \rho\to 0.$ (4.25)
+Here $b_{1},\dots,b_{N}$ are all distinct nonzero possible values of $\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})$, and $p_{0},p_{1},\dots,p_{N}$ are polynomials of $t^{1},\dots,t^{\ell}$.
+By usingLemma 4.2 again, we get $p_{j}=0$ for $j=1,\dots,N$ and $z^{r}=p_{0}$, so $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$.
+The theorem is proved.
+∎
 Remark 4.1.
-This theorem shows that the proper generators are polynomials of the flat coordinates of $\eta$.
-It deserves to note that this property also holds true for the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and on the extended affine Weyl groups [12, 14].
-For the Frobenius manifold structures constructed in [9] and [12], the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s, while for the ones constructed in [14], $t^{\alpha}$ may involve radical expressions of $z^{j}$’s.
-For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell})$ for $\ell\geq 5$.
-From Theorem 4.2, one can verify that each $g^{\alpha\beta}$ is a rational function of $t^{1},\cdots,t^{\ell}$.
-Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\overline{\mathcal{M}}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
-$\eta^{\gamma\nu}\frac{\partial^{3}F}{\partial t^{\alpha}\partial t^{\beta}\partial t^{\nu}}=c^{\gamma}_{\alpha\beta},\quad\mathcal{L}_{\tilde{E}}F=2F+\textrm{quadratic functions of}\ t^{1},\dots,t^{\ell},$
-up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$, here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.53).
-From the relations (3.38) and (3.53) it follows that we can choose the potential $F$ such that
-$\eta^{\alpha\nu}\eta^{\beta\xi}\frac{\partial^{2}F}{\partial t^{\nu}\partial t^{\xi}}=\frac{\kappa}{d_{\alpha}+d_{\beta}}g^{\alpha\beta}.$ (4.24)
-Therefore, all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
+The above theorem shows that the pencil generators are polynomials of the flat coordinates of $\eta$.
+This property is shared by the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and of the extended affine Weyl groups [12, 14].
+The key distinction lies in the inverse relationship: while the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s for the Frobenius manifold structures constructed in [9] and [12], they may involve radical expressions of $z^{j}$’s for the ones constructed in [14].
+For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell},\{y_{1},\dots,y_{\ell}\})$ for $\ell\geq 5$.
 Corollary 4.4.
-Define the covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$ as (3.59). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\overline{\mathcal{M}}$.
+All degrees $d_{\alpha}$ are positive rational numbers.
+Proof.
+By Lemma 3.5, it is sufficient to prove that $d_{\alpha}\in\mathbb{Q}$.
+From Theorem 4.3 we know that each $z^{r}=z^{r}(t),r=1,\dots\ell$ is a quasi-homogeneous polynomial of $t^{1},\dots,t^{\ell}$ of degree $\theta_{r}$, so for any of its monomial term of the form $c(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}$ we have
+$n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}.$
+Since $z^{1}(t),\dots,z^{\ell}(t)$ are functionally independent, among their monomial terms we can find $\ell$ terms
+$f_{k}=c_{k}(t^{1})^{n_{k1}}\cdots(t^{\ell})^{n_{k\ell}},\quad k=1,\dots,\ell$
+which are functionally independent, and have degrees $\theta_{m_{1}},\dots,\theta_{m_{\ell}}$. Thus we obtain the following system of linear equations for $d_{1},\dots,d_{\ell}$:
+$n_{k1}d_{1}+\cdots+n_{k\ell}d_{\ell}=\theta_{m_{k}},\quad k=1,\dots,\ell.$ (4.26)
+Due to the functional independence of $f_{1},\dots,f_{\ell}$, the Jacobian determinant
+$\det\begin{pmatrix}\frac{\partial f_{1}}{\partial t^{1}}&\cdots&\frac{\partial f_{1}}{\partial t^{\ell}}\\ \vdots&\ddots&\vdots\\ \frac{\partial f_{\ell}}{\partial t^{1}}&\cdots&\frac{\partial f_{\ell}}{\partial t^{\ell}}\end{pmatrix}=\frac{f_{1}\cdots f_{\ell}}{t^{1}\cdots t^{\ell}}\det\begin{pmatrix}n_{11}&\cdots&n_{1\ell}\\ \vdots&\ddots&\vdots\\ n_{\ell 1}&\cdots&n_{\ell\ell}\end{pmatrix}$
+does not vanish identically.
+It implies that the system of equations (4.26) has non-degenerate coefficient matrix, hence it determines $d_{1},\dots,d_{\ell}$ uniquely in $\mathbb{Q}$.
+The corollary is proved.
+∎
+Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widetilde{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$, hence $\pi_{1}(\mathcal{M}_{D})$ acts on those functions.
+Denote the stable subgroup of this action by
+$H=\{\gamma\in\pi_{1}(\mathcal{M}_{D})\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),\ \forall p\in\widetilde{\mathcal{M}}_{D},~\alpha=1,\dots,\ell\},$ (4.27)
+then $H$ corresponds to a covering space $\widehat{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$. Concretely, we may define an equivalence relation $\sim$ on $\widetilde{\mathcal{M}}_{D}$ by
+$p\sim p^{\prime}\iff\pi(p)=\pi(p^{\prime}),\quad t^{\alpha}(p)=t^{\alpha}(p^{\prime}),\quad\alpha=1,\dots,\ell.$ (4.28)
+Here $\pi\colon\widetilde{\mathcal{M}}_{D}\to\mathcal{M}_{D}$ is the canonical projection. Then we have
+$\widehat{\mathcal{M}}_{D}=\widetilde{\mathcal{M}}_{D}/\sim$ (4.29)
+with covering map
+$\widehat{\pi}:~\widehat{\mathcal{M}}_{D}\to\mathcal{M}_{D}.$
+Corollary 4.5.
+Define the covering space $\widehat{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$ as (4.29). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\widehat{\mathcal{M}}_{D}$.
 Proof.
 To prove that $(t^{1},\cdots,t^{\ell})$ are global coordinates, it suffices to show that
-$(t^{1},\cdots,t^{\ell})\colon\overline{\mathcal{M}}\to\mathbb{C}^{\ell}$ (4.25)
-is injective. Suppose $p,p^{\prime}\in\overline{\mathcal{M}}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~{}\alpha=1,\cdots,\ell$, then by using Theorem 4.2 we know that
+$(t^{1},\cdots,t^{\ell})\colon\widehat{\mathcal{M}}_{D}\to\mathbb{C}^{\ell}$ (4.30)
+is injective. Suppose $p,p^{\prime}\in\widehat{\mathcal{M}}_{D}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~\alpha=1,\cdots,\ell$, then by using Theorem 4.3 we know that
 $z^{r}(t^{1}(p),\cdots,t^{\ell}(p))=z^{r}(t^{1}(p^{\prime}),\cdots,t^{\ell}(p^{\prime})).$
-This implies that $\overline{\pi}(p)=\overline{\pi}(p^{\prime})$, where $\overline{\pi}:\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is the canonical projection.
-Therefore $p\sim p^{\prime}$, i.e., they are identified on $\overline{\mathcal{M}}$.
-Hence (4.25) is injective, and we proved the corollary.
-∎
-In conclusion, we have the following commutative diagram:
+This implies that $\widehat{\pi}(p)=\widehat{\pi}(p^{\prime})$, hence $p\sim p^{\prime}$, i.e., they are identified on $\widehat{\mathcal{M}}_{D}$. Therefore, (4.30) is injective. ∎
+Corollary 4.6.
+Suppose $F$ is the prepotential of $\widehat{\mathcal{M}}_{D}$, then all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
+Proof.
+From Theorem 4.3, one can verify that each $g^{\alpha\beta}=g(\mathrm{d}t^{\alpha},\mathrm{d}t^{\beta})$ is a rational function of $t^{1},\cdots,t^{\ell}$.
+Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\widehat{\mathcal{M}}_{D}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
+$\eta^{\gamma\nu}\frac{\partial^{3}F}{\partial t^{\alpha}\partial t^{\beta}\partial t^{\nu}}=c^{\gamma}_{\alpha\beta},\quad\mathcal{L}_{\tilde{E}}F=2F+\textrm{quadratic functions of}\ t^{1},\dots,t^{\ell},$
+up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$.
+Here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.56).
+From the relations (3.41) and (3.56), we can choose the potential $F$ such that
+$\eta^{\alpha\nu}\eta^{\beta\xi}\frac{\partial^{2}F}{\partial t^{\nu}\partial t^{\xi}}=\frac{\kappa}{d_{\alpha}+d_{\beta}}g^{\alpha\beta}.$ (4.31)
+Therefore, all second derivatives of $F$ are rational functions of $t^{1},\dots,t^{\ell}$. ∎
+4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
+The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and can be extended directly to the case of generalized Frobenius manifolds.
+Let us first recall this defintion.
+Let $M$ be an $\ell$-dimensional generalized Frobenius manifold with flat metric $\eta$ and intersection form $g$. Consider the Gauss-Manin equation
+$\nabla_{g}\mathrm{d}v=0.$ (4.32)
+for the flat coordinates of $g$. Denote
+$\Sigma=\{z\in M\mid\det g(z)=0\},$
+then the solutions of (4.32) live on the universal covering $\mathcal{V}$ of $M\setminus\Sigma$.
+Suppose $v^{1},\dots,v^{\ell}$ be flat coordinates of $g$, then they span, together with constant functions, the solution space of (4.32).
+For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, $\gamma$ acts on each $v^{\alpha}$ by analytic continuation.
+Since this action preserves solutions of (4.32), $\gamma^{*}v^{\alpha}$ is still a flat coordinate.
+Therefore, $\gamma$ yields an affine transformation
+$\gamma^{*}v^{\alpha}=A_{\beta}^{\alpha}(\gamma)v^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.33)
+on the solution space of (4.32). Here the linear part $A^{\alpha}_{\beta}$ is orthogonal w.r.t. the metric $g$. Let
 […]
-In other words, upon taking coordinates $(t^{1},\cdots,t^{\ell})$ on $\overline{\mathcal{M}}$ and $(z^{1},\cdots,z^{\ell})$ on $\mathcal{M}\setminus D$, the projection $\overline{\pi}$ coincides with the natural map $\mathrm{Spec}(\mathbb{C}[t^{1},\cdots,t^{\ell}])\to\mathrm{Spec}(\mathbb{C}[z^{1},\cdots,z^{\ell}])$.
-4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
-The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and it can be extended directly to generalized Frobenius manifolds. Let us first recall this defintion.
-Let $M$ be an $\ell$-dimensional generalized Frobenius manifold, of which the flat metric is $\eta$ and the intersection form is $g$, and let $t^{1},\cdots,t^{\ell}$ be flat coordinates of $\eta$.
-Consider the system of equations
-$\frac{\partial^{2}u}{\partial t^{\alpha}\partial t^{\beta}}=\Gamma_{\alpha\beta}^{\epsilon}\frac{\partial u}{\partial t^{\epsilon}},\quad\alpha,\beta=1,\dots,\ell$ (4.26)
-for the flat coordinates of $g$.
-Here $\Gamma_{\alpha\beta}^{\epsilon}$ are the Christoffel symbols of the Levi-Civita connection of $g$.
-Denote $\Sigma=\{t\in M~{}|~{}\det g(t)=0\}$, then the solutions $u$ live on the universal covering $\widehat{M\setminus\Sigma}$ of $M\setminus\Sigma$.
-Define the solution space
-[…]
-then we know that $\dim\mathcal{S}=n$, and there exist flat coordinates $u^{1},\cdots,u^{\ell}$ of $g$ such that
-$\mathcal{S}=\mathrm{span}\{u^{1},\cdots,u^{\ell}\}.$ (4.28)
-For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, we have a transformation
-$\gamma^{*}u^{\alpha}=A_{\beta}^{\alpha}(\gamma)u^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.29)
-on $\mathcal{S}$ by analytic continuation.
-Denote by $E^{\ell}$ the dual space of $\mathcal{S}$, then it has coordinates $u^{1},\cdots,u^{\ell}$.
-Hence the transformation (4.29) yields a transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$.
-Since $E^{\ell}$ is an $\ell$-dimensional complexified Euclidean space $\mathbb{R}^{\ell}\otimes\mathbb{C}$ equipped with the metric $g$, and the action by $\pi_{1}(M\setminus\Sigma)$ preserves the metric $g$, this transformation on $E^{\ell}$ is an isometry.
-Hence we have obtained a representation
-$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}.$ (4.30)
+and $E^{\ell}$ be the dual space of $\mathcal{S}$.
+Then $v^{1},\dots,v^{\ell}$ span $\mathcal{S}$, and they also give a coordinate system on $E^{\ell}$.
+The transformation (4.33) yields an affine transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$.
+This transformation on $E^{\ell}$ is an isometry w.r.t.
+the metric $g$, hence we obtain a representation
+$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}$ (4.35)
+of the fundamental group of $M\setminus\Sigma$.
 Definition 4.7 ([9], Appendix G.).
 The image of $r$ given in (4.35) is called the monodromy group of $M$, which is denoted by
-$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.31)
-Now we proceed to compute the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Let
-$\overline{\Sigma}=\{t\in\overline{\mathcal{M}}\mid\det\left(g^{\alpha\beta}(t)\right)=0\},$
-and $\mathcal{U}(\mathcal{M})$ be the universal covering space of ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$.
-The flat coordiantes of $g$ are $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$, and they are functions on $\mathcal{U}(\mathcal{M})$.
-Moreover, $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so we can view the complexified Euclidean space $E^{\ell}$ as $V\otimes\mathbb{C}$.
+$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell}.$ (4.36)
+In our construction, the inner product space $E^{\ell}$ can be identified with $V\otimes\mathbb{C}$. Let
+$\Sigma=\left\{z\in\mathcal{M}_{D}\mid\det\left(g^{ij}(z)\right)=0\right\},$
+and $\mathcal{U}(\mathcal{M}_{D})$ be the universal covering space of $\mathcal{M}_{D}\setminus\Sigma$.
+Then the flat coordinates of $g$ are $x^{1}\bigr|_{\lambda=0},\cdots,x^{\ell}\bigr|_{\lambda=0}$, which are functions on $\mathcal{U}(\mathcal{M}_{D})$.
+Moreover, $x^{1}\bigr|_{\lambda=0},\cdots,x^{\ell}\bigr|_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so the complexified Euclidean space $E^{\ell}$ is canonically isomorphic to $V\otimes\mathbb{C}$.
 Hence we obtain a representation
-$r\colon\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.32)
+$r\colon\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell,\mathbb{C})\ltimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.37)
 where
 $\gamma^{*}x^{i}\bigr|_{\lambda=0}=A_{j}^{i}(\gamma)x^{j}\bigr|_{\lambda=0}+B^{i}(\gamma).$ (4.38)
-Here the image $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ is the monodromy group of $\overline{\mathcal{M}}$.
-However, this group is difficult to compute, and we will instead consider a group bigger than it.
+Here the image $r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))$ is the monodromy group of $\mathcal{M}_{D}$.
+Now we proceed to compute the monodromy group of the generalized Frobenius manifolds $\mathcal{M}_{D}$ and $\widehat{\mathcal{M}}_{D}$.
+We will end this subsection by proving that
+$\mathrm{Mono}(\mathcal{M}_{D})\cong\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$
+and $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is a subgroup of $\mathrm{Mono}(\mathcal{M}_{D})$.
+Proposition 4.8.
+There exists a natural embedding $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})\hookrightarrow\mathrm{Mono}(\mathcal{M}_{D})$.
+Proof.
 Recall that the covering map
-$\overline{\pi}\colon\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ (4.34)
-is nondegerated, so from
-$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)$ (4.35)
-we know that $\left(g^{\alpha\beta}(t)\right)$ degenerates if and only if $\left(g^{ij}(z(t)\right)$ degenerates. In other words, denote
-$\Sigma^{\prime}=\{z\in\mathcal{M}\setminus D\mid\det\left(g^{ij}(z)\right)=0\},$
-then we have $(\overline{\pi})^{-1}(\Sigma^{\prime})=\overline{\Sigma}$, and the map
-$\overline{\pi}:\overline{\mathcal{M}}\setminus\overline{\Sigma}\to\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ (4.36)
-is also a covering map, hence
-$\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\leq\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})).$ (4.37)
-Since ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$ is a covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, they share the same universal covering space $\mathcal{U}(\mathcal{M})$, and both $\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})$ and $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ act on $\mathcal{U}(\mathcal{M})$.
-Note that $g^{ij}(z)$ are polynomials of $z^{1},\cdots,z^{\ell}$, so in particular they are single valued functions on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
-This implies that the action of $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ on $\mathcal{U}(\mathcal{M})$ also leaves the metric $g$ invariant, so the representation (4.32) can be extended to a representation
-$r\colon\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.38)
-Thus it follows from (4.37) that
-$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))~{}\leq~{}r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))),$ (4.39)
-i.e., the monodromy group of $\overline{\mathcal{M}}$ is a subgroup of $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$.
-We are to show in what follows that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
+$\widehat{\pi}\colon\widehat{\mathcal{M}}_{D}\to\mathcal{M}_{D}$ (4.39)
+is nondegerate, so from
+$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)^{2}$ (4.40)
+it follows that $\left(g^{\alpha\beta}(t)\right)$ is degenerate if and only if $\left(g^{ij}(z(t))\right)$ is degenerate. In other words,
+$(\widehat{\pi})^{-1}(\Sigma)=\widehat{\Sigma}:=\left\{t\in\widehat{\mathcal{M}}_{D}\mid\det\bigl(g^{\alpha\beta}(t)\bigr)=0\right\},$
+and the map
+$\widehat{\pi}:\widehat{\mathcal{M}}_{D}\setminus\widehat{\Sigma}\to\mathcal{M}_{D}\setminus\Sigma$ (4.41)
+[…]
+$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})\hookrightarrow\mathrm{Mono}(\mathcal{M}_{D})$ (4.42)
+as subgroups of $\mathrm{Iso}(V\otimes\mathbb{C})$. The proposition is proved. ∎
+Now we proceed to show that $\mathrm{Mono}(\mathcal{M}_{D})$ is isomorphic to $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$. Recall that
+$\varphi_{0}\colon V\otimes\mathbb{C}\to\mathcal{M}$ (4.43)
+is a branched covering, which induces an unbranched covering map
+$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)\to\mathcal{M}_{D}\setminus\Sigma.$ (4.44)
+Since $\mathcal{U}(\mathcal{M}_{D})$ is the universal covering space of $\mathcal{M}_{D}\setminus\Sigma$, it is also the universal covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$.
+We denote the corresponding covering map by
+$j_{0}\colon\mathcal{U}(\mathcal{M}_{D})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma),\quad s\mapsto(x^{1}\bigr|_{\lambda=0}(s),\cdots,x^{\ell}\bigr|_{\lambda=0}(s)).$ (4.45)
+We have the following commutative diagram of covering spaces:
+$\mathcal{U}(\mathcal{M}_{D})$$\widehat{\mathcal{M}}_{D}\setminus\widehat{\Sigma}$$\mathcal{M}_{D}\setminus\Sigma$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$$\widehat{\pi}$$j_{0}$$\varphi_{0}$
+All the maps in the diagram are covering maps.
+To determine the image of $\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$ under the map $r$ that is introduced in (4.37), we are to use the isomorphism
+$\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)\cong\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M}_{D})}{\mathcal{M}_{D}\setminus\Sigma}\right),$ (4.46)
+here $\mathrm{Deck}(X/Y)$ is the deck transformation group on the covering space $X$ over $Y$.
+We will also denote by $\gamma$ the deck transformation on $\mathcal{U}(\mathcal{M}_{D})$ induced by $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$.
+Lemma 4.9.
+Let $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$, then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$ such that
+$j_{0}\circ\gamma=p\circ j_{0}.$ (4.47)
+In other words, we have the following commutative diagram:
+[…]
+Proof.
+The uniqueness follows from the fact that the covering map $j_{0}$ is surjective.
+Now we show that $p=r(\gamma)$ satisfies the commutative diagram.
+Since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M}_{D})$, $p=r(\gamma)$ satisfies (4.47).
+We then only need to verify that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$.
+Since $\gamma$ is a deck transformation over $\mathcal{M}_{D}\setminus\Sigma$, we have
+$\varphi_{0}\circ j_{0}=\varphi_{0}\circ j_{0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ j_{0}.$ (4.48)
+Since $j_{0}$ is surjective, it follows that
+$\varphi_{0}=\varphi_{0}\circ r(\gamma).$
+The lemma is proved. ∎
+Proposition 4.10.
+The monodromy group $r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))$ can be represented as
+$r(\pi_{1}(\mathcal{M}_{D}\setminus\Sigma))=\mathrm{Deck}\bigg(\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)}{\mathcal{M}_{D}\setminus\Sigma}\bigg).$ (4.49)
+In particular, every deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$ can be extended to an isometry on $V\otimes\mathbb{C}$.
+Proof.
+For any $\gamma\in\pi_{1}(\mathcal{M}_{D}\setminus\Sigma)$, we know from Lemma 4.9 that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$.
+On the other hand, suppose $p$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ over $\mathcal{M}_{D}\setminus\Sigma$.
+Since $\mathcal{U}(\mathcal{M}_{D})$ is their universal covering space, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M}_{D})\to\mathcal{U}(\mathcal{M}_{D})$.
+Therefore, we have the following commutative diagram:
+[…]
+Thus by using Lemma 4.9, we have $p=r(\sigma)$. The proposition is proved. ∎
+Now we are ready to compute the monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$.
+Theorem 4.11.
+For the generalized Frobenius manifold $\mathcal{M}_{D}=\mathcal{M}(R,\omega)$, its monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\ltimes\mathbb{Z}^{\ell}$.
+More precisely,
+$\mathrm{Mono}(\mathcal{M}_{D})\cong\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}.$ (4.50)
+Here $\mathrm{Stab}_{W}(\omega)$ stands for the stabilizer of $\omega$ in $W(R)$.
+We first prove three lemmas.
+Lemma 4.12.
+Let $\mathbb{F}\hookrightarrow\mathbb{E}$ be a field extension, $x_{1},\dots,x_{m}\in\mathbb{E}$ be some nonzero algebraic elements over $\mathbb{F}$.
+Suppose $a_{1},\dots,a_{m}\in\mathbb{E}$ such that $a_{1}x_{1}^{n}+\cdots+a_{m}x_{m}^{n}\in\mathbb{F}$ holds for any $n\in\mathbb{Z}_{>0}$, then it also holds for $n=0$, i.e., we have $a_{1}+\cdots+a_{m}\in\mathbb{F}$.
+Proof.
+Since $x_{i}\not=0$, the minimal polynomial $p_{i}$ of $x_{i}$ over $\mathbb{F}$ has a nonzero constant term (even if $x_{i}\in\mathbb{F}$).
+Let $p=p_{1}\cdots p_{m}$, then $a_{1}p(x_{1})+\cdots+a_{m}p(x_{m})=0$.
+Since $p$ has a nonzero constant term, we have $a_{1}+\cdots+a_{m}\in\mathbb{F}$.
+∎
+Lemma 4.13.
+For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we have
+$Q_{\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$ (4.51)
+Proof.
+For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we denote
+$\rho=\rho_{1}\omega_{1}+\cdots+\rho_{\ell}\omega_{\ell},\quad\rho_{j}\in\mathbb{Z}.$ (4.52)
+We first consider the case when each $\rho_{j}>0$. In what follows, we briefly call this case by $\rho>0$. Note that
+$\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\rho,\sigma(\mathbf{x})-c\omega)}=\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}})+c\sigma(\omega)-c\omega)}\in\mathscr{A}^{W}[\lambda^{\varepsilon/\kappa}],$ (4.53)
+where $\varepsilon=\mathrm{gcd}\{(\omega_{r},\alpha_{r})\mid r=1,\dots,\ell\}$. Indeed, (4.53) is $W_{a}(R)$-invariant, and equation (2.18) implies that
+$\sigma(\omega)-\omega=-\sum_{r=1}^{\ell}n_{r}\alpha_{r}$
+for some $n_{r}\in\mathbb{Z}_{\geq 0}$, which indicates that $(\rho,\sigma(\omega)-\omega)\in\varepsilon\mathbb{Z}_{\leq 0}$.
+Since $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, by putting $\lambda=0$ in (4.53), we obtain
+$\sum_{\begin{subarray}{c}\sigma\in W(R)\\ (\rho,\sigma(\omega)-\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left[z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right].$
+Therefore, we only need to show that for any $\sigma\in W(R)$, $(\rho,\sigma(\omega)-\omega)=0$ implies $\sigma(\omega)=\omega$. In fact, if $(\rho,\sigma(\omega)-\omega)=0$, we have
+$(\rho,\sigma(\omega)-\omega)=(\rho,-\sum_{r=1}^{\ell}n_{r}\alpha_{r})=\sum_{r=1}^{\ell}\rho_{r}n_{r}(\omega_{r},\alpha_{r})=0.$
+Since $\rho_{r}>0$, $n_{r}\geq 0$ and $(\omega_{r},\alpha_{r})>0$, $r=1,\dots,\ell$, we have $n_{r}=0$, $r=1,\dots,\ell$, hence $\sigma(\omega)=\omega$.
+Therefore, $Q_{\rho}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0})$ for any $\rho>0$.
+In particular, we know that for any $\rho>0$ and $\sigma\in\mathrm{Stab}_{W}(\omega)$, $\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}$ is algebraic over $\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0})$.
+In fact, since for any $n\in\mathbb{Z}_{>0}$, we have $n\rho>0$, hence
+$Q_{n\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(n\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
+In other words, all the Newton sums of $\{\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\mid\sigma\in\mathrm{Stab}_{W}(\omega)\}$ belong to the field $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$, hence so are their symmetric polynomials
+$r_{n}=\sum_{\begin{subarray}{c}I\subseteq\mathrm{Stab}_{W}(\omega)\\ \#I=n\end{subarray}}~\prod_{\sigma\in I}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
+This implies that they are algebraic over $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$.
+Now let us consider the general case.
+For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we take $\xi\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$ and $\xi>0$ such that $\rho+\xi>0$.
+Then for any $n\in\mathbb{Z}_{>0}$, we have $\rho+n\xi>0$, hence
+$Q_{\rho+n\xi}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho+n\xi,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
+Since $\mathrm{e}^{2\pi i(\xi,\sigma(\hat{\mathbf{x}}))}$ is algebraic over $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$, it follows from Lemma 4.12 that
+$Q_{\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
+Therefore, we have proved the lemma. ∎
+Lemma 4.14.
+Let $R$ be a root system, $\omega$ be a weight and $\mathscr{A}$ be the corresponding $\lambda$-Fourier polynomial ring.
+Denote by $\mathscr{K}=\mathrm{Frac}(\mathscr{A})$ the fraction field of $\mathscr{A}$, then $W(R)$ acts on $\mathscr{K}$, and
+$\mathrm{Inv}_{\mathscr{K}}\left(\mathrm{Stab}_{W}(\omega)\right)=\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda\right).$
+Here $\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))$ refers to the fixed subfield of $\mathscr{K}$ with respect to the action of $\mathrm{Stab}_{W}(\omega)$.
+In particular, the field extension $\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda)\hookrightarrow\mathscr{K}$ is Galois.
+Proof.
+From (2.31) and (2.36) we know that
+$z^{j}\bigr|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$, (4.54)
+where
+$H_{j}=\{\sigma\in W(R)\mid(\omega_{j},\omega-\sigma(\omega))=0\}.$ (4.55)
+Since for any $\phi\in\mathrm{Stab}_{W}(\omega)$, $\sigma\mapsto\sigma\phi$ induces a permutation on $H_{j}$, we have
+$z^{j}\big|_{\lambda=0}\in\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega)).$
+Thus we only need to prove that
+$\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))\subseteq\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda\right).$
+We first show that
+$\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell]^{\mathrm{Stab}_{W}(\omega)}\subseteq\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$ (4.56)
+In fact, for any $q\in\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell]^{\mathrm{Stab}_{W}(\omega)}$, we denote
+$q=\sum_{\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}}A_{\rho}\mathrm{e}^{2\pi i(\rho,\hat{\mathbf{x}})}$,
+where only finitely many $A_{\rho}$ is not zero. Since $q$ is $\mathrm{Stab}_{W}(\omega)$-invariant, we have
+[…]
+By using Lemma 4.13, we arrive at (4.56).
+We next show that
+$\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}\subseteq\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$ (4.57)
+Recall that $W(R)$ does not act on $\mathscr{A}$, since its action may introduce terms with negative power of $\lambda$.
+However, its subgroup $\mathrm{Stab}_{W}(\omega)$ acts on $\mathscr{A}$.
+In fact, for each $q\in\mathscr{A}$, we can write it as
+$q(\hat{\mathbf{x}};\lambda)=\sum_{r}\lambda^{r}q_{r}(\hat{\mathbf{x}}),\quad q_{r}\in\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell].$
+Then an element $\phi\in\mathrm{Stab}_{W}(\omega)$ acts on $q$ by
+$),\displaystyle(\phi^{*}q)(\hat{\mathbf{x}};\lambda)=\sum_{r}\lambda^{r}q_{r}(\phi(\mathbf{x})-c\omega)=\sum_{r}\lambda^{r}q_{r}(\phi(\hat{\mathbf{x}}$)), (4.58)
+where we used the fact that $\phi(\omega)=\omega$.
+Therefore, $\mathrm{Stab}_{W}(\omega)$ indeed acts on $\mathscr{A}$.
+Note that (4.58) also implies that if $q\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$, then each $q_{r}$ is also $\mathrm{Stab}_{W}(\omega)$-invariant, hence it follows from (4.56) that
+$q_{r}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}),$
+and $q\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda)$.
+Finally we show that for any $p/q\in\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))$, where $p,q\in\mathscr{A}$, we have
+$\frac{p}{q}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$
+In fact, we have
+$\frac{p}{q}=\frac{p\cdot\prod_{\phi\in\mathrm{Stab}_{W}(\omega)\setminus\{\mathrm{id}\}}\phi^{*}q}{\prod_{\phi\in\mathrm{Stab}_{W}(\omega)}\phi^{*}q}=:\frac{\tilde{p}}{\tilde{q}}$.
+Here $\tilde{q}\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$, hence $\tilde{p}\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$. Therefore, by using (4.57) we have
+$\frac{p}{q}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$
+The lemma is proved. ∎
+Proof of Theorem 4.11.
+Due to Proposition 4.10, we only need to determine all the isometric actions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ that preserve $\mathcal{M}_{D}\setminus\Sigma$, i.e.
+actions on coordinates $x^{i}$ that preserves $z^{j}\bigr|_{\lambda=0}$.
 Recall that
-$\varphi_{0}\colon V\otimes\mathbb{C}\to\mathcal{M}$ (4.40)
-is a quotient map. By eliminating a divisor, we obtain a covering map
-$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})\to\mathcal{M}\setminus(D\cup\Sigma^{\prime}).$ (4.41)
-Since $\mathcal{U}(\mathcal{M})$ is the universal covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, it is also a covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$, and the covering map is given by
-$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\colon\mathcal{U}(\mathcal{M})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime}),\quad p\mapsto(x^{1}\bigr{|}_{\lambda=0}(p),\cdots,x^{\ell}\bigr{|}_{\lambda=0}(p)).$ (4.42)
-In conclusion, we have the following commutative diagram of covering spaces:
-$\mathcal{U}(\mathcal{M})$¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathrm{pr}_{1}$$\overline{\pi}$$\mathrm{pr}_{2}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\varphi_{0}$
-All the maps in the diagram are covering maps. From the property of universal covering spaces, we have
-$\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))=\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\right).$ (4.43)
-Here $\mathrm{Deck}(X/Y)$ is the deck transformation group on $X$ over $Y$, where $X$ is a covering space over $Y$.
-Relation (4.43) implies that, every element $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ induces a deck transformation on $\mathcal{U}(\mathcal{M})$.
-We have the following lemma for $r(\gamma)$.
-Lemma 4.4.
-Let $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, and let us view $\gamma$ as a deck transformation on $\mathcal{U}(\mathcal{M})$.
-Then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ such that
-$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\circ\gamma=p\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.44)
-In other words, we have the following commutative diagram.
-$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$\exists!~{}p=r(\gamma)$$\varphi_{0}$$\varphi_{0}$$\gamma$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$
-Proof.
-Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, the map $p$ satisfying (4.44) is uniquely determined.
-Moreover, since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M})$, $p=r(\gamma)$ satisfies (4.44).
-Thus in order to prove the lemma, it suffices to verify that $r(\gamma)$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$.
-As $\gamma$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, we have
-$\varphi_{0}\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}=\varphi_{0}\circ\hat{\mathbf{x}}\Big{|}_{\lambda=0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.45)
-Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, it follows that
-$\varphi_{0}=\varphi_{0}\circ r(\gamma).$
-Therefore, $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. The lemma is proved. ∎
-Proposition 4.5.
-The group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ can be represented as
-$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Iso}(V\otimes\mathbb{C})\cap\mathrm{Deck}\bigg{(}\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\bigg{)}.$ (4.46)
-The right hand side consists of all the isometries on $V\otimes\mathbb{C}$ whose restrictions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ are deck transformations over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
-Proof.
-For any $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, $r(\gamma)$ is of course in $\mathrm{Iso}(V\otimes\mathbb{C})$.
-Moreover, by using Lemma 4.4, we know that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
-On the other hand, suppose $p$ be an isometry of $V\otimes\mathbb{C}$ and also be a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
-By the general lifting lemma of covering spaces, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M})\to\mathcal{U}(\mathcal{M})$, since
-$p_{*}\left(\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\right)_{*}\pi_{1}(\mathcal{U}(\mathcal{M}))=\{e\}.$ (4.47)
-Therefore, we have the following commutative diagram:
-$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$p$$\varphi_{0}$$\varphi_{0}$$\sigma$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$
-Thus by using Lemma 4.4, we have $p=r(\sigma)$. The proposition is proved. ∎
+$z^{j}\bigr|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$. (4.59)
+For any given element $\tilde{\phi}\in\mathrm{Iso}(V\otimes\mathbb{C})$, we decompose it into linear and the affine parts as follows:
+$\tilde{\phi}(\hat{\mathbf{x}})=\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.60)
+Then $\tilde{\phi}$ acts on the the Fourier polynomials $z^{j}$ by
+$(\tilde{\phi}^{*}z^{j})\big|_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.61)
+$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.62)
+where $H_{j}$ is defined in (4.55). Therefore, the action preserves all $z^{j}\bigr|_{\lambda=0}$, $j=1,\dots,\ell$ if and only if both of the following conditions hold:
+  1. (1)
+    For each $j=1,\dots,\ell$, $\phi$ induces a permutation on $\{\sigma^{-1}(\omega_{j})\mid\sigma\in H_{j}\}$ by right action $\sigma\mapsto\sigma\phi$.
+  2. (2)
+    for each $j=1,\dots,\ell$, $\sigma\in H_{j}$,
+    $(\sigma^{-1}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
+    is an integer.
+We now proceed to respectively show that condition (1) is equivalent to $\phi\in\mathrm{Stab}_{W}(\omega)$, and condition (2) is equivalent to $v^{1},\dots,v^{\ell}\in\mathbb{Z}$.
+Suppose $\phi\in\mathrm{Stab}_{W}(\omega)$, then $\sigma\mapsto\sigma\phi$ induces a permutation on $H_{j}$, hence on $\{\sigma^{-1}(\omega_{j})\mid\sigma\in H_{j}\}$.
+On the other hand, if condition (1) holds, we are to show that $\phi\in\mathrm{Stab}_{W}(\omega)$.
+We consider the canonical action of $\phi$ on $\mathscr{K}$ given by $\phi^{*}q(\mathbf{x},\lambda)=q(\phi(\mathbf{x}),\lambda)$.
+According to Lemma 4.14, we know that
+$\mathrm{Gal}\left(\mathscr{K}/\mathbb{C}\left(z^{1}|_{\lambda=0},\dots,z^{\ell}|_{\lambda=0};\lambda\right)\right)=\mathrm{Stab}_{W}(\omega).$
+Condition (1) implies that $\phi^{*}$ fixes $\mathbb{C}\left(z^{1}|_{\lambda=0},\dots,z^{\ell}|_{\lambda=0};\lambda\right)$, hence we only need to show that $\phi^{*}(\mathscr{K})\subseteq\mathscr{K}$.
+Note that $\mathscr{K}$ is generated by $\mathrm{e}^{2\pi ix_{1}},\dots,\mathrm{e}^{2\pi ix_{\ell}}$ and $\lambda$ as $\mathbb{C}$-algebra, and we have $\phi^{*}\lambda=\lambda$, it suffices to show that $\phi^{*}\mathrm{e}^{2\pi ix_{1}}\in\mathscr{K}$.
+In fact, by considering $\mathrm{id}\in H_{j}$, $j=1,\dots,\ell$, we know from condition (1) that
+$\phi^{*}(\omega_{j})=\sigma_{j}(\omega_{j}),\quad\exists~\sigma_{j}\in H_{j}.$
+Then in particular, for each $j=1,\dots,\ell$, we have
+$\phi^{*}\mathrm{e}^{2\pi ix^{j}}=\phi^{*}\mathrm{e}^{2\pi i(\omega_{j},\mathbf{x}-c\omega)}=\mathrm{e}^{2\pi i(\omega_{j},\phi(\mathbf{x})-c\omega)}=\mathrm{e}^{2\pi i(\sigma_{j}^{-1}(\omega_{j}),\hat{\mathbf{x}})}$.
+Here we used $(\sigma_{j}^{-1}(\omega_{j})-\omega_{j},\omega)=0$. From (2.20) we know that for each $r=1,\dots,\ell$, $(\sigma_{j}^{-1}(\omega_{j}),\alpha_{r}^{\vee})$ is an integer, so
+$\phi^{*}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi i\sum_{r}x^{r}(\sigma_{j}^{-1}(\omega_{j}),\alpha_{r}^{\vee})}\in\mathscr{A},\quad j=1,\dots,\ell.$
+It follows immediately that $\phi^{*}(\mathscr{K})\subseteq\mathscr{K}$. Hence $\phi\in\mathrm{Stab}_{W}(\omega)$.
+Suppose $v^{1},\dots,v^{\ell}\in\mathbb{Z}$, then for any $\sigma\in H_{j}$, we know from (2.20) that for each $i=1,\dots,\ell$ $(\sigma^{-1}(\omega_{j}),\alpha_{i}^{\vee})$ is an integer, so is their $\mathbb{Z}$-combination $(\sigma^{-1}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$.
+On the other hand, if condition (2) holds, then for any $j=1,\dots,\ell$, by considering $\mathrm{id}\in H_{j}$, we obtain that $v^{j}\in\mathbb{Z}$.
+Therefore, we have $\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$. ∎
+By using Proposition 4.8, we have the following corollary.
+Corollary 4.15.
+The monodromy group of $\widehat{\mathcal{M}}_{D}$ is isomorphic to a subgroup of $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$.
+In general, the monodromy group of $\widehat{\mathcal{M}}_{D}$ is not isomorphic to that of $\mathcal{M}_{D}$, and we will give a counterexample in Sect.
+5.
+However, in some cases we have $\widehat{\mathcal{M}}_{D}=\mathcal{M}_{D}$, so they have isomorphic monodromy groups.
+We will also see such examples in Sect.
+5.
 Remark 4.2.
-Note that $z^{1},\cdots,z^{\ell}$ are coordinates on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, so intuitively speaking, $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(z^{1},\cdots,z^{\ell})$.
-Similarly, $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(t^{1},\cdots,t^{\ell})$.
-From Theorem 4.2 we know that, if $t^{1},\cdots,t^{\ell}$ are preserved under some transformations, so are $z^{1},\cdots,z^{\ell}$.
-This fact implies the relation (4.39).
-We are ready to present the following theorem.
-Theorem 4.6.
-For the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$, the group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\rtimes\mathbb{Z}^{\ell}$.
-More precisely,
-$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))\cong\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.48)
-Here $\mathrm{Stab}(\omega)$ is the stabilizer of $\omega$ in $W(R)$.
-Proof.
-Due to Proposition 4.5, we only need to determine all the isometric actions on $x^{i}\bigr{|}_{\lambda=0}$ that preserve $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., preserve $z^{1},\cdots,z^{\ell}$.
-From (2.31) and (2.41) it follows that
-$z^{j}\bigr{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (4.49)
-Here
-$H_{j}=\{\sigma\in W(R)\mid(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\}$
-is a subgroup of $W(R)$. For any given element $\tilde{\phi}\in\mathrm{Iso}(E^{\ell})$, we decompose it into linear and the affine parts as follows:
-$\tilde{\phi}(\mathbf{t})=\phi(\mathbf{t})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.50)
-It acts on the the Fourier polynomials $z^{j}$ by
-$\tilde{\phi}^{*}z^{j}\big{|}_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.51)
-$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.52)
-so it preserves $z^{j}\bigr{|}_{\lambda=0}$, $j=1,\cdots,\ell$ if and only if $\phi$ induces a permutation on $H_{j}$ by right action $\sigma\to\sigma\phi$ and
-$(\sigma^{*}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
-is always an integer for each $\sigma\in H_{j}$, $j=1,\cdots,\ell$. These conditions are equivalent to
-$\phi\in\bigcap_{j=1}^{\ell}H_{j},\quad v^{1},\cdots,v^{\ell}\in\mathbb{Z}.$ (4.53)
-This implies that
-$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\bigcap_{j=1}^{\ell}H_{j}\rtimes\mathbb{Z}^{\ell}.$ (4.54)
-Now let us compute $\bigcap_{j=1}^{\ell}H_{j}$. Note that $\sigma\in H_{j}$, $j=1,\cdots,\ell$ if and only if
-$(\omega_{j},\omega)=(\omega_{j},\sigma(\omega)),\quad j=1,\cdots,\ell.$ (4.55)
-Since $\omega_{j},j=1,\cdots,\ell$ span $V\otimes\mathbb{C}$, (4.55) is equivalent to $\sigma(\omega)=\omega$. Therefore,
-$\bigcap_{j=1}^{\ell}H_{j}=\mathrm{Stab}(\omega).$ (4.56)
-It follows from (4.54) and (4.56) that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$. The theorem is proved. ∎
-By using (4.39), we have the following corollary.
-Corollary 4.7.
-The monodromy group of the generalized Frobenius manifold $\mathcal{M}(R,\omega)$ is isomorphic to a subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
-It comes out a natural question to ask: is $\mathrm{Mono}(\mathcal{M}(R,\omega))$ exactly isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$?
-Unfortunately, the answer to this question is negative, and we will see counterexamples in Sec.
-5.
-However, in some special cases, the isomorphism holds true.
-One simple case is when ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$.
-In this case we have ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}\cong(\mathcal{M}\setminus D)\setminus\Sigma^{\prime}$, hence
-$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))=r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.57)
-We will also see examples which belong to this case in Sec. 5.
+The group $\mathrm{Stab}_{W}(\omega)$ is a parabolic subgroup of $W(R)$ [6]. Explicitly, we have
+$\mathrm{Stab}_{W}(\omega)=W_{S^{c}}=\langle\sigma_{i}\mid i\in S^{c}\rangle,$ (4.63)
+where $S$ is given by (2.29) and $S^{c}$ is its complement in $\{1,\cdots,\ell\}$.
 5. Examples
-[…]
+In this section, we give some typical examples for our construction.
+We assume that $e_{1},\dots,e_{n}$ give an orthonormal basis of the Euclidean space $\mathbb{R}^{n}$ for $n\in\mathbb{N}$ with inner product (,)($\cdot\,$$\cdot$), and the $\ell$-dimensional space $V$ will be $\mathbb{R}^{\ell}$ or a subspace of $\mathbb{R}^{n}$ for a certain $n>\ell$ endowed with the induced inner product (,)($\cdot\,$$\cdot$).
+For convenience of the presentation, we will use $x_{i},y_{i},z_{i},t_{\alpha}$ to denote $x^{i},y^{i},z^{i}$ and $t^{\alpha}$ respectively.
+We will also use $E$ to denote the Euler vector $\tilde{E}$ of the generalized Frobenius manifolds given in (3.53).
 We start with the simplest example.
 Example 5.1 ($(R,\omega)=(A_{1},\omega_{1})$).
-Let $R$ be the root system of type $A_{1}$, $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
+Let $R$ be the root system of type $A_{1}$, and $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
 $\alpha_{1}=e_{1}-e_{2}.$
@@
 $\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1},\quad\mathbf{x}\in V.$
-In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence
-$\lambda=\mathrm{e}^{-2\pi ic}.$
-The Weyl group $W(R)$ acting on $V$ is generated by the reflections
-$\sigma_{1}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{1},\mathbf{x})\alpha_{1},\quad\mathbf{x}\in V.$
-Then we have
+In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence $\lambda=\mathrm{e}^{-2\pi ic}$. The Weyl group $W(R)$ acting on $V$ is generated by
 $\sigma_{1}(\mathbf{x})=c\omega_{1}+(-c-x_{1})\alpha^{\vee}_{1},\quad x_{1}\mapsto-c-x_{1}.$
 The invariant Fourier polynomial has the form
-$Y_{1}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
-and we take
-$y_{1}=e^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
-We have
-$(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2},$
-and
+$Y_{1}=\mathrm{e}^{2\pi i\left(\omega_{1},\mathbf{x}\right)}+\mathrm{e}^{2\pi i\left(\omega_{1},\sigma_{1}(\mathbf{x})\right)}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
+hence
+$y_{1}=\mathrm{e}^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
+The metric (,)($\cdot\,$$\cdot$) on $V$ is given by $(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2}$, hence
 $g^{11}_{\lambda}=\frac{1}{4\pi^{2}}(\operatorname{d\!}y_{1},\operatorname{d\!}y_{1})=\frac{1}{4\pi^{2}}\frac{\partial y_{1}}{\partial x_{1}}(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})\frac{\partial y_{1}}{\partial x_{1}}=-\frac{1}{2}y_{1}^{2}+2\lambda.$
-Therefore, here $\{y_{1}\}$ is a pencil generator, and we obtain a flat pencil
+Therefore, $y_{1}$ is a pencil generator, and we obtain a flat pencil
 $\eta^{11}=2,\quad g^{11}=-\frac{1}{2}y_{1}^{2}.$
-Now we consider the flat coordinate $t_{1}=y_{1}/\sqrt{2}$ of $\eta$, then in the new coordinate the flat pencil has the form
+The metric $\eta$ has the flat coordinate $t_{1}=y_{1}/\sqrt{2}$. In this new coordinate the flat pencil has the form
 $\eta^{11}=1,\quad g^{11}=-\frac{1}{2}t_{1}^{2}.$
@@
 Moreover, the flat coordinate of $g$ is given by
-$x^{1}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}(\log(\sqrt{2}t_{1})).$
-It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $t_{1}$. Hence the monodromy group of the generalized Frobenius manifold is given by
-$\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathbb{Z}.$
-We considered the construction of generalized Frobenius manifolds for different types of affine Weyl groups with appropriately chosen weights.
+$x^{1}\big|_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}\log(\sqrt{2}\,t_{1}).$
+It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $y_{1}$, and equivalently $t_{1}$, hence the monodromy groups of $\mathcal{M}_{D}$ and $\widehat{\mathcal{M}}_{D}$ are given by
+$\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathbb{Z}.$
+Indeed, $\mathrm{Stab}_{W}(\omega_{1})$ is trivial and we have $\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathrm{Stab}_{W}(\omega_{1})\ltimes\mathbb{Z}$.
+We have applied the above-mentioned construction of generalized Frobenius manifolds to different types of affine Weyl groups with appropriately chosen weights.
 In what follows we will give some representative examples, which are chosen based on the following points:
   1. 1.
-    In Example 5.2, the basic generators $y^{1},\cdots,y^{\ell}$ are pencil generators. The covering space ¯$\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is trivial, thus
-    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$
+    In Example 5.2, the basic generators $y_{1},y_{2}$ are pencil generators, and
+    $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathrm{Mono}(\mathcal{M}_{D})~\cong~\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}.$
   2. 2.
-    In Example 5.3, the basic generators $y^{1},\cdots,y^{\ell}$ are also pencil generators. However, unlike the previous example, the monodromy group
-    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{\ell}$
-    is a proper subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
+    In Example 5.3, the basic generators $y_{1},y_{2},y_{3}$ are also pencil generators, while $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is a proper subgroup of $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$.
   3. 3.
-    In Example 5.4, the basic generators $y^{1},y^{2},y^{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y^{1},y^{2}$ and $y^{3}$, i.e.,
+    In Example 5.4, the basic generators $y_{1},y_{2},y_{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y_{1},y_{2}$ and $y_{3}$, i.e.,
     $z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=y_{3}.$
-    Here $a_{1},a_{2}$ are some constants, and we find three possible values for them.
+    Here $a_{1},a_{2}$ are some constants, and there are three possible values for them.
   4. 4.
-    In Example 5.5, $y^{1},y^{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
+    In Example 5.5, $y_{1},y_{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
     $z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
   5. 5.
-    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight.
-This complicates the construction of pencil generators and the generalized Frobenius manifold structure.
-These examples involve most issues that we may come across in the computations.
-Something interesting issues will also happen when we consider other types of affine Weyl groups.
-More details will be presented in our subsequent publications.
+    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight. This complicates the construction of pencil generators.
+  6. 6.
+    In Example 5.7, we present a case with much more complicated construction of pencil generators.
+Compared with the former ones, the resulting prepotential in this case is not a Laurant polynomial in flat coordinates.
+More examples will be presented in our subsequent publications [20].
 Example 5.2 ($(R,\omega)=(A_{2},\omega_{2})$).
@@
 $\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
-We choose $\omega=\omega_{2}$. Let us introduce the coordinates $x_{1},x_{2}$ of $V$ by
+Choose $\omega=\omega_{2}$, and introduce the coordinates $x_{1},x_{2}$ of $V$ by
 $\mathbf{x}=c\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
-The Weyl group $W(R)$ acting on $V$ is generated by the reflections
-$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
-The invariant $\lambda$-Fourier polynomials are given by
+We have the invariant Fourier polynomials
 $Y_{1}$ $=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{2}-x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{2}+\frac{2}{3}c\right)},$
 $Y_{2}$ $=\mathrm{e}^{2\pi i\left(x_{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{1}-x_{2}-\frac{1}{3}c\right)}.$
-Since $\kappa=1$, we have
-$\lambda=\mathrm{e}^{-2\pi ic}.$ (5.1)
-We take
+Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$ and the invariant $\lambda$-Fourier polynomials
 $y_{1}$ $=\lambda^{\frac{1}{3}}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\lambda\mathrm{e}^{-2\pi ix_{2}},$
 $y_{2}$ $=\lambda^{\frac{2}{3}}Y_{2}=\mathrm{e}^{2\pi ix_{2}}+\lambda\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}.$
-The metric on $V$ that is induced from $\mathbb{R}^{3}$ is given by
+The metric on $V$ is given by
 $\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix},\quad(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})={\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}}^{-1}=\frac{1}{3}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
@@
 $\left(g_{\lambda}^{ij}\right)=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{2}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&3\lambda-\frac{1}{3}y_{1}y_{2}\\[3.0pt] 3\lambda-\frac{1}{3}y_{1}y_{2}&2\lambda y_{1}-\frac{2}{3}y_{2}^{2}\end{pmatrix}$.
-Then $y^{1},y^{2}$ are pencil generators, and the corresponding flat pencil is given by
+Then $\{y_{1},y_{2}\}$ is a set of pencil generators, and the corresponding flat pencil is given by
 $\left(\eta^{ij}\right)=\begin{pmatrix}0&3\\ 3&2y_{1}\end{pmatrix},\quad\left(g^{ij}\right)=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&-\frac{1}{3}y_{1}y_{2}\\[3.0pt] -\frac{1}{3}y_{1}y_{2}&-\frac{2}{3}y_{2}^{2}\end{pmatrix}$.
-The metric $\eta$ has the flat coordinates
+The metric $\eta$ has flat coordinates
 $t_{1}=\frac{1}{3}y_{1},\quad t_{2}=y_{2}-\frac{1}{6}y_{1}^{2},\quad$
@@
 $e=-\frac{2}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{1}}-\frac{6t_{1}}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{2}}.$
-We also have $e^{\flat}=-\mathrm{d}\log y_{2}$.
-Note that in this example, not only $y_{1},y_{2}$ are polynomials of $t^{1}$ and $t^{2}$, but also $t^{1},t^{2}$ are polynomials of $y_{1}$ and $y_{2}$.
-This implies that the monodromy between $t^{1}$, $t^{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$.
+We also have $e=-\mathrm{grad}_{\eta}\log y_{2}$.
+Note that in this example, not only are $y_{1},y_{2}$ polynomials of $t_{1}$ and $t_{2}$, but $t_{1}$ and $t_{2}$ are also polynomials of $y_{1}$ and $y_{2}$.
+This implies that the monodromy between $t_{1}$, $t_{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., $\widehat{\mathcal{M}}_{D}=\mathcal{M}_{D}$.
 Therefore,
-$\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))=\mathrm{Stab}(\omega_{2})\rtimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\rtimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
+$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Stab}_{W}(\omega_{2})\ltimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\ltimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\ltimes\mathbb{Z}^{2}.$
 More explicitly, $\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))$ is generated by
-$\sigma_{1}\colon$ $x_{1}\to x_{2}-x_{1},\quad x_{2}\to x_{2},$
-$T_{1}\colon$ $x_{1}\to x_{1}+1,\quad x_{2}\to x_{2},$
-$T_{2}\colon$ $x_{1}\to x_{1},\quad x_{2}\to x_{2}+1.$
+$\sigma_{1}\colon$ $x_{1}\mapsto x_{2}-x_{1},\quad x_{2}\mapsto x_{2},$
+$T_{1}\colon$ $x_{1}\mapsto x_{1}+1,\quad x_{2}\mapsto x_{2},$
+$T_{2}\colon$ $x_{1}\mapsto x_{1},\quad x_{2}\mapsto x_{2}+1.$
 Example 5.3 ($(R,\omega)=(C_{3},\omega_{3})$).
@@
 Then we have the coroots
-$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3}.$
-The fundamental weights are given by
+$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3},$
+and the fundamental weights
 $\omega_{1}=\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\alpha^{\vee}_{3},\quad\omega_{2}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+2\alpha^{\vee}_{3},\quad\omega_{3}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+3\alpha^{\vee}_{3}.$
@@
 $\mathbf{x}=c\omega_{3}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
-Since $\kappa=2$, we have
-$\lambda=\mathrm{e}^{-4\pi ic},$
-then the $\lambda$-Fourier invariant polynomials are given by
+Since $\kappa=2$, we have $\lambda=\mathrm{e}^{-4\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
 $y_{1}=\lambda^{\frac{1}{2}}(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{3}{2}}\xi_{1}\xi_{2}\xi_{3},$
@@
 $\xi_{3}$ $=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi i(x_{3}-x_{2})}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi i(x_{3}-x_{2})}.$
-The metric (,)(\ ,\ )( , ) has the form
+The metric on $V$ has the form
 $\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&1\end{pmatrix}^{-1}=\begin{pmatrix}1&1&1\\ 1&2&2\\ 1&2&3\end{pmatrix}$.
@@
 The flat coordinates $t_{1},t_{2},t_{3}$ of $\eta$ are given by the polynomials
-$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{2}^{3}+t_{1}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
+$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{1}^{3}+t_{2}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
 In these flat coordinates the metric $\eta$ has the form
@@
 $e=$ $\,\frac{3}{2\left(t_{1}^{3}+t_{2}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right)}$
-×$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
-We also have $e^{\flat}=-\frac{1}{2}\mathrm{d}\log y_{3}$.
+$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
+We also have $e=-\frac{1}{2}\mathrm{grad}_{\eta}\log y_{3}$.
 Note that, if we take the flat coordinates $v_{1},v_{2},v_{3}$ of $\eta$ via a linear transformation
@@
 the Euler vector field and the unit vector field have the form
-$E=\frac{1}{2}v^{1}\frac{\partial}{\partial v^{1}}+\frac{1}{2}v^{2}\frac{\partial}{\partial v^{2}}+\frac{1}{2}v^{3}\frac{\partial}{\partial v^{3}},\quad e=-\frac{1}{2v^{1}}\frac{\partial}{\partial v^{1}}-\frac{1}{2v^{2}}\frac{\partial}{\partial v^{2}}-\frac{1}{2v^{3}}\frac{\partial}{\partial v^{3}}.$
+$E=\frac{1}{2}v_{1}\frac{\partial}{\partial v_{1}}+\frac{1}{2}v_{2}\frac{\partial}{\partial v_{2}}+\frac{1}{2}v_{3}\frac{\partial}{\partial v_{3}},\quad e=-\frac{1}{2v_{1}}\frac{\partial}{\partial v_{1}}-\frac{1}{2v_{2}}\frac{\partial}{\partial v_{2}}-\frac{1}{2v_{3}}\frac{\partial}{\partial v_{3}}.$
 It is clear that, up to a factor, we have the isomorphism
@@
 $x_{1}\big|_{\lambda=0}=$ $\frac{1}{2\pi i}\log(2v_{1}),\quad x_{2}\big|_{\lambda=0}=\frac{1}{2\pi i}\left(\log(2v_{2})+\log(2v_{1})\right),$
-$x_{3}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\left(\frac{1}{2}\log(2v_{3})+\frac{1}{2}\log(2v_{2})+\frac{1}{2}\log(2v_{1})\right),$
-we know that the monodromy group is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
-$\mathrm{Mono}(\mathcal{M}(C_{3},\omega_{3}))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{3}~{}<~{}\mathrm{Stab}(\omega_{3})\rtimes\mathbb{Z}^{3}.$ (5.3)
-Here $\mathrm{Stab}(\omega_{3})$ is nontrivial since $\sigma_{1},\sigma_{2}\in\mathrm{Stab}(\omega_{3})$.
-This example shows that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ is not always isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
+$x_{3}\big|_{\lambda=0}=$ $\frac{1}{4\pi i}\left(\log(2v_{3})+\log(2v_{2})+\log(2v_{1})\right),$
+we know that the monodromy group of $\widehat{\mathcal{M}}_{D}$ is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
+$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})~\cong~\{e\}\ltimes\mathbb{Z}^{3}~<~\mathrm{Stab}_{W}(\omega_{3})\ltimes\mathbb{Z}^{3}.$ (5.2)
+Here $\mathrm{Stab}_{W}(\omega_{3})=\langle\sigma_{1},\sigma_{2}\rangle\cong S_{3}$ is not trivial.
+In fact, in this case we have $\widehat{\mathcal{M}}_{D}/S_{3}\cong\mathcal{M}_{D}$.
+This example shows that $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is not isomorphic to $\mathrm{Mono}(\mathcal{M}_{D})$ in general.
 Example 5.4 ($(R,\omega)=(B_{3},\omega_{1})$).
@@
 $\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
-Since $\kappa=1$, we have
-$\lambda=\mathrm{e}^{-2\pi ic}.$
-The $\lambda$-Fourier invariant polynomials are given by
+Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
 $y_{1}=\lambda(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{1}{2}}\tilde{\xi}_{1}\tilde{\xi}_{2}\tilde{\xi}_{3}.$
@@
 $\tilde{\xi}_{3}$ $=e^{\pi i(2x_{3}-x_{2})}+e^{-\pi i(2x_{3}-x_{2})}.$
-The metric (,)(\ ,\ )( , ) has the form
+The metric on $V$ has the form
 $\left((dx_{i},dx_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-2\\ 0&-2&4\end{pmatrix}^{-1}=\begin{pmatrix}1&1&\frac{1}{2}\\ 1&2&1\\ \frac{1}{2}&1&\frac{3}{4}\end{pmatrix}$.
-In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. Let us consider the proper generators with the form
-$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=z_{3},$
-here $a_{1},a_{2}$ are two constants. Then there are three possible choices of $a_{1}$ and $a_{2}$ such that $z_{1},z_{2}$ and $z_{3}$ form a set of pencil generators, they are given by
-$a_{1}=6,\quad a_{2}=-12.$ (5.4)
+In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. A set of proper generators should have the form
+$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=y_{3},$
+where $a_{1},a_{2}$ are two constants. We have the following three choices for $a_{1}$ and $a_{2}$:
+$a_{1}=6,\quad a_{2}=-12,$ (5.3)
 $a_{1}=2,\quad a_{2}=4,$ (5.4)
@@
 In these coordinates we have
-$(\eta^{ij})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}.$
-We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},1)$ with potential
+$(\eta^{\alpha\beta})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}$.
+We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{1}$ with potential
 $F=$ $\,\frac{t_{1}^{2}t_{2}}{3t_{3}}-\frac{t_{1}^{2}t_{3}^{2}}{12}-\frac{2t_{1}t_{2}^{3}}{27t_{3}^{2}}-\frac{1}{36}t_{1}t_{2}^{2}t_{3}+\frac{1}{360}t_{1}t_{2}t_{3}^{4}-\frac{t_{1}t_{3}^{7}}{30240}$
@@
 $e$ $=-\frac{6}{D}\left(\left(40t_{1}+20t_{2}t_{3}^{2}+t_{3}^{5}\right)\frac{\partial}{\partial t_{1}}+\left(40t_{2}+20t_{3}^{3}\right)\frac{\partial}{\partial t_{2}}+40t_{3}\frac{\partial}{\partial t_{3}}\right).$
-Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
+Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e=-\mathrm{grad}_{\eta}\log z_{1}$.
 Now we consider the $a_{1}=2,a_{2}=4$ case. We have
@@
 In these coordinates we have
-$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
-We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},2)$ with potential
+$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$.
+We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{2}$ with potential
 $F=$ $\,-\frac{t_{2}^{4}}{48}+\frac{t_{1}^{3}}{24t_{3}}-\frac{1}{4}t_{1}t_{2}^{2}t_{3}-\frac{t_{1}^{2}t_{3}^{2}}{12}+\frac{t_{2}^{2}t_{3}^{4}}{24}+\frac{t_{1}t_{3}^{5}}{180}-\frac{t_{3}^{8}}{2268}.$
@@
 $e$ $=\frac{1}{3t_{2}^{2}-6t_{1}t_{3}-2t_{3}^{4}}\left(\left(6t_{1}+8t_{3}^{3}\right)\frac{\partial}{\partial t_{1}}-6t_{2}\frac{\partial}{\partial t_{2}}+6t_{3}\frac{\partial}{\partial t_{3}}\right).$
-We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
+We also have $e=-\mathrm{grad}_{\eta}\log z_{1}$.
 Finally, let us consider the $a_{1}=-2,a_{2}=4$ case. We have
@@
 In these coordinates we have
-$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix},$
-and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$. In all cases, the monodromy groups are given by
-$\mathrm{Mono}(\mathcal{M}(B_{3},\omega_{1},k))=\mathrm{Stab}(\omega_{1})\rtimes\mathbb{Z}^{3}=\langle\sigma_{2},\sigma_{3}\rangle\rtimes\mathbb{Z}^{3}\cong D_{4}\rtimes\mathbb{Z}^{3},\quad k=1,2,3.$
+$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$,
+and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$, hence the third generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{3}$ is isomorphic to the second one.
 Example 5.5 ($(R,\omega)=(G_{2},\omega_{2})$).
@@
 $\mathbf{x}=c\,\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}=(x_{1}+3c)\alpha^{\vee}_{1}+(x_{2}+6c)\alpha^{\vee}_{2},$
-Since $\kappa=3$, we have
-$\lambda=\mathrm{e}^{-6\pi ic}.$
-The $\lambda$-Fourier invariant polynomials are given by
-$y_{1}=$ $\,\mathrm{e}^{-6\pi ic}\left(\mathrm{e}^{2\pi i(x_{1}+3c)}+\mathrm{e}^{-2\pi i(x_{1}+3c)}+\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}\right.$
-$\,\left.+\mathrm{e}^{2\pi i(x_{1}-x_{2}-3c)}+\mathrm{e}^{-2\pi i(x_{1}-x_{2}-3c)}\right)$
-$=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}$
-$+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
-$y_{2}=$ $\,e^{-12\pi ic}\left(\mathrm{e}^{2\pi i(x_{2}+6c)}+\mathrm{e}^{-2\pi i(x_{2}+6c)}+\mathrm{e}^{2\pi i(3x_{1}-2x_{2}-3c)}+\mathrm{e}^{-2\pi i(3x_{1}-2x_{2}-3c)}\right.$
-$\,\left.+\mathrm{e}^{2\pi i(x_{2}-3x_{1}-3c)}+\mathrm{e}^{-2\pi i(x_{2}-3x_{1}-3c)}\right)$
-$=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}$
-$\,+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
-The metric (,)(\ ,\ )( , ) has the form
+Since $\kappa=3$, we have $\lambda=\mathrm{e}^{-6\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
+$y_{1}=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
+$y_{2}=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
+The metric on $V$ has the form
 $\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1\\ -1&\frac{2}{3}\end{pmatrix}^{-1}=\begin{pmatrix}2&3\\[3.0pt] 3&6\end{pmatrix}$.
 In the $y_{1},y_{2}$ coordinates the metric does not depend at most linearly on $\lambda$, hence $y_{1}$ and $y_{2}$ are not pencil generators.
-Proper generators $z_{1}$ and $z_{2}$ may be introduced by
+Proper generators $z_{1}$ and $z_{2}$ can be introduced by
 $z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
 Then in the $z_{1},z_{2}$ coordinates we have
-$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$
+$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$.
 The metric $\eta$ has flat coordinates
@@
 $e$ $=-\frac{2}{3t_{2}\left(4t_{1}^{2}+81t_{2}^{4}\right)}\left((4t_{1}^{2}+243t_{2}^{4})\frac{\partial}{\partial t_{1}}+4t_{1}t_{2}\frac{\partial}{\partial t_{2}}\right).$
-In this case, $\mathrm{Stab}(\omega_{1})=\langle\sigma_{1}\rangle$, where
-$\sigma_{1}(x_{1},x_{2})=(-x_{1}+x_{2},x_{2}).$
-However, one has
-$t_{1}=\frac{z_{1}}{2\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}},\quad t_{2}=-\frac{1}{3}\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}.$
-Hence the actions of $\sigma_{1}$ on $t_{1}$ and $t_{2}$ is not trivial. In fact,
-[…]
-Here the value of $3]{-1}$ depends on the choice of the monodromy region of $f(z)=z^{1/3}$. Therefore we have
-$\mathrm{Mono}(\mathcal{M}(G_{2},\omega_{2}))=\{e\}\rtimes\mathbb{Z}^{2}.$
+We also have $e=-\frac{1}{3}\mathrm{grad}_{\eta}\log z_{2}$.
 Example 5.6 ($(R,\omega)=(A_{3},\omega_{1}+\omega_{3})$).
@@
 $\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}-e_{4}.$
-We have $\alpha^{\vee}_{i}=\alpha_{i},i=1,2,3$, and the fundamental weights are given by
+We have $\alpha^{\vee}_{i}=\alpha_{i},~i=1,2,3$, and the fundamental weights are given by
 $\omega_{1}=\frac{3}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{1}{4}\alpha^{\vee}_{3},\quad\omega_{2}=\frac{1}{2}\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{1}{2}\alpha^{\vee}_{3},\quad\omega_{3}=\frac{1}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{3}{4}\alpha^{\vee}_{3}.$
@@
 $\mathbf{x}=c(\omega_{1}+\omega_{3})+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3},\quad\mathbf{x}\in V.$
-Since $\kappa=1$, we have
-$\lambda=\mathrm{e}^{-2\pi ic}.$
-We have the $\lambda$-invariant Fourier polynomials
+Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. We have the invariant $\lambda$-Fourier polynomials
 $y_{1}=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{2})}+\lambda^{2}\mathrm{e}^{-2\pi ix_{3}}+\lambda\mathrm{e}^{-2\pi i(x_{2}-x_{3})},$
@@
 $\,+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{3})}+\mathrm{e}^{2\pi i(x_{1}-x_{2}+x_{3})},$
-$y_{3}=$ $\,\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{2\pi ix_{3}}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
-The metric on $V$ that is induced from $\mathbb{R}^{4}$ is given by
+$y_{3}=$ $\,\mathrm{e}^{2\pi ix_{3}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
+The metric on $V$ is given by
 $\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix},\quad\left(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j}\right)={\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)}^{-1}=\frac{1}{4}\begin{pmatrix}3&2&1\\ 2&4&2\\ 1&2&3\end{pmatrix}$.
@@
 $z_{3}$ $=\frac{1}{6}t_{1}^{4}-t_{1}^{2}t_{2}+\frac{1}{2}t_{2}^{2}+t_{1}t_{3};$
-in these coordinates the metrics $\eta$ has the form
-$\left(\eta^{ij}\right)$ $=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
+in these coordinates the metric $\eta$ has the form
+$\left(\eta^{\alpha\beta}\right)$ $=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$.
 Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{3},\omega_{1}+\omega_{3})$ with potential
@@
 and the unit vector field
-$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{3}t_{1}^{2}-36t_{2}^{2}t_{1}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{2}t_{3}t_{1}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right.$
-$\,\left.+\left(-8t_{1}^{7}-60t_{3}t_{1}^{4}+120t_{2}^{2}t_{1}^{3}-72t_{3}^{2}t_{1}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$
+$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{1}^{2}t_{3}-36t_{1}t_{2}^{2}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{1}t_{2}t_{3}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right$.
+$\,\left.+\left(-8t_{1}^{7}-60t_{1}^{4}t_{3}+120t_{1}^{3}t_{2}^{2}-72t_{1}t_{3}^{2}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$
 where
-$D=\left(t_{1}^{4}-6t_{2}t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{2}\ t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right).$
-We also have $e^{\flat}=-(\mathrm{d}\log z_{1}+\mathrm{d}\log z_{3})$. The monodromy group is given by
-$\mathrm{Mono}(\mathcal{M}(A_{3},\omega_{1}+\omega_{3}))=\mathrm{Stab}(\omega_{1}+\omega_{3})=\langle\sigma_{2}\rangle\rtimes\mathbb{Z}^{3}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
+$D=\left(t_{1}^{4}-6t_{1}^{2}t_{2}+6t_{1}t_{3}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{1}^{2}t_{2}+6t_{1}t_{3}+3t_{2}^{2}\right).$
+We also have $e=-\mathrm{grad}_{\eta}(\log z_{1}+\log z_{3})$.
+Example 5.7 ($(R,\omega)=(D_{4},\omega_{2})$).
+Let $R$ be the root system of type $D_{4}$, and $V=\mathbb{R}^{4}$. We take the simple roots
+$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}-e_{4},\quad\alpha_{4}=e_{3}+e_{4}.$
+Then we have the coroots $\alpha_{i}^{\vee}=\alpha_{i},~i=1,2,3$, and fundamental weights
+$\omega_{1}=$ $~\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\frac{1}{2}\alpha_{3}^{\vee}+\frac{1}{2}\alpha_{4}^{\vee},\quad\omega_{2}=\alpha_{1}^{\vee}+2\alpha_{2}^{\vee}+\alpha_{3}^{\vee}+\alpha_{4}^{\vee},$
+$\omega_{3}=$ $~\frac{1}{2}\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\alpha_{3}^{\vee}+\frac{1}{2}\alpha_{4}^{\vee},\quad\omega_{4}=\frac{1}{2}\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\frac{1}{2}\alpha_{3}^{\vee}+\alpha_{4}^{\vee}.$
+Let us take $\omega=\omega_{2}$ and introduce coordinates $x_{1},x_{2}$ of $V$ by
+$\mathbf{x}=c\,\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}+x_{4}\alpha^{\vee}_{4}.$
+Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. We have the invariant $\lambda$-Fourier polynomials
+$y_{1}=$ $~\lambda(\xi_{1}+\xi_{2}+\xi_{3}+\xi_{4}),$
+$y_{2}=$ $~\lambda^{2}(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{1}\xi_{4}+\xi_{2}\xi_{3}+\xi_{2}\xi_{4}+\xi_{3}\xi_{4}),$
+$y_{3}=$ $~\frac{\lambda}{2}(\psi_{1}\psi_{2}\psi_{3}\psi_{4}-\phi_{1}\phi_{2}\phi_{3}\phi_{4}),$
+$y_{4}=$ $~\frac{\lambda}{2}(\psi_{1}\psi_{2}\psi_{3}\psi_{4}+\phi_{1}\phi_{2}\phi_{3}\phi_{4}).$
+Here we denote $v_{1}=x_{1}+c$, $v_{2}=x_{2}-x_{1}+c$, $v_{3}=x_{4}+x_{3}-x_{2}$, $v_{4}=x_{4}-x_{3}$, and
+$\xi_{k}=\mathrm{e}^{2\pi iv_{k}}+\mathrm{e}^{-2\pi iv_{k}},\quad\psi_{k}=\mathrm{e}^{\pi iv_{k}}+\mathrm{e}^{-\pi iv_{k}},\quad\phi_{k}=\mathrm{e}^{\pi iv_{k}}-\mathrm{e}^{-\pi iv_{k}},\quad k=1,2,3,4.$
+The metric on $V$ is given by
+$\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)=\left(\begin{array}[]{cccc}2&-1&0&0\\ -1&2&-1&-1\\ 0&-1&2&0\\ 0&-1&0&2\\ \end{array}\right),\quad\left(\mathrm{d}x_{i},\mathrm{d}x_{j}\right)=\left(\begin{array}[]{cccc}1&1&\frac{1}{2}&\frac{1}{2}\\ 1&2&1&1\\ \frac{1}{2}&1&1&\frac{1}{2}\\ \frac{1}{2}&1&\frac{1}{2}&1\\ \end{array}\right).$
+In the coordinates $y_{1},\dots,y_{4}$ we have
+$\left(g_{\lambda}^{ij}\right)=\frac{1}{4\pi^{2}}\left(\begin{array}[]{cccc}16\lambda^{2}-y_{1}^{2}+2y_{2}&3\lambda y_{3}y_{4}-y_{1}y_{2}&4\lambda y_{4}-\frac{1}{2}y_{1}y_{3}&4\lambda y_{3}-\frac{1}{2}y_{1}y_{4}\\ 3\lambda y_{3}y_{4}-y_{1}y_{2}&g_{\lambda}^{22}&3\lambda y_{1}y_{4}-y_{2}y_{3}&3\lambda y_{1}y_{3}-y_{2}y_{4}\\ 4\lambda y_{4}-\frac{1}{2}y_{1}y_{3}&3\lambda y_{1}y_{4}-y_{2}y_{3}&16\lambda^{2}-y_{3}^{2}+2y_{2}&4\lambda y_{1}-\frac{1}{2}y_{3}y_{4}\\ 4\lambda y_{3}-\frac{1}{2}y_{1}y_{4}&3\lambda y_{1}y_{3}-y_{2}y_{4}&4\lambda y_{1}-\frac{1}{2}y_{3}y_{4}&16\lambda^{2}-y_{4}^{2}+2y_{2}\\ \end{array}\right),$
+where $g_{\lambda}^{22}=-64\lambda^{4}+4\lambda^{2}y_{1}^{2}+4\lambda^{2}y_{3}^{2}+4\lambda^{2}y_{4}^{2}-24\lambda^{2}y_{2}+2\lambda y_{1}y_{3}y_{4}-2y_{2}^{2}$. We introduce the pencil generators
+$z_{1}=y_{1}+8\lambda,\quad z_{2}=y_{2}+24\lambda^{2}+\lambda(2y_{1}+2y_{3}-2y_{4}),\quad z_{3}=y_{3}+8\lambda,\quad z_{4}=y_{4}-8\lambda.$
+In the new coordinates, we obtain a flat pencil of metrics $(g^{ij})+\lambda(\eta^{ij})$. The flat coordinates $t_{\alpha}$ of $\eta^{ij}$ are given by polynomials
+$z_{1}$ $=\frac{8}{3}t_{1}^{4}+2t_{1}t_{4}+\frac{8}{3}t_{2}^{4}+2t_{2}t_{3},$
+$z_{2}$ $=\frac{4}{9}t_{1}^{8}-\frac{4}{3}t_{1}^{5}t_{4}+\frac{56}{9}t_{1}^{4}t_{2}^{4}+\frac{20}{3}t_{1}^{4}t_{2}t_{3}+t_{1}^{2}t_{4}^{2}+\frac{20}{3}t_{1}t_{2}^{4}t_{4}+2t_{1}t_{2}t_{3}t_{4}+\frac{4}{9}t_{2}^{8}+t_{2}^{2}t_{3}^{2}-\frac{4}{3}t_{2}^{5}t_{3},$
+$z_{3}$ $=-\frac{4}{3}t_{1}^{4}+8t_{1}^{2}t_{2}^{2}+2t_{1}t_{4}+2t_{2}t_{3}-\frac{4}{3}t_{2}^{4},$
+$z_{4}$ $=\frac{4}{3}t_{1}^{4}+8t_{1}^{2}t_{2}^{2}-2t_{1}t_{4}+\frac{4}{3}t_{2}^{4}-2t_{2}t_{3}.$
+In these coordinates the metric $\eta$ has the form
+$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\end{pmatrix}$.
+Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{3},\omega_{1}+\omega_{3})$ with potential
+$F$ $=\frac{1}{t_{1}^{4}-t_{2}^{4}}\Bigg(-\frac{t_{1}^{12}}{1134}-\frac{t_{1}^{9}t_{4}}{180}+\frac{t_{1}^{8}t_{2}t_{3}}{36}-\frac{13t_{1}^{8}t_{2}^{4}}{1134}-\frac{t_{1}^{6}t_{4}^{2}}{24}+\frac{t_{1}^{5}t_{2}^{4}t_{4}}{30}-\frac{t_{1}^{5}t_{2}t_{3}t_{4}}{12}+\frac{t_{1}^{4}t_{3}^{3}}{96t_{2}}-\frac{t_{1}^{4}t_{2}^{2}t_{3}^{2}}{24}$
+$\quad-\frac{t_{1}^{4}t_{2}^{5}t_{3}}{30}+\frac{13t_{1}^{4}t_{2}^{8}}{1134}-\frac{t_{1}^{3}t_{4}^{3}}{96}+\frac{t_{1}^{2}t_{2}t_{3}t_{4}^{2}}{16}+\frac{t_{1}^{2}t_{2}^{4}t_{4}^{2}}{24}+\frac{t_{1}t_{2}^{5}t_{3}t_{4}}{12}-\frac{t_{1}t_{2}^{2}t_{3}^{2}t_{4}}{16}-\frac{t_{1}t_{2}^{8}t_{4}}{36}$
+$\quad+\frac{t_{2}^{3}t_{3}^{3}}{96}+\frac{t_{2}^{6}t_{3}^{2}}{24}+\frac{t_{2}^{9}t_{3}}{180}+\frac{t_{2}^{12}}{1134}-\frac{t_{2}^{4}t_{4}^{3}}{96t_{1}}\Bigg).$
+the Euler vector field
+$E=\frac{1}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{4}t_{2}\frac{\partial}{\partial t_{2}}+\frac{3}{4}t_{3}\frac{\partial}{\partial t_{3}}+\frac{3}{4}t_{4}\frac{\partial}{\partial t_{4}},$
+and the unit vector field
+$e=\frac{1}{D}$ $\left(6t_{1}\left(2t_{1}^{4}-3t_{4}t_{1}-10t_{2}^{4}-3t_{2}t_{3}\right)\frac{\partial}{\partial t_{1}}+6t_{2}\left(-10t_{1}^{4}-3t_{4}t_{1}+2t_{2}^{4}-3t_{2}t_{3}\right)\frac{\partial}{\partial t_{2}}\right$.
+$-2\left(16t_{2}^{7}-30t_{3}t_{2}^{4}+112t_{1}^{4}t_{2}^{3}+120t_{1}t_{4}t_{2}^{3}+9t_{3}^{2}t_{2}+30t_{1}^{4}t_{3}+9t_{1}t_{3}t_{4}\right)\frac{\partial}{\partial t_{3}}$
+$-2\left.\left(16t_{1}^{7}-30t_{4}t_{1}^{4}+112t_{2}^{4}t_{1}^{3}+120t_{2}t_{3}t_{1}^{3}+9t_{4}^{2}t_{1}+30t_{2}^{4}t_{4}+9t_{2}t_{3}t_{4}\right)\frac{\partial}{\partial t_{4}}\right),$
+where
+$D=4t_{1}^{8}+56t_{1}^{4}t_{2}^{4}+4t_{2}^{8}+60t_{1}^{4}t_{2}t_{3}-12t_{2}^{5}t_{3}+9t_{2}^{2}t_{3}^{2}-12t_{1}^{5}t_{4}+60t_{1}t_{2}^{4}t_{4}+18t_{1}t_{2}t_{3}t_{4}+9t_{1}^{2}t_{4}^{2}.$
+We also have $e=-\mathrm{grad}_{\eta}\log z_{2}$.
 6. Conclusions
-Given a root system $R$, a weight $\omega$ and a set of pencil generators $z^{1},\cdots,z^{\ell}$, we constructed a generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$.
-We proved that $z^{1},\cdots,z^{\ell}$ depend polynomially on the flat coordinates $t^{1},\cdots,t^{\ell}$ of the generalized Frobenius manifold, we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold can be embedded into the subgroup $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
-In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with an appropriately chosen weight $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail.
+Given a root system $R$, a weight $\omega$ and a set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators, we present an approach to construct generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$, and we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold is isomorphic to the subgroup $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
+In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with appropriately chosen weights $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail.
 For the case $(A_{\ell},\omega_{\ell})$, the generalized Frobenius manifold structure that we will present in [20] coincides with the one constructed in [7], where this generalized Frobenius manifold structure is constructed in terms of a super-potential given by a Fourier polynomial.
@@
   1. i)
-    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrary weight $\omega$?
+    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrarily chosen weight $\omega$?
   2. ii)
-    Precise description of monodromy groups: While we have shown that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ embeds into $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$, can this subgroup be characterized more explicitly?
+    Relation to other Frobenius manifold constructions: Are these generalized Frobenius manifolds related to the Frobenius manifolds constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups?
   3. iii)
-    Relation to other Frobenius manifold constructions: How do these generalized Frobenius manifolds relate to earlier constructions on the orbit spaces of the finite and the extended affine Weyl groups?
+    Almost duality: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
   4. iv)
-    Almost duality and geometric interpretations: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
+    Legendre transformations: What are the properties of Legendre transformations [23, 27] for these generalized Frobenius manifolds?
   5. v)
-    Legendre transformations: What are the properties of Legendre transformations [23, 27] for these generalized Frobenius manifolds?
-  6. vi)
-    Associated integrable systems: Which integrable hierarchies already known in the theory of solitons are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?
+    Associated integrable systems: Which integrable hierarchies already known in the theory of integrable systems are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?
+Appendix A Proof of Lemma 3.5
+We are to prove that
+$\mathrm{Re}(d_{\alpha})>0,\quad\alpha=1,\dots,\ell.$ (A.1)
+To this end, we first show a general version of Theorem 4.1, the conclusion of which is not based on the assertion of Lemma 3.5.
+Since we do not assume a priori that all degrees $d_{\alpha}$ have positive real part, we may come across resonance cases when solving the differential equation of quasi-homogeneous periods, i.e., it could happen that
+$d_{\beta}-d_{\alpha}=j\kappa,\quad j\in\mathbb{Z}$
+for some $\alpha$ and $\beta$, so in such a case the quasi-homogeneous periods may contain logarithmic terms. In what follows we follow the notations of Sect. 4.
+Lemma A.1.
+For each quasi-homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a quasi-homogeneous period $u\in\mathcal{S}_{\Lambda}$ such that
+$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho),$ (A.2)
+where $h^{\alpha}(z,\rho)$ has the form
+$h^{\alpha}(z,\rho)=t^{\alpha}(z)+\sum_{j=1}^{\infty}\rho^{j}\Big(h^{\alpha}_{j,0}(z)+\log\rho~h^{\alpha}_{j,1}(z)\Big).$ (A.3)
+Proof.
+To find a solution $u$ of the form (A.2) for the equations (4.5) and (4.8), we need to solve the following system of equations:
+$\frac{\partial\vec{\xi}}{\partial z^{i}}$ $=\Gamma_{i}(\rho)\vec{\xi},\quad i=1,\cdots,\ell,$ (A.4)
+$\kappa\rho\frac{\partial\vec{\xi}}{\partial\rho}$ $=(A(\rho)-d_{\alpha})\vec{\xi},$ (A.5)
+where
+$\vec{\xi}=\left(\frac{\partial h^{\alpha}}{\partial z^{1}},\dots,\frac{\partial h^{\alpha}}{\partial z^{\ell}}\right)^{T},\quad(\Gamma_{i}(\rho))_{j}^{k}=\Gamma_{ij}^{k}(\rho).$
+We expand $A(\rho)$ into a power series
+$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots$
+and expand $\vec{\xi}$ into the form given by (A.3) as
+$\vec{\xi}=\vec{\xi}_{(0)}+\sum_{j=1}^{\infty}\rho^{j}\Big(\vec{\xi}_{(j,0)}+\log\rho~\vec{\xi}_{(j,1)}\Big).$
+Then the equation (A.5) is equivalent to
+$(A-d_{\alpha})\vec{\xi}_{(0)}$ $=0,$ (A.6)
+$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j,1)}$ $=-\sum_{r=1}^{j-1}A_{r}\vec{\xi}_{(j-r,1)},\quad j=1,2,\dots,$ (A.7)
+$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j,0)}$ $=\kappa\vec{\xi}_{(j,1)}-A_{j}\vec{\xi}_{(0)}-\sum_{r=1}^{j-1}A_{r}\vec{\xi}_{(j-r,0)},\quad j=1,2,\dots,$ (A.8)
+here (A.6) is automatically satisfied.
+Note that $A$ is diagonalizable due to the assumption that $\nabla E$ is diagonalizable, hence for each $\zeta\in\mathbb{C}$, $\ker(A-\zeta)\oplus\mathrm{Im}(A-\zeta)=\mathbb{C}^{\ell}$.
+Therefore, for each $j=1,2,\dots$, whether $d_{\alpha}+j\kappa$ is an eigenvalue of $A$ or not, we can solve $\vec{\xi}_{(j,0)}$ and $\vec{\xi}_{(j,1)}$ uniquely from the equations (A.7) and (A.8).
+Then we obtain a solution $\vec{\xi}$ of (A.5) by induction on $j$.
+Since $\vec{\xi}_{(0)}$ solves (A.4), and equations (A.4) are compatible with (A.5), we know that $\vec{\xi}$ also solve (A.4).
+The lemma is proved.
+∎
+It follows from the above Lemma that for each index $\alpha$, there exist constants $C^{\alpha}_{\beta}$ and $S^{\alpha}$, such that
+$v^{\alpha}(z,\rho)=C^{\alpha}_{\beta}\check{x}^{\beta}(z,\rho)+S^{\alpha}.$ (A.9)
+Now we focus on the behaviour of the right hand side of (A.9) when $\rho\to 0$. Note that the Laurant polynomials $Z_{1},\cdots,Z_{\ell}$ given in (4.6) induce the canonical projection
+$Z=(Z_{1},\cdots,Z_{\ell})\colon\mathbb{C}^{\ell}\to\mathbb{C}^{\ell}/W_{a}(R),$
+where $\check{x}^{1},\dots,\check{x}^{\ell}$ are coordinates on $\mathbb{C}^{\ell}$ and the affine Weyl group $W_{a}(R)$ acts canonically on $\mathbb{C}^{\ell}=\mathbb{R}^{\ell}\otimes\mathbb{C}$.
+Denote the inverse map of $Z$ by $\check{X}$, then we have
+$\check{x}^{\beta}(z,\rho)=\check{X}^{\beta}(\rho^{\theta_{1}/\kappa}z_{1},\cdots,\rho^{\theta_{\ell}/\kappa}z_{\ell}).$
+Note that $Z$ is a branched covering map, hence there exists a neighbourhood $U$ of $(Z_{1},\dots,Z_{\ell})=(0,\dots,0)$ such that its preimage is a disjoint union of some open subsets of $\mathbb{C}^{\ell}$, on each one of which merely one point is mapped to $(Z_{1},\dots,Z_{\ell})=(0,\dots,0)$.
+Since $(\rho^{\theta_{1}/\kappa}z_{1},\dots,\rho^{\theta_{\ell}/\kappa}z_{\ell})$ lies in $U$ for sufficiently small $\rho$, we know that on each monodromy branch, $\check{x}(z,\rho)=(\check{x}^{1}(z,\rho),\dots,\check{x}^{\ell}(z,\rho))$ converges to a constant point when $\rho\to 0$, regardless of the value of $z=(z_{1},\dots,z_{\ell})$.
+In particular, this constant point is one of the common roots of $Z_{r}$ for $r=1,\cdots,\ell$.
+Therefore, the right hand side of (A.9) converges to a constant $V^{\alpha}$ when $\rho\to 0$ on each monodromy branch.
+It follows from Lemma A.1 and (A.9) that
+$\rho^{d_{\alpha}/\kappa}\left(t^{\alpha}(z)+\sum_{j=1}^{\infty}\rho^{j}\Big(h^{\alpha}_{j,0}(z)+\log\rho~h^{\alpha}_{j,1}(z)\Big)\right)\to V^{\alpha},\quad\textrm{when}\ \rho\to 0.$
+Since $t^{\alpha}(z)$ is not a constant, we have $V^{\alpha}=0$ and $\rho^{d_{\alpha}/\kappa}\to 0$ when $\rho\to 0$ on each monodromy branch.
+This implies that $\mathrm{Re}(d_{\alpha}/\kappa)>0$, hence $\mathrm{Re}(d_{\alpha})>0$.
+The theorem is proved.
+Acknowledgement
+This work is supported by NSFC No. 12571266.